A Complete Bibliography of Publications in The Computer Journal: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 2.04

Title word cross-reference

1 − 2 [OW80]. 2 [MF88, TH82, Zak84]. 3 [Dob84, Zak84]. 4 [Lu89, Pik81]. d [RDT80, EHH86]. K [Er83a]. M [AO80b]. n [BJ80, Var80, Wat81]. P [AF88]. T [Er87b].

-Ary [Er87b]. -Bit [Pik81]. -component [TH82]. -Compressed [AF88]. -dimensional [Wat81]. -Inhibition [RDT80]. -Level [Dob84]. -partition [AO80b]. -peg [Lu89]. -sided [BJ80].

1 [Cra86]. 1000 [Mil82b]. 11 [WD80]. 117 [WH87]. 121 [Hun85, Wic87]. 122 [Spo85]. 1975 [FJA80]. 1977 [FE80, NQ80]. 1978 [FJA80, WD80]. 1979 [BB80a, BL81a, Eme80, FE80, FJA80, NQ80, Pau80b, VP80, WD80]. 1980 [CF81, JR81, NPCA81]. 1981 [FSWL82, Kil81b, NPCA81]. 1984 [Ano85m].

2 [Bai88, Tra84]. 2900 [IS81]. 2nd [CF81, Che82, Lea80].

3 [Ree83].

68 [GP80].

77 [VP80].

80 [HW89].

Abstract [BL81b, CGK80, Fur89, Les83a, Mai85, MTW88, RvS88, Rob88, YP86, CFL84, WS86, Yel89]. Abstraction [FM86b, JL80, KGC85, Ree80c]. Abstractions [Sho81, Wel80a, SC87]. Access [GP88, KS82, Kie88, LW85b, MK82, SW84, O’S88, ZR88]. Access-Control [KS82]. Accessing [CN82, Mul87, WN81]. Accounting [Bar82a, Kol83]. Achieving [Lyo85]. Acid [CCL87]. ACM [Ano80e, Har80]. Acquisition [AB84, Bow83, KC88]. Action [WHF83]. active [HTW89]. Activity [Tul85, BWM88]. actor [KL89]. Ada [BM86, Bro84a, Hun85, RB85, SW83, Wic82, Wic87]. Adapted [Ver87]. Adaptive [Akl87, BM88, FA88, Mäk89, Wit80]. Adder [Fen87, Sal87]. Adding [LE86]. Address [Par88, SAA88]. Addressing [Bra85, Bra86b, CFL84]. Adequacy [DW88]. Admin [PPB84a, PPB84b, PPB83]. Administration [Cha82, Aki88]. admissible [Er89a]. Adoption [AB84]. Advanced [Ing83, NQ80, WW83, Bai88]. advances [Mur83, Wis82]. advice [BM80]. Affine [Skl81]. after [SPMD83]. agglomerative [EHW89]. Aggregate [Bow84]. Aggregate-Incomplete [Bow84]. Aid [DP88, Mul81, Tra82, Tra84, HM80a, KH81, Tom82]. Aided [FR80]. Algebra [BBK87, Bar84, Bar85]. Algebraic [Bir89, HK88, FM85]. Algebraically [Fos86]. ALGOL [GP80, DM85]. Algols [FJA80]. Algorithm [Abb80, Ano87a, BGR86, BMT85, Bro83, CL83b, CD80, CHPS85, CG85, CD85, DMS^{+}89, DB81, EL89, Er83a, Er84d, Er88a, Er88b, EN82, Fro81, GC87, Hat84, HPR88, HPB84, HWW87, IZP81, IP86, JBJ84, Les83b, McC82, Pie82, PG82, PK85, SM88, VR81b, Wil85a, Wil83d, Wil86a, Zak84, ZOG85, Ano80c, Com85, ET88, Er88c, Er89b, Hut86a, Iga80, KR89, Lut80, NSW87, SE82, Sne80, SAA88, Thi81, VR81a, Wes80a, Wil86b, Hun85, Spo85, WH87, Wic87]. algorithmic [Oul87]. Algorithms [Akl87, Ano80a, Ano81a, Ano81b, Ano82a, Ano84a, Ano84b, BE82, BF81, Bri85, CR87, CL83a, CV86, DM85, Dro87, HM83, ME88b, Mur85, NS87, Oxl89, Ree84a, Ric88, Wil83e, Wit80, Atk80, FL80, GP80, Mur83, WM84, O’K85]. allocated [Dav82]. allocating [KH81]. Allocation [Bas86, BGR86, CC87, CC88, CD82, Leu82a, McI82a, Pfl84, Ree83, Ree84b, Ree87, Ree80b, Wan80]. Alternative [Hou86, MD87, Sch85, Zar80]. Ambiguity [Sch85, MV88]. Among [BT81]. Analyses [Ree84a, O’K85]. Analysing [LE86]. Analysis [ABD84, AB84, BS86, BF81, Bri82, CC87, CGK80, DP88, EW80, FSW85, HD82, Ing80, Jac80c, Lar85, Leu83a, LW85a, Lom88, MG85, Mil81a, PW81a, Pag84, Par82, PAW86, Pfl84, Pot82, Ric88, Sco82, SAA88, WHF82, ZE88, ZOG85, Coo83, Gea80, Ham80a, HM80a, JC86, Leu82b, Lut80, Mum80, SI80, Spu88, Wat80]. Analyst [BS87]. Analytic [BD81, FS81]. Analytical [Sto86]. and/or [Kol83]. Anne [NQ80]. Annotated [SS80]. Announcement [Ano81c, Ano84c, Ano85a, Ano85b, Ano86a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p]. Announcements [Ano83a, Ano85c, Ano85d, Ano85e, Ano85f, Ano86b, Ano86c, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano88g, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano88s, Ano89q, Ano89r, Ano89s, Ano89t, Ano89u, Ano89v, Ano89w, Ano89x, Ano89y]. Ansi [BP81]. Anthony [VP80]. APL [JK80]. Apl80 [Ray82]. applicability [Ros82]. Application [Sel88, CAC82, GF87, SA81, Wat81]. Applications [All88, Bar83a, EHH86, Kol83, SD82, SM88, Har88, Bri82]. Applied [BES83, FJA80]. Approach [AMN81, ABC^{+}83, Ben82a, BR86, Bro84a, Che86a, Che86b, Col83, Er83c, Er84c, FE80, FW86, Fis83, KB87, ST89, ST82, Tse87a, UVS87, WT84, WL87, Wil84, AO80b, BD80, Er82b, Fun85, FM85, HM80b, JC88, Pro80b, Qui83, Rus86, Spu88, WN88, Mil82a]. Approaches [WHF82, Coo83, EWH86, Gea80, KH87, Wat80]. appropriate [EWH86]. Approximate [BWM88, Wes80b]. Approximation [BMT85, KR89]. Arbitrary [HPR88, Pag82b, AO87]. ARCA [MHS85]. Architectural [PGM84]. Architecture [Bow84, BWW81, DAODT85, UHS85, WW83, BBKR89, BB80b, CFL84, MMH87, ME80]. Architectures [Tre82, TAP89, CFS89]. arcs [Abb80]. Area [GP88, IEH^{+}85, Mul81]. Areas [OW83b, dS80]. Ariadne [Cra86]. Ariadne-1 [Cra86]. Arithmetic [Clu80, Goo80, SD82, Vel82, Spo85]. Arjuna [DPSW89]. arrangements [VR81a]. Arrays [Bla82, ME88a, ZBS81, Iga80, Mul87]. Arthur [WD80]. Artificial [Bur80b]. Ary [Er87b]. aspects [Mum80]. Assembly [WD80, WD80]. Assessing [Rze85]. Assessment [ABBS83, LW83]. Assisted [Smi83, LM88]. association [EDS86]. Associative [Lea86, MMH87, Eme80]. Atre [NPCA81]. attachment [PW81b]. Attitudes [EDS86, JC86, Lem80]. Attribute [BDS82, PK86, WM83, KC88]. Auditing [Cha81a]. augmenting [Hut86a]. Authors [Ano82d, Ano81m]. AUTO [TCL89]. AUTO-DFD [TCL89]. Automated [Hai83]. Automatic [Cho82, Clo86, Inc87, Rai87, Ray80, SI80, Sch82a, FW83, RB86]. Autumnal [FM86a]. Average [Bra85, Bra86b]. AVL [OW80]. Award [Ano85m, HT81]. Axiomatic [FW86, Pra84]. axiomatising [Gib88].

B [NPCA81, BD84, LG86, MZ86, Mul81, TB86, ZOG85]. B-Splines [MZ86]. B-Tree [TB86, ZOG85]. B-Trees [BD84, LG86, Mul81]. Backtrack [Hel84]. Backtracking [Er87c, Irv84, Er87a]. Balanced [EA83, GOW83]. Balfour [VP80]. Barron [NPCA81]. Base [AK81, Bak82b, Cha82, Dav81, Fro85, GW88, Jon82, NPCA81, PSD^{+}81, PJ80, SM88, VP80, Wel80a, Fro86, Mil81b, Pro80b]. Based [Cha87, DW88, ES81, FSWL82, Fur89, LW83, Lom88, MK85, MT86, OM88, Ric88, WC86, YP86, BWM88, BMNS87, CFL84, Fun84, GGH^{+}87, KL89, Ken80, KGC85, MV88, Mic82, TAP89, YY89, MBC^{+}86]. bases [Ano80d, MMH87]. Basic [Fre82]. Basic-Pack [Fre82]. Basis [Bog83, Mac81, Mar82]. Batched [MKH86, MKB87]. BCPL [Bro82]. BCS [Ano80e, Har80]. Be [AC84, PJ89]. BEDSOCS [ER80, Har82b, SB81]. behaviour [BM84, CM89]. Behavioural [Les83a]. Berry [Hil82]. best [CP87]. best-fit [CP87]. Between [Bra86b, Bra85]. Beyond [IS81]. bibliographical [YAH80]. Bibliography [SS80]. Binary [Bri85, CC88, Clu80, EA83, Fro82, FP82, Les83b, Mah86, MKH86, Pal86, Pal88, Sal87, Sha81c, Sip88, ZOG85, CM89, Er89b, FL80, Hat84, RM89]. Binary-Relational [Fro82]. Binary-ripple [Sal87]. Binding [ABM88]. Bisection [CZ82, SE82]. Bit [Pik81]. bivalent [Mot80]. blackboard [Cra86]. BLEND [MP85, Sha82a, SPMD83]. BLEND-LINC [SPMD83]. Block [Hou86, TH82]. Block-Structured [Hou86]. Blocked [CN82]. Blocking [KW84]. Blood [Sca85]. Bloom [Mul87, SPC89]. Board [RHS89]. Book [Add80, Add81, Ait83, Ale81, And81, Bak80, Bar80, BB80a, Bar83b, Bar82b, Bea81, Bel81b, BL81a, Bel82, Ben81, Ben82b, Bly80, Bra80a, Bri81a, Bri81b, Bri81c, Bri82, Bro89a, Bur80a, Bur80b, CF81, Cha81b, Che82, Cla81, Clu81, Cor81, Day80, De 80, Den80a, Den80b, Eas80, Edw81, Eme80, Eva81, FSWL82, FE80, Fel80, Fis80, For80, FJA80, Fre81a, Fre81b, Geo81, Gil82, Gil83a, Gou81, Gra80, Hil82, Hol81, Hop84, Ing81, Jac80d, Jen84, Jon81a, JR81, Jon81b, Ken81, Kil81a, Kil81b, Kin81, Lan84, Lav81, Leh81, Lin81, Lov81, Mil82a, MH81, Mux81, Mux84, New80, NQ80, New81, NPCA81, Pan81, Par83a, Par83b, Pau80a, Pau80b, Pot81, Pro80c, Pro80d, Qui80, Qui81, Rad81a, Rad81b, Ree80a, Ree81, Ril81, Ros81a, Ros81b, Sha81a, Sha81b, Sha82b, Sim81, Sma81, Tho80c, Tho80d, Tho81]. Book [Val80, VP80, Wal80, Wal81, Wel80b, Wel84a, Wel84b, Wil83a, Wil83b, Wis80, WD80, Wis81, Wis82, Woo81]. Books [Ano81d, Ano81e, Ano81f, Ano81g]. Boolean [Bla82]. Bottom [UVS87]. Bottom-Up [UVS87]. bound [MMB87]. boundary [Thi81]. Bounded [CD82, DE86]. BPL [Wil82b]. branch [Tho80a]. branches [Wig81b]. Break [AB84]. Break-Even [AB84]. Breaking [HM83]. Brelsford [FSWL82]. Bresenham [PG82]. British [Ano81h]. Brodie [JR81]. brother [OW80]. Brown [NQ80, Bur80b]. Browsing [LSV89, DB88, MP85]. Broyden [CG85]. Bubble [Wri83]. Bucket [Lyo85]. Buckley [NQ80]. Buddy [CR87, PBH86]. Buffer [DE86]. Buffered [HPB84, WH87]. Buffering [Wri82]. building [Kec87]. Bulterman [Er88c]. bus [Mar82]. Business [Che82, Lic89, Mil82a, Mum85, EDS86, FE80]. Byte [BS89].

C [Ben82b, Mil82a, SAM88]. CAD [Bog83]. CAD/CAM [Bog83]. CAL [GF87]. Calculate [CZ82]. Calculating [Oxl89]. Calculation [Bir89, Hun85, SAA88, Wic87]. calculations [MMB87]. calibration [GP81]. Calingaert [Gil82]. Call [Ano81h, Ano82b, Ano84d, Ano85g, Ano85h]. calling [BB80b, ME80]. Calls [Ano80b, Ano81i, PJ86]. CAM [Bog83]. CAMPS [LM88]. can [PJ89]. Cannot [RHS89]. canonical [Dee80, DNV81]. Capabilities [FC87]. Capability [APW86, MT86, CFL84]. Capability-Based [MT86]. Carberry [BL81a]. Carry [Fen87, Sal87]. Carry-completion-detection [Sal87]. carte [LSV89]. Cartesian [CC88]. Case [CCL87, Lar82, Mum85, TB82, Cla88, Erk84]. catalogues [Bir81]. Categoricity [Les83a]. cautious [Qui83]. Caxton [Ben82b]. Cellular [ZBS81]. Centralized [IP86]. Centre [Muc82]. Centrenet [IEH^{+}85]. centres [Jam81]. Centroid [Mot85]. Certain [FD88]. CFG [HKPW89]. CGI [Tom82]. Chained [Sur86]. Chainmail [Ree87]. Chains [Bra88, Leu83a, MKP87]. Chairman [Lan80]. Challenge [NQ80]. Chalmers [AJ89]. Change [Mul81]. changing [Bri80a, Bri80b, Ham80b]. channel [RM89]. Character [WW82, YAH80, Gri80]. Characterise [Fro85]. Characteristics [Lle87, RS88]. Check [War86]. Checking [ABM88, Cow86, Hou83]. Chess [Mic81, RHS89, BM80]. China [GGH^{+}87]. Choice [Sto86]. choose [EWH86]. Chow [Pau80b]. Chromatic [Col82]. Chuen [Pau80b]. Chuen-Yen [Pau80b]. Ciphers [HM83]. Circuits [Clu80]. Circulating [WI84]. clashes [Stu84]. Class [CD82, IW86]. Classes [Er89a, FD88]. Classical [Van86]. Classification [MD87, CD80]. Classificatory [AW85]. Clauses [Ber87a]. Clear [GGH^{+}87]. Clock [CF88]. Closed [Bra88]. Closure [FD88]. Clues [SD86]. Cluster [Sco82]. Clustering [BMSA88, FM86b, Mur85, Ree84b, EHW89, Mur83]. clusters [SAM88]. CMOS [Fen87]. Co [Bra88, WG83]. Co-expressions [WG83]. Co-Relationship [Bra88]. Coalesced [CV86]. Cobol [Eme80, Ing80, Ing83, Tri80, Tri83, TL85a]. CODASYL [Aki88, GE85]. Code [Akm86, EW80, Vel82]. Codes [Lle83, Lle85, War86, RM89]. Coding [Pik81, PG82, YAH80]. Cognitive [Lav83, BWM88]. collection [NSW87]. Collision [Sam81]. Colour [Er84d, FS87, Er84a]. Colouring [DB81, EL89, Sel88, Wil85a, WM84]. Colours [Wil85a]. COMAR [HKPW89]. Combinations [Akl87]. Combinator [Bur87a]. Combinatorial [Fle82, FM83]. Combining [THR82]. Comma [Lle83, Lle85]. Comma-Free [Lle83, Lle85]. Command [MA89]. Comment [BK83b, BK83a, O’K85, Wil82a, Rei81]. Comments [Bre81b, Mic81]. commercial [Ham81b]. communicating [Nat86]. Communication [Flo80, PB86, Sha82b, Sil81, Mar82]. Communications [BB80a, Par88]. Compact [Sto86]. Compaction [Bre82, Col87, QY89]. Comparative [Laz81, KH87]. Comparing [Min88]. Comparison [DF87, EHW89, JW82, Lom88, MK82, Pra88, Ric88, GGS89, RD80, Wic82]. Comparison-Based [Lom88, Ric88]. Compilation [AMN81, Ber87a, BY89, Smi83, Wil89]. Compiler [AJ89, BY89, Rob80b, BHY89, SC87, SS81]. Compiler-Compilation [BY89]. Compilers [Cap80]. Complete [Pag81]. completer [Zar80]. completion [Sal87]. Complex [BGR86, ZBS81, BM80, HM80a]. Complexity [BR86, CCD86, CCS87, MK85, Nav87, Pra84, Ree84a, O’K85]. Component [UVS87, TH82]. Components [Fro85]. Composite [FJA80, McL85]. Compositional [RL89]. Compressed [AF88, Woo84, Mul87]. Compression [BD82, CH87, FM83, Lle87, Pik81, Rai87, BM89, Mäk89]. Computability [Goo85a]. Computation [Er83c, ME88a]. Computational [Jam89, Pau80b]. Computations [JJ86]. Compute [Bra85, Bra86b]. Computer [Ano81i, Bar82a, BL81a, Ben82d, Bri82, BK83a, CF81, CD81, Cav83, Cha81a, CW82, Eme80, EM81b, FSWL82, FR80, FJA80, FS81, GK84, GG86, Gil83c, Jac80a, JJ86, JR81, KVW89, Laz81, Lea80, Lic89, Mar80, Mit82, MTPL88, Par89, QY89, RHS89, Smi83, Tho80b, TF87, UHS85, Wel82, Woo82, BP82, BK83b, Cap86, Er86b, GGH^{+}87, HM80a, Hea81, HM80b, JC88, Jam81, Kec87, KH81, MR85, Wat82, WN81, LM88]. Computer-Assisted [Smi83, LM88]. Computer-Based [FSWL82]. Computerised [BD81]. Computers [Ben85, Flo80, FRS^{+}87, Fre82, Kec87, Sow87]. Computing [Abb80, Bow81, Col82, Er82a, Er83a, FSWL82, Muc82, PK83, RS88, THR82, Wat81, Wic89a, WN86, Bur87b, Noo84, Pro83, CF81]. concept [BPG87, Fun84, SA81]. Concepts [Fro85, HS82, Rob80b]. Conceptual [FL85, Fro83, Min88, Sha81c, SC87]. Concrete [BBK87]. Concurrency [HM84, IP86, PPB84b, RvS88, Wet81]. Concurrent [BIY88, HW89, Hul86, KS85, Shi85, Bur87b, KL89, KS82]. Condition [Gra87]. Conference [Ano81h, Ano83c, Ano85i, Ano85j, Ano85k, Ano80d]. Conferences [Ano83b]. configuration [MBC^{+}86]. Connected [UVS87]. Connection [Bra88]. Considerations [AB84, Wri83, PGM84]. considered [Lip80]. Consistency [BKK84, LW85a]. Consistent [UHS85]. Constant [Mul85]. constituent [JC86]. Constrained [Mur85, Wri82]. Constraints [Orm88, Sho81, PPB84c]. Construct [Lo84]. Constructing [FL89, EL82]. Construction [EA83, Lle85, Rob80b, SS80]. Constructs [DE86]. containing [WC85]. Contemporary [FSW85]. Context [AMN81, Ram85, SEPK89, EDS86]. Context-free [SEPK89]. Contiguity [Mur85]. Contiguity-Constrained [Mur85]. Continuations [All88]. contour [RB86]. Contouring [Sch82a, RB86, ST81]. Control [Ben82c, Ben82d, ER86a, EM81b, FK86, GP88, GN83, IP86, KS82, MK85, MT83, PPB84b, RvS88, THR82, TF87, Ken80, O’S88, SP82, YAH80, Den80b]. controller [SA81]. Controlling [O’S88]. Conventions [HH80]. Convergence [Jam87]. Converging [Jam89]. Conversion [CCD86, Par89, SW83, Ste86]. copies [KH81]. Coping [BR86]. Copying [Har81a]. Corecs [PPB84a]. corporate [Car86, Ham80a]. Correct [Bra80b]. Correctable [TB86]. Correction [Sub85, WT84]. Correctness [Lew84, MP89, O’S88, PJ86]. corrector [Ano80c, Sne80, Wes80a]. Correspondance [Bel81a, Har81c, Rob83, Wig81a]. Correspondence [Apo89, Bae86, BM83, BM81, Bow87, Bre81a, Bre81b, Bro86, Bro88a, Bro88b, BJ80, Cai83, CP86, Coa83, Cr081, Cro84, Du 86, Els86, Eng81, Er84b, Er85a, Fel84, Fer86, Fin83, Fro87, Gro84, Har83, Har81b, HT81, Hea84, Hea85, Hea86, Hil80, Hil81, Hol84, Hub88, Hub89, Hun81a, Hun81b, Ibb81, Jam88, JL80, Kil84, Kio81, KS87, Lav86, Lei88, Lis84, Lun86, Mac83, Mac85, Mea83, Mep84, ME80, MR82, Mit89b, MM89, Mos86, Phi84, Rec84, Ree80c, Ree84c, Rob80a, Roh88, Ros87, RvB89, Sha85, Sne80, SB81, Tay80, Tay81, Tho80a, Tow87, Tsi86, Wak81, Wat82, Whe89, Wig81b, Wlo89, Bir81, Lem80, SI80, Zar80]. Corrigenda [Ano81j]. Corrigendum [Ano84e]. Cosets [PPB84a]. Cost [ABBS83, CN82, Lea86]. Cost-Effective [Lea86]. Could [AC84]. counter [Sow87]. countries [Noo84]. Course [NQ80]. Cover [RHS89]. Covering [SM88]. CPU [MMB87]. CPU-bound [MMB87]. creation [Noo84]. Criteria [DW88, KH81, Lom88, Rze85]. Criterion [Tse87b]. Crossword [Ber87a, BY89, Smi83, SD86, Wil89, SS81]. Cryptic [SD86]. Cryptography [Gro82]. CSP [Hul86]. Currency [LW85a]. Current [Cha81a, Tri80, WHF82, Coo83]. Curve [MZ86]. Curve-Smoothing [MZ86]. Curves [BT81, Col87, FS87, GM83, QY89]. cutting [AO80a]. Cweb [Thi86]. Cybernetic [FE80]. Cycle [DM83, Har84b, Ros82]. cycles [Ros82]. Cyclic [Er84c, Pag84]. Cyclide [McL85].

D [FSWL82, JR81, NPCA81, VP80]. D. [Bri82]. DAP [CHPS85, DM85, KB87]. DAP-Algol [DM85]. Data [AK81, Bak82b, BIY88, Bel84, Bir81, Bow84, Bow83, BL81b, Cha82, CH87, Dav81, DW88, Dea82, FM86b, Gol85, HM84, Inc87, IS81, Jon82, Kol83, KA86, Les83a, LW85a, Lic89, Lle87, MG85, Mai85, Min88, Par82, PB86, PAW86, PSD^{+}81, PPB83, PJ80, QY89, Sch82a, Sha82b, Sho81, Sub85, ST82, THR82, VP80, Wel80a, Wis82, YP86, YY89, Zah81, Ano80d, BM89, BP81, Dee80, Ham81b, HKPW89, Hut86a, JL80, Ken85, KGC85, KH87, Mäk89, Mil81b, Pro80b, Ree80c, RB86, Spu88, TCL89, TP89, WS86, ZR88, Bar82b, BB80a, NPCA81]. data-compression [Mäk89]. Data-Flow [PB86]. Database [ABM88, Bak82a, BMT85, Bow84, CN82, CCL87, DAODT85, HS82, IP86, JR81, LW83, Mai85, Orm88, PF86, RvS88, Seg87, She85, Aki88, Ber87b, Bro82, DNV81, DTIR88, ET88, Fun84, GE85, HK82, WS86, ZR88]. database/administration [Aki88]. Databases [Ano81h, Ano83c, BMSA88, Bra86c, Bra86a, Bra88, Dee85, Kie88, LSV89, LW85a, Sch83, Car86, CAC82, FW83, WN88]. Dataflow [PK85, Sow87]. Davis [Eme80]. DBMS [CS85, PPRB89, PPB84c]. Deadlock [TF87]. Decimal [SD82]. Decision [Lew84, Sch85, TB82, MV88]. Decisions [AB84, Che82]. Decomposition [Den80b, DĎ88a, FW86, NS87, PG81, SGW82, Wil82a]. Deduction [Pau85]. Defined [Gol85, GN83, Har84a]. Defining [Mum85]. Definition [Fro83, Geh83, MM84, PPB83, ST82, Gal87]. Definitions [Pag81]. degree [MF88]. Degrees [Mah86]. Delaunay [Wat81]. Delay [Bas86]. Delays [CD82]. Deletion [CV86]. Delphi [Clo87]. DeMarco [TP89]. Denotational [All83, MA89, All85]. Department [GG86]. Dependencies [Bra86c, Bra86a, Bra88, Jaj86, HW88]. dependency [CAC82, O’S88]. depth [Er89a]. depth-first [Er89a]. Derivation [Dro87, Roh87, WW82]. Derrick [Eme80]. described [PJ89]. Describing [SWB87, JC86]. Description [DBH85, Mot80, Mot85, Uck85]. Descriptive [BS86, GGS89]. descriptor [All82]. Descriptors [BB81, BB82, IS81]. Design [BMT85, Bri80a, Bri80b, Bri82, CF81, CD81, CS85, Dav81, FR80, GM83, Gol82, Hai83, HS83, HM84, Hut86b, JK80, Jon82, Le 88, Mac81, Mil81a, MT86, Mum84a, NPCA81, Par82, Ran86, Sal87, SWB87, TL85a, TL85b, TAP89, Wel80a, Woo86, BWM88, BBKR89, DNV81, EH88, Gea80, GP81, JC86, JL80, KGC85, MV88, Mum84b, Ree80c, RB85, Ros82, Rus86, SI80, Teu87, Wat80, BB80a, Eme80, FJA80, JR81]. Designing [Kol83]. Designs [Clo86]. destination [Mit89c]. Detailed [LW83]. Detecting [Jan88]. Detection [EW80, WT84, BD80, Cze88, Sal87]. Determination [SC81, HK82, SE82]. detranslation [HM80b]. Developing [Bog83, Sel88, BWM88, Noo84]. Development [Ano80e, Bog83, Dea82, DM83, DM85, Har80, AFWH88, Har88, RB85, Ros82, SC87, Tyl88]. Devices [Sam81]. devolved [Car86]. Dewey [Foo86]. DFD [GGS89, TCL89]. Diagnosable [Lom88]. diagonal [CL83b]. Diagrams [DP88, Min88, TP89]. Dialogue [HS82, RB85]. diameter [MF88]. Dictionaries [FM83]. Dictionary [GN88, MG85, Zah81, Spu88]. Differencing [ME88b]. Differential [BD81]. Difficult [Bel84]. Diffusion [LW85b]. Digital [Ben82c, CF81, Clu80, DBH85, Lea80, WL87, Mor89]. digitised [AO87, Lut80]. digraphs [MF88]. dimension [Com85]. dimensional [Ken85, Wat81]. Dimensions [EHH86]. direct [Mad80, Spo85]. Directed [HS83, Sil81, PJ86]. Directions [Hop80]. Directory [Zve80, DTIR88]. diregular [MF88]. Dirichlet [Bow81]. Disc [CC87, CC88, Kie88, Leu83a, LW83, Leu83b, LW85b, PW81a, Leu82b, Per80]. Disciplines [MK82]. Discrete [Fle82, Sch82b]. Discussion [Bir81, Lem80, SI80, Wat80, Zar80]. Displaying [NW87]. Distance [Bra85, Bra86b]. Distributed [BIY88, Dav81, DAODT85, HPR88, IP86, LW85a, Mah84, MT86, NH89, PB86, RvS88, Sil82, TB82, BPG87, Ber87b, BBKR89, BP82, Bur87b, CFS89, DTIR88, ET88, Mar82, Nat86]. Distribution [Har84b, Mah86, Ree83, Com85, Ree80b]. Distributions [SB85]. Distributive [McC82]. DMC [BM89]. Document [Fur89, Mac81, All82, EHW89, SvR83]. Documentation [CW82]. Documents [Bro89b, SPC89, Mul87]. Does [Hut84]. domain [Fro86]. DOPLs [Lee82]. Double [ME88a, Wri82]. Double-Pipeline [ME88a]. Doubly [Sur86]. Douglas [CF81]. Down [ABD84, Bar87, HW88]. Dr. [Fox87]. Drawn [Les83b]. Driven [SEPK89]. DSA [BS86]. Dumping [Fro81]. Duties [WR86]. Dynamic [AB84, Bar83a, BGR86, Che86a, Che86b, CH87, DC84, Dav84b, Leu82a, McI82a, Mul85, Pfl84, RL82, Ree84b, Wan80, Dav84a, Fro81]. dynamically [Dav82].

E. [Mil82a]. Earthwords [Oxl89]. economic [Mor89]. ed [CF81, FSWL82]. edged [WW82]. edited [Eme80, JR81, NQ80, NPCA81]. Editing [AMN81]. Edition [Lea80]. editor [Sha82b]. Editorial [Ano87m, Col89, Coo89a, Ham81a, Ham82, Ham84, Ham89, Wad89, Wic89a, Ham80a]. Edn [Che82]. EFDM [KA86]. Effect [Kol83, RM89]. Effective [Jon82, Lea86]. effects [SvR83]. Efficiencies [Jam89]. Efficiency [LW83]. Efficient [Cze88, Er87c, EA83, Har81a, Har82a, Mul85, Ram85, Sal87, SSDS89, Sip88, Ano80c, CP87, Er87a, Ken85, Sne80, Wes80a]. Eiffel [Coo89b]. eigensolution [CHPS85]. Eigenvalue [BES83]. Eigenvalues [CZ82, SC81, CL83b, SE82]. Eight [RHS89]. Electre [ER86a]. Electronic [BB89, Col89, Gou80, GGH^{+}87, Mit89c, Sha82a, SPMD83]. Element [She85, ST81]. Elemental [Abe85]. Elements [Ben82d, Gil82, Gil83b, CL83b]. Eliminating [HY81]. Embedded [Bro84b]. Empirical [IH86a, IH86b, PW81a, EDS86]. Employing [Tyl88]. Encoded [OW83a]. Encoding [Bae84b, Sto86, Bae84a, CP87, KR89]. encryption [YY89]. end [WN81, Bak82b]. End-User [Bak82b]. Engineering [Ben82f, BR86, Col80, Fal81, Rus86, Tyl88]. English [WN81]. Enhanced [Par88]. Enhancement [TL85a]. Enhancing [TL85b]. Enough [Lan85, AFWH88]. Enquiry [Sta85]. Entities [Sha81c, Com85]. Entity [DP88, FM86b, Par82, Ros82]. Entity-Relationship [Par82, DP88]. entropy [Fun84, Fun85]. Enumerating [Akl87, Er85b, Pal86]. Enumeration [SM88]. Environment [BIY88, UPT89, Bri80a, Bri80b, DB88]. Environmental [GW88]. Environments [KS85, RS88, GF87]. Equation [Jam87, Shi89]. Equations [BD81, DAEH81, Abb80, IZP81, JBJ84, RD80]. Equifrequent [YW82]. Erratum [Ano80c]. Erroneous [Kol83]. Error [ABBS83, BDS82, KS85, Mul81, WT84]. Errors [Lei86, BD80]. essential [JC86]. Estimating [CN82, LW85b]. Ethernet [Par88]. Eurographics [Tea82]. Evaluating [RS87]. Evaluation [Cav83, IH86a, IH86b, Mah84, Pap82, YW82, Ber87b, Rei81, Wes80b]. Evaluative [Mic81]. Even [AB84, Com85]. Event [Ree84a, O’K85]. Events [Ano84f, Ano85l]. Evolution [ST89, FE80]. examination [AW85, Stu84]. Example [Muc82, Mum85, Par82]. examples [GF87]. Exclusion [HPR88, PJ89]. executable [GCS89]. Execution [WC86, Gal87, Teu87]. Exegesis [Sma88]. Exercise [HS83, Oxl89]. expansion [SvR83]. Expected [Bri85, Lar82, Mah86]. Experience [BM88, FR87, SA81]. Experiences [Thi86, VR81b]. Experiment [CF88, PRO80a]. Experimental [YP86, Aki88]. Experiments [HM83, Mic82, GP81]. Expert [BM81, Mic80, BWM88]. Explaining [CM89]. Explicit [BK83a, Woo82, BK83b]. Exploiting [Ben85]. exponential [Lip80]. Expression [DE86]. Expressions [HD82, Vel82, Hou83, Rei81, WG83]. Extended [KA86, WM83, Hou83]. extensibility [Lay85]. Extensible [NF80]. Extension [Pra88, Gal87]. external [TH82].

F [BL81a, FE80, Sha82b]. F. [Sha82b]. F.R.S. [Fox87]. Facilitating [Fro85]. Facilities [PPB83]. Facility [TL85a, PPRB89, Uck85]. factor [BA88]. Factoring [MD84]. Factors [Le 88]. Faculty [Sel88]. Families [FS87]. Fast [DMS^{+}89, Er83a, Er88a, Fen87, GC87, Mad80, Ull88, UPT89, Coo89c]. Fast-Carry [Fen87]. Fault [Ben82e]. Favour [DM83]. Feature [FSW85]. FEDOS [She85]. feedback [SvR83]. Fewer [Pri81]. Fibonacci [Er83a]. Fields [WN86]. figures [BJ80]. File [DC84, Dav84b, HK80, Lar85, LW85c, MHS85, Rai87, SW84, Sur86, Wri83, All82, Dav84a, KH81, MKP87, TH82]. Files [Bra86b, BHK85, CC87, CC88, Cro80, KW84, Lar82, Bra85]. Filled [Par89, Dav82]. Filling [Col87]. Filter [SPC89]. filters [Mul87]. Finding [Jam89, JP86, CL83b]. Findler [Eme80]. Finite [Lle83, She85]. Firms [Lic89]. First [Ano81h, Ree87, Yel89, Er89a, SA81]. First-Fit [Ree87]. Fisher [Eme80]. Fit [Leu82a, Pag82b, Ree83, Ree87, CP87, Ree80b]. Fitting [Oxl85]. Five [Wil85a]. Fixed [Bla82, PW81a, SPC89]. Fixed-Size [SPC89]. flexible [GF87]. Floating [RS88, Wic89b]. Floating-Point [RS88]. Flow [Deo83, Min88, PB86, THR82, TCL89, TP89]. Flowchart [PG81, SGW82, Wil82a, Wil83c]. Flowcharting [Mil81a]. Fluid [Pau80b, Sta85]. Form [DP88, Geh83, Hou83, Oul87, Dix85]. Formal [Coo89c, DF87, MM84, Rob88, Tse87a, Wic89a, Wic89b, FM85, Lei87, TP89, TL85c]. Formalisms [Fro85]. format [BP81, HKPW89]. Formatting [Cho82]. formed [CCD86, Er83b]. Forms [FR80]. Formula [CC88]. Formulae [Jam89]. Formulation [BGR86]. Formulations [BT81]. Forthcoming [Ano81k, Ano81l, Ano82c, Ano84f, Ano85l]. Fortran [VP80, CL83b, CHPS85, HPB84, IZP81, WH87, BP81, Fis83, Mil82a, Sch82b, SW83, FJA80]. Foster [Ben82b]. foundation [TP89]. Foundations [BL81a]. Four [AW85, BT81]. Fourth [DP88]. FP [TAP89]. FP-based [TAP89]. Fragmentation [Leu83a, Leu82b]. fragments [All82]. Frame [YP86]. framework [EWH86, KH87, Mil81b, Noo84, Ros82]. Fredholm [RD80]. Free [Lle83, Lle85, Ree80b, Ree83, BP81, Pac88, SEPK89]. free-format [BP81]. Friendly [SLH83]. front [WN81]. Full [Dix85]. fully [Yel89]. Function [GP88, Zar80]. Functional [Bur87b, Bur88, EB89, Har82a, HK88, Hug89, JS89, KA86, Orm88, Pey89, Top82, Wad89, BHY89, CAC82, FL89]. Functionality [Sha86]. Functions [Er82a, Sha81c]. Fundamental [Lei86]. Fundamentals [Pot82, FSWL82, WD80]. Further [Goo85b]. Fusion [AW85].

G [Den80b, FJA80, NQ80]. game [Bra80b, EDS86, Mic82]. game-learning [Mic82]. Games [RHS89]. Gamma [Fal81]. garbage [NSW87]. gate [Sal87]. Gates [Fen87, RDT80]. general [BKK87]. Generalised [CC87, DAODT85, Dee80]. generalization [Er84a]. Generalized [Er84d, FD88, Hat84]. Generalizing [HTW89]. generatable [Er89a]. generate [VR81a]. Generating [CF89, Dix85, DMS^{+}89, Er88a, Er88b, McI82b, McL85, SM88, Top82, Er83b, Er89b, FL80, Iga80]. Generation [BHK85, GB83, IW86, Inc87, LB86, Mid87, MF82, SD86, Spo85, VR81b, Vel82, YW82]. Generator [Gra87]. Generators [Gri88, LB86]. Gentle [CF81]. Geographic [Abe85, LG86]. Geometric [FR80, Oxl89, WT84]. Geometry [Bra86c, Bra86a]. German [Muc82]. Getting [Atk84]. Gill [WD80]. Gilmour [FE80]. Glass [BL81a]. Goal [UPT89]. Goal-Oriented [UPT89]. Goblin [Wil88]. GOTO [Pag82a, WC85]. GOTOs [Tay81]. GPSS [Sch82b]. GPSS-Fortran [Sch82b]. Grammar [EB89, Sto86, Wil85b]. Grammars [BDS82, HS83, PK86, WM83, KC88, Pag82a]. Graph [DB81, EL89, BBKR89, GP80]. Graphic [BK83a, GM83]. Graphically [CR83, PJ89]. Graphics [Col87, CW82, Jac80a, JR81, KVW89, Par89, QY89, Rey83, Woo82, BK83b]. Graphs [ADD86, Wil85a, WM84]. Grid [FA88]. Grids [Bra86c, Bra86a]. Group [YW82]. groups [Atk80]. Growth [LW85c]. Guard [BS84]. Guide [Bak82b, Eva83]. Guidebook [BL81a]. guidelines [Wal82].

H [BL81a, Bur80b, Fox87, Kil81b]. Hagen [Kil81b]. Halstead [Pra88]. Halting [Hut84]. Handbook [FE80]. Handling [Sam81, WN88]. Hanoi [Er82b, Er84a, Er84c, Er84d, GC87, Lu89, Roh87]. Hardware [Laz81, PK85]. Hartley [NQ80]. Hash [BD84, Bra85, Bra86b, Lar82, Lyo85, Sam81, Fro81, Mad80]. Hashing [BD82, CL86, CC87, CV86, CHK85, Lar85, Mul85, RL82]. Head [PW81a]. Heads [PW81a]. Height [GOW83]. Height-Ratio-Balanced [GOW83]. help [GF87]. Herbert [VP80]. Heterogeneous [BS84]. Heuristic [Ben82a, AO80a, Hut86a]. Hi [Lic89]. Hi-Tech [Lic89]. hidden [Tom82]. hierarchic [EHW89]. hierarchical [Coo89c, Mur83]. Hierarchically [Lei83]. Hierarchy [FK86, KVW89, MM84]. High [Geh83, IEH^{+}85, Tre82, TL85b, BPG87, Ken80, Wal82]. High-Level [TL85b, BPG87, Wal82]. Higher [Hen88]. Higher-order [Hen88]. Hill [NPCA81]. historical [Mor89]. History [Les83b, Lay85]. Holder [Zah81]. Holes [KVW89]. HOPE [WS86]. Horizontally [Seg87]. Horn [Ber87a]. Horsley [BB80a]. Human [BA88, FE80, Lic89, MTPL88, Rey83, Tul85, JC88, Mic82]. Human-Computer [MTPL88, JC88]. Hutchison [FSWL82]. Hybrid [Gol85, Gom81, NSW87]. Hypergraph [OM88]. Hypertext [Rit89]. Hyphens [Pri81].

I/O [MK82, Per80]. Ibbett [Eme80]. IBM [Sha86]. ICL [IS81]. Icon [Gri80, WG83]. Identification [Tse87a, O’S88]. Identifying [Tse87b]. Identities [Bir89]. ideographic [WN81]. II [All85, Mic82]. ILP [Kil81b]. Image [AF88, OW83b]. Images [OW83a, AO87, Lut80]. Impact [LE86, Ham80b, Ham81b, Hea81, Jam81, Nyg80, Wat82]. imperative [Bai88]. Implementation [CS85, DAT87, ER80, Er87c, FA88, Har82b, MA89, NF80, SSDS89, Sip88, SB81, TB86, TL85a, Ull88, WF89, CFL84, Cze88, Er87a, RB85, Ros82, Sma88, ZR88, NPCA81]. Implementations [Hul86, Pey89]. Implementing [FRS^{+}87, GP80, Mäk89]. implicant [Iga80]. implications [Smi86]. Implicit [SK85, Shi89]. Important [Sha86]. Improved [GN88, Leu82a, MM84, ME88a, ZOG85]. Improving [Fro86, Wes80b]. in/one [CCD86]. Inclusion [CW82]. Incompatible [WR86]. Incomplete [Bow84, WN88]. incompleteness [MV88]. Increasing [BR86, Laz81]. Incremental [AMN81, Bre82, Bro84a, Ear80, PW81b]. Independence [BL81b]. Index [BMT85, Col82, HK85b, JW82, PF86, Zve80, All82, Fro81]. Indexes [HK85a, Mul87]. Indexing [Foo86, HK80, MM84, SB85]. Indicators [FR87, Ree88, Er86b]. Indirection [Gol85]. Individual [Le 88]. Inducing [CD82]. Industry [Bar82a]. Inference [Clo87, HW88]. INFERNO [Qui83]. informal [FM85]. Information [Abe85, AB84, Ano80e, Ano81m, Ano82d, AFWH88, Cha87, CCS87, FSW85, Geh83, Hai83, Ham89, Har80, HN88, Lan85, NH89, Rze85, Sca85, Sca89, Sch83, Sta85, Tul85, VS87, WHF83, Van86, Cap86, Cla88, RSS87, Ros82, SI80, Spu88, WN88, Wil80]. Information-theoretic [CCS87]. Inheritance [Hut89, KL89]. inhibition [RDT80]. input [BP81]. insertion [Zak84]. Instances [Mai85]. Instruction [Bae84b, BS89, Mil82a, Bae84a]. instructional [Rus86]. Integer [PK83, Wil89, FL80, Pro83]. Integers [MD84]. Integral [DAEH81, RD80]. Integration [DM83, HKPW89]. Integrity [Bak82a, FW83, PPB84c]. Intelligence [Kha80, Bur80b]. Intelligent [Kie88, RSS87, TCL89]. Intentions [Lo84]. Interacting [CR83]. Interactive [ER80, GK84, Rey83, SLH83, BWM88, Har82b, SB81]. Interconnected [MCL89]. Interconnecting [ZE88]. Interface [AK81, Dix85, Fro85, MTPL88, PJ80, GE85, GF87, Jam80, Shi89]. Interfaces [BM88, Hop81, SLH83, Dee80, RB85]. interference [Qui83]. Intermediate [Kil81b, SW83]. Internal [FL85]. International [Ano80d, Ano83c]. Interpolation [MKB87]. interpretation [Ber87b, CFS89]. Interpreter [MA89]. interpreters [FL89]. Interprocess [PB86, Mar82]. Interrupted [Wil83d]. Interval [VT87]. Introducing [Gou80, HK80]. Introduction [Eme80, Pau80b, Bai88, Lan80, TL85c]. Introductory [Eva83]. Inverse [Ull88]. Inversion [Pie82, Stu84]. Inverted [JW82]. investigation [Mil81b, Tal84]. Investment [AB84]. Invitation [Cai82]. IPSEs [Le 88]. irregular [BJ80]. ISAM [MKP87]. isolation [JBJ84]. issue [Ano81d, Ano81e, Ano81f, Ano81g]. Issues [VP80]. Iterative [BE82, DMS^{+}89, Er84d, Jam87, Jam89, Lu89, Roh87, vdW88].

J [BB80a, BL81a, CF81, FJA80, Fox87, JR81, Kil81b, NQ80, WD80]. James [Bel82]. jobs [GP81]. Johnson [VR81b]. Johnson-Trotter [VR81b]. Join [Bra86c, Bra86a, Bra88]. Joins [Seg87]. Joint [Ano80e, Har80]. journal [BB89, Sha82a, SPMD83, Ano81i]. JSD [GGS89]. Judicial [Cla88]. Jumping [ASG79, Atk84, Hil80, Ing82, MR82, Rob80a, Tay80]. Justification [Pri81].

Kennedy [CF81]. Key [BD82, CC87, SB85, Hun85, TH82, Wic87]. Khalil [BL81a]. kind [RD80]. kinds [Hut89]. Kit [BS87, AFWH88]. Klint [Kil81b].

Knapsack [SM88, AO80b]. Knowledge [BMNS87, Fro85, KC88, PK86, Fro86, MMH87, Rus86]. Knowledge-based [BMNS87]. Knuth [Thi86]. Koffmann [NPCA81]. Kuo [Sha82b].

L [BL81a, Che82, NPCA81]. L. [Che82]. Labelling [SK85, UHS85, UVS87]. labels [Pag82a]. Laboratory [FSWL82]. LAN [Par88]. Language [EB89, ER80, ER86a, Fro83, Har84a, HS83, Hil82, Hil83, JK80, Kil81b, Lav83, MM84, NF80, NPCA81, PJ80, TL85a, WD80, YP86, BPG87, BHY89, FL89, GCS89, Har82b, Ken80, Lay85, Ser81a, Ser81b, SB81, Wan80, Wil82b, WN81]. Languages [ABM88, Bur88, Cow86, FD88, Har82a, HK88, HH80, Hou86, Pag81, Pey89, SEPK89, TL85b, WF89, KL89, PRO80a, Tal84, TL85c, Wal82, Yel89, FSWL82, Tre82]. Lano [BB80a]. Lapin [Che82]. Laplace [Pie82]. Large [FM83, Rit89, Har88, Kec87, MMH87, TH82]. large-scale [Har88]. Late [Kol83]. Latency [Sam81]. Lattice [Pal88]. Lazy [Wad89, FL89, AJ89]. Lazy-ML [AJ89]. LCF [Pau85]. Learned [Pau85]. Learning [Ing80, MG85, WHF83, Mic82, FSWL82]. Least [ABBS83]. Least-Cost [ABBS83]. Leathrum [BL81a]. Leaves [OW83b]. Left [Vel82]. Legal [Lei86]. Length [Goo85b, Lle85, RM89]. Lengths [Bra85, Bra86b]. Lessons [Les83b, Lic89, Pau85]. Letter [CL86]. Letter-oriented [CL86]. Level [DBH85, Dob84, FR87, Geh83, Tre82, TL85b, Wil85b, Aki88, BPG87, EB89, Ken80, Pag82a, Wal82]. Levy [BL81a]. Lewin [CF81]. Lexicographic [Er87b, Er87c, Er88b, Irv84, Er87a, Spo85, Wil83d]. Lexicographically [Er85b, Er83b]. library [Teu87]. Life [Sca85, Ros82]. Like [Har84b]. LINC [SPMD83]. Line [Bre82, Ear80, PG82]. Linear [Abe85, BE82, DD87, DD88b, HY81, Jak85, Lar85, Wil85a, CF89, Wit80]. lines [CP87]. LISP [Sip88]. Listing [Er87b]. Literate [Knu84, Thi86]. Literature [Sco82]. Livelock [Pac88]. Livelock-free [Pac88]. LKBs [MMH87]. Load [EM81b]. Local [GP88, IEH^{+}85, MHS85, Dee80]. Locality [Leu83b, BM84]. Locally [ABBS83, TB86]. Location [EHH86]. Locking [BS84, Wet81]. Löf [BK87]. Logic [Eme80, Gol82, Lei86, MK85, Sal87, Spi83, War86, Van86, CFS89, Gib88, FSWL82]. Logic-Based [MK85]. Logical [Wel80a, JL80, Ree80c, Tal84]. lookup [Mad80]. Loop [Tak86]. Loops [MT83, SP82]. Loss [Sch83]. Lumped [Ree84b]. Lumped-state [Ree84b].

M [BL81a, Den80b, FSWL82, NQ80]. Machine [GN88, Kha80, PF86, Tre82, PGM84, RB86, Clo87, WD80]. machine-readable [RB86]. Machines [Shi85]. Macro [Bro84b, TL85b, Lay85]. Maddison [Bre81b]. Magnetic [Wri83]. mail [Mit89c]. Main [LE86]. mainframe [Hea81, Wat82]. Maintainable [Hut86b]. Maintenance [BIY88, CCD86, FW83, HK82]. Makes [Sha86]. Making [TB82, Coo89b]. Management [Har82a, Lic89, PSD^{+}81, Wil84, Wil80, Wis82, Ber87b, MBC^{+}86, Noo84, RSS87, NPCA81, VP80]. Manager [Hek89, Dix85]. Manipulation [FS87, Uck85]. manufacturers [Hea81, Wat82]. map [BKK87]. Mapping [FL85]. mappings [Gri80]. Mark [Bro83, CG85, Wil86a, Wil86b]. Mark-Scaling [Bro83, Wil86a, Wil86b]. Markov [CH87, Leu83a, LW83, Lle87]. Markovian [Pfl84]. marks [Var80]. Martin [Bel82, Mil82a, BK87]. Martin-Löf [BK87]. Marwick [VP80]. Match [Zve80]. Matching [PK85, Coo89c]. Mathematical [LW85c, LM88, JR81]. Matrices [CZ82, SC81, CHPS85]. Matrix [ME88a, CL83b, SE82]. Matter [Hut84]. Matters [Hug89]. Maximum [Wet81]. Mean [Bra85, Bra86b, Deo83]. means [Sow87]. Measure [Pra84]. Measurement [CCD86, RS88, BM84, JC88]. Measurements [Muc82]. Measures [RS87]. Measuring [Lem80, JC86]. Mechanics [Pau80b]. Mechanism [KS82, BP82]. mechanisms [O’S88]. mechanization [Mic82]. median [Erk84]. median-of-three [Erk84]. Medium [GP88, MD84]. Medium-Sized [MD84]. Meek [FJA80, NPCA81, HWW87]. Meet [Mum85]. Meeting [Ano81k]. Meetings [Ano82c, Ano81l]. Memories [Dob84]. Memory [Bur87a, LE86, Pfl84, Ree87, Ver87, Wri83, BBKR89, Dav82]. Merging [DD87, DĎ88a, DD88b, Wri82]. Message [Rob88]. messaging [GGH^{+}87]. Method [CC87, CC88, Fro85, McL85, Sub85, HWW87, Kec87, Lei87, Sow87]. Methodologies [CD81, FSW85, SWB87]. Methodology [BR86, HS83, Jac80a, TAP89]. Methodology-Directed [HS83]. Methods [Che82, Fle82, JR81, Laz81, LCP81, RvS88, Wic89a, EHW89, JC86, Spo85]. Metholodology [Rob80b]. Metric [DW88]. Metrication [FW86]. Metrics [Ben82g, Nav87]. Michael [Bar82b]. micro [MR85]. micro-computer [MR85]. Microcode [Fos86]. Microcomputer [Dea82, EM81a, Ham81b]. Microcomputers [CS85]. Microelectronics [Jam81]. Microprocessor [ES81, Hor83, WW83, MMB87]. Microprocessor-Based [ES81]. Microprocessors [Ben82c, Hea81, Wat82, NQ80]. Microprogram [LCP81]. Millington [Bri82]. MIMD [ZE88]. minicomputer [Ken80]. minicomputers [Hea81, Wat82]. Minimal [CL86, SM88]. Minimaxing [Mic81]. Minimize [Deo83]. Minimum [MF88]. MIS [EDS86]. Missing [Kol83]. ML [AJ89, MA89]. Model [DC84, FM86b, FS81, Goo85b, KA86, Leu83b, Mac81, PW81a, RL89, Ree84b, Ree87, RvS88, Rob88, Zah81, CM89, Dee80, Fun84, Gom81, KGC85, Leu82b, MKP87, Per80, ZR88]. Modeling [Rad82]. Modelling [CH87, LM88, Par82, BWM88, KH87, SC87, ZR88]. Models [Bar83a, Bas86, Dav84b, FRS^{+}87, Fur89, LW83, LW85b, LW85c, Lle83, Min88, OM88, ZE88, Coo89c, Dav84a, HTW89, Ros82]. modern [BHY89]. modest [Tho80a]. Modified [MM84, SE82, SAA88]. MODULA [Wan80, Bai88]. Modula-2 [Bai88]. Modular [Eme80]. Modules [DAEH81, MG88]. Modulo [CC87, CC88]. Monitor [CR83]. Monitoring [PAW86]. Morphic [FD88]. Morris [Eme80]. movements [Nyg80]. Moving [PW81a, Rit89]. MU5 [Eme80, BB82]. Mu6P [WW83]. Multi [CC87, Mit89c, Par88, Ken85]. Multi-destination [Mit89c]. multi-dimensional [Ken85]. Multi-Disc [CC87]. Multi-LAN [Par88]. multidimensional [JC88]. Multimedia [NH89]. Multiple [BHK85, CC87, GP88, Lo84, MMB87, MKP87, NSW87]. Multiplexing [MCL89]. Multiplexor [ES81]. Multiprocessor [BWW81, Bur87a, MCL89]. Multisection [BES83]. Multivalued [Jaj86]. Mutual [HPR88, PJ89]. Myers [FJA80]. Mystical [Sta85].

N [Eme80, FSWL82, Kil81b]. NAND [RDT80]. NAND/NOR [RDT80]. National [Ano81h, Noo84]. Nations [Bog83]. Natural [Pau85, PJ80, FL89]. Nature [Sca89, Tul85]. Necessity [Dun82]. Need [BS89]. Needed [GG86]. Needs [Mum85]. Neglected [Ben82b]. Nested [MT83, SP82]. Nesting [CD85]. Net [Rad82, Gal87]. Network [EM81b, FRS^{+}87, HPR88, IEH^{+}85, MHS85, AO87, CF89, Fun85, Ken80, KH81, PPRB89, PPB84c]. Networks [EM81a, Gil83c, GP88, Har84b, Lea80, SK85, Wel82, Lip80, Sha82b, Eme80]. Neural [FRS^{+}87]. Nomenclature [Wil83c]. Non [Er88b, Goo80, Goo85a, HH80, Wey82, WF89, Van86, Hat84]. Non-Classical [Van86]. Non-Overflow [Goo80]. non-recursive [Hat84]. Non-Regular [Er88b]. Non-Strict [WF89]. Non-Testable [Wey82]. Non-Underflow [Goo80, Goo85a]. Nondeterminism [Bur88]. Nonlinear [Mot85, Vaj83, Abb80, IZP81, Mot80]. Noot [Kil81b]. NOR [RDT80, Sal87]. NOR-gate [Sal87]. Normal [DP88]. Normalization [AC84]. Notation [DBH85, Hul86]. Note [Bae84b, Che86b, CG85, Dav84b, DMS^{+}89, EL89, Er82a, Er87c, FP82, Jam87, Lu89, Mic81, Mum84a, Oxl89, Pac88, PK83, Rey83, Sal87, SAM88, WH87, Wic87, Wil86a, BM89, Dav82, Dav84a, Er83b, MT83, Pro83]. Notes [AB84, Akm86, Bra86a, BS84, BK83a, BKK84, CG85, Fro86, GC87, IH86b, Irv84, Jam89, Lo84, MD84, NPS89, PK83, SEPK89, Tse87b, CC88, Hut84, ME88a, Pal88, RvS88]. Notice [Ano88t]. Notices [Ano86d]. Novel [WL87]. Nucleic [CCL87]. Null [HY81]. Number [Cha87, McI82a, MT83, PW81a, SD82, SP82]. Numbers [Er83a].

O [DD88b, MK82, Per80]. Obituary [Fox87]. OBJ [GCS89]. Object [Coo89a, Cro80, DE86, FA88, Fur89, HW89, LSV89, Coo89c, KL89, Yel89, ZR88]. Object-Based [Fur89]. Object-Oriented [Coo89a, HW89, LSV89, KL89, Yel89, ZR88]. Objects [Hop81, CFL84, DPSW89]. Office [Geh83, Gou80, Tyl88]. Offshore [She85]. One [Lyo85, Com85, CCD86]. One-in [CCD86]. One-in/one-out [CCD86]. one-out [CCD86]. Only [CD82, Sal87]. Open [Bra85, Bra86b]. Operating [BMNS87, Gil83b, Hop81, JS89, MT86, Gil82]. Operations [Abe85, Bla82, GN83, OW83a, OW83b, Rey83, Ull88, Wet81]. Operator [BBK87]. Optically [MCL89]. Optimal [Akl87, BGR86, CD85, HK85b, KW84, Leu82a, Pag82b, Pap82, Sur86, Bra80b, EL82, KR89]. Optimal-Fit [Leu82a]. Optimisation [Den80b, LW85a]. Optimising [Seg87, Thi87, BHY89]. Optimization [HD82, Vaj83, Tal84]. optimum [HK82]. Options [Tri80]. Order [BD81, Er83a, Er87c, Er88b, Irv84, SB85, Er87a, Hen88, Wil83d]. Order- [Er83a]. Ordered [DY80, Er85b, Er89a, BPG87]. ordering [Spo85].

Organisation [Cha81a, HK85b, Sca85, Sur86]. Organisational [Le 88]. organisations [EH88]. Organization [DC84, Dav84b, Dav84a]. Organizations [CN82]. Oriented [Coo89a, HW89, KB87, LSV89, UPT89, Bai88, CL86, KL89, TAP89, Yel89, ZR88]. OUFDM [Lei87]. output [Fro86, HPB84, WH87]. Overflow [Bra85, Bra86b, Goo80, MKP87]. Overlapping [BHK85]. Overlay [SAM88, BKK87].

P [Bur80b, Kil81b, NQ80]. P. [Gil82]. Pack [Fre82]. package [GP80]. Packages [SLH83]. Packing [Ben82a]. Pad [SKW85]. Paper [DMR88]. Papers [Ano84d, Ano85g, Ano85h, Ano80b, Bre81b]. paradigms [KH87]. Parallel [Akl87, BE82, BES83, Clo87, CCL87, DM85, Er83c, EW80, FRS^{+}87, GB83, JJ86, Lea86, Pey89, WI84, BBKR89, CFS89, PGM84]. Parallelism [Bar85]. parameter [Abb80]. Parametric [GM83]. parenthesis [Er83b]. Parity [War86]. Parser [Sto86]. Parsing [SEPK89]. Part [Ree80b, Ree83, Rob80b, Tra84]. Partial [Goo80, NS87, Zve80]. Partial-Match [Zve80]. Participative [Mum84b, Mum84a]. partition [AO80b]. Partitioned [Sch83, Seg87]. Partitioning [FM83]. Partitions [DMS^{+}89, Er88a, FL80]. PASCAL [BKK87, Wic82, AK81, Cai82, DE86, Jan88, KS82, Mil82b, WC85, WC86, NPCA81]. Pascal/1000 [Mil82b]. Password [APW86]. Password-Capability [APW86]. Patches [McL85]. Path [ADD86, DE86, Kie88, DE86]. Paths [HK80]. Pattern [PK85, Tra82, Tra84]. Patterns [LW85b, ZBS81]. PDP [WD80]. PDP-11 [WD80]. Peak [EM81b]. Pedagogical [Gra87]. peg [Lu89]. Perfect [CL86, CHK85]. Performance [Bri85, CC87, CC88, FS81, GN88, IEH^{+}85, IP86, JW82, Lar82, Lar85, Mah84, MK82, Muc82, Mul85, NPCA81, Rey83, Ver87, GE85, JC88, Leu82b, WM84]. Performances [ES81]. Perkins [Mil82a]. Perlog [MG88]. Permutation [Atk80, Er87c, Irv84, MF82, VR81b, Er87a, Erk84]. Permutations [Akl87, GB83, Top82, CF89, Er89a, Spo85, Var80, Wil83d]. Persistence [MG88]. Persistent [ABC^{+}83, DB88, DPSW89]. Personal [Wil84, Ham80a]. Perspective [Gil83b, Gil82]. Pertaining [Wri83]. Peter [FJA80]. Petri [Gal87, Rad82]. Phenomenon [FE80]. philosophy [Cap86]. Physical [BMT85]. Pictorial [DP88]. picture [PW81b]. Pictures [Kil81b]. Pieces [Ben82a, RHS89]. Pipeline [ME88a]. pis [Cr081]. PL [FJA80]. PL/1 [FJA80]. Placement [Bel84, CW82, Dee85, Pag84]. Plagiarism [Jan88]. Planar [Wil85a, WM84]. playing [Bra80b]. Plotters [Ear80, PW81b]. PMS [DBH85]. Point [Bla82, BKK84, EHH86, RS88, Wic89b]. Point-in-Polygon [BKK84]. Pointer [vdW88]. Points [BJ80, Kio81, Sch82a, dS80, HK82]. Policy [Ano84g]. Polish [Rei81]. Polygon [Bra88, BKK84]. Polygonal [Par89]. Polygons [KVW89, Kio81, dS80]. Polyhedral [FA88]. Polymorphic [Cow86, Har84a]. polynomial [Wes80b]. polyphase [EL82]. polytopes [Wat81]. Port [Sil81]. Portability [NQ80]. portable [Wal82]. Porting [Cro80]. Portions [Pag81]. Ports [BPG87]. Possible [Ric88, Hea81, Wat82]. Postage [Mit89a]. postfix [Rei81]. Power [JJ86].

Practical [Ben82f, CHK85, Hut86b, Muc82, Par82, ST82]. Practice [Ben82e, Hor83, EL82, HN88, Tay81]. Precedence [Er82a]. Preci [DNV81, BD82, DAODT85, WS86]. PRECI* [DAT87]. precompilation [Ber87b]. predicting [Per80]. prediction [Wit80]. predictor [Ano80c, Sne80, Wes80a]. predictor-corrector [Ano80c, Sne80, Wes80a]. Premature [Hun81b]. preparation [Smi86, Wal82]. prescribed [CL83b]. presentation [RSS87]. Prevention [TF87]. Pricing [EM81b, GK84]. Primary [HK80]. prime [Iga80]. Primer [Gro82]. Principles [Ben82e, BBKR89, BB81, Clo87, Hut86b, TL85b, FSWL82, JR81]. Priority [BBK87, FK89, GP88, Sil82]. Prize [Ano81o]. Probes [Lyo85]. Problem [BES83, Bar84, BGR86, BR86, Col83, GC87, Hut84, Mit89a, Sel88, TAP89, Wil83c, AO80a, Bai88, Er82b, Er88c, HM80b, KH81, Mot80, NPCA81]. Problem-oriented [TAP89, Bai88]. Problem-Solving [BR86]. Problems [ADD86, Bel84, BWW81, KV84, Mur85, NPS89, SM88, UHS85, AO80b, BKK87, Fun85, PJ89]. Procedural [CD85, HH80]. Procedure [BB80b, Leu82a, ME80, HW88, PJ86]. Procedures [RS88]. Process [BBK87, Ben82d, CZ82, Lav83, TF87]. Processable [EW80]. Processe [Sil82]. processes [Nat86]. Processing [CCL87, Dea82, HM84, Lea86, Lic89, MK82, Wis82, ET88, Ham81b, MMH87, SI80, Sow87, Thi81]. processor [NSW87, TCL89, WN81]. Processors [Bro84b, Lea86, MF82, TL85b, War86, Lay85]. Product [CC88, ME88a]. Production [Lei83, BB89]. Productivity [Laz81]. products [EH88, Tyl88]. Program [Bas86, Bir89, CCD86, CR83, CCS87, Eme80, FW86, FK86, Gol82, Ing80, LB86, McI82b, Mot85, RS87, Tse87a, Tse87b, Woo86, BM80, BKK87, GCS89, Hut86a, MP89, Mot80, Sow87]. programme [Sha82a]. Programmed [Mil82a, Lei87]. Programmer [FR87, GN83]. Programmer-Defined [GN83]. Programming [All83, All85, ABC^{+}83, ABM88, Bar83a, Ben82b, BGR86, Bur88, DE86, EB89, FSWL82, Fis83, Fle82, Gri88, HS83, Hel84, Hil83, Hug89, Jac80b, JS89, Knu84, Lee82, Lei86, LM88, MK85, Mil82b, Mot85, Pag81, Pey89, Ree88, Rob80b, Spi83, Thi86, Tri80, Wad89, Wil89, WD80, WF89, Bai88, BPG87, BM80, Bur87b, Er86b, Lay85, Mot80, Wal82, Wan80, Wil82b, Hil82, NPCA81, NPCA81, VP80]. Programming-Models [Bar83a]. Programs [CGK80, Dun82, Fre82, HW89, Jan88, Lew84, Oul82, Sto86, Thi87, Top82, Wey82, Bra80b, BM84, CFS89, HM80b, SAM88, WC85]. project [SPMD83, GGH^{+}87]. Projections [Ull88]. Prolog [Foo86, KB87, MG88, WC86]. Prolog-Based [WC86]. prolonged [CM89]. Proof [Lew84, BM86, PJ86]. Proofs [Pau85]. Propagation [KS85]. Properties [FD88, Pal88, MP89]. proposal [Coo89b, Hou83, Tho80a]. proposed [Tri83]. propositional [Gib88]. Protection [Cha87, Sho81, Wil84]. Protein [CCL87]. Protocol [BS84, NH89]. Protocols [Pac88, Sha82b]. prototype [DAT87, SS81]. Prototypes [DM83]. Prototyping [MD87, UPT89, Har88]. Proving [MP89]. Public [Mit82]. publications [Smi86]. Publishing [Col89, Dun82, BB89]. Purpose [ASG79, Ing82, Hil80, MR82, Rob80a, Tay80]. Puzzle [RG86]. Puzzles [Smi83].

QMC [Rob88]. Quad [BK83a, Woo82, Woo84, BK83b]. Quadratic [Jam87, ST81]. Quadtree [OW83a, OW83b, Wil88]. Quadtrees [Abe85, FM86a, Ste86, UVS87, BKK87]. Quadtress [AF88]. Quality [RS87]. Quantification [NPS89]. Quantitative [Che82]. Quasi [YW82]. Quasi-Equifrequent [YW82]. Queries [LG86, Pap82]. Query [OM88, ET88, PPRB89, Ser81a, Ser81b, SvR83, Tal84, TL85c]. Queueing [Col83, Lip80]. Queues [Har84b]. Queuing [Sil82]. Quickshunt [McC82]. quicksort [Erk84, Ver87]. Quindiagonal [CZ82, SC81, SE82].

R [BB80a, Bur80b, FSWL82]. R. [Hil82]. Rainbow [SKW85, WSG^{+}84]. Random [EN82, IH86a, IH86b, Mah86, NW87, Ree83, SW84, WN86, Ree80b]. Range [LG86]. Rank [SB85]. Ranking [Er87b, Pal86]. Rapid [IW86]. Rapidly [Jam89]. Raster [Col87, Par89]. Ratio [GOW83]. Raulefs [FJA80]. RCC [NF80]. Re [AW85, WI84]. Re-Circulating [WI84]. Re-examination [AW85]. readable [RB86]. Readers [KRR82]. Reading [DMR88, MP85]. Readings [FSWL82]. Ready [BBK87, Deo83]. Ready-Trace [BBK87]. Real [ER86a, JP86, CL83b, CHPS85, Lut80, Tom82, Ben82b]. Real-Time [JP86, Ben82b]. realisation [WS86]. Recognition [KVW89, Ray80, Tra82, Tra84, ZBS81]. Recognizing [Jaj86]. reconfiguration [BP82]. record [YAH80]. Recording [MTPL88, PAW86]. recordings [Bir81]. Records [Bra85, Bra86b, CN82]. Recovery [ABBS83, BDS82, Har81a, KS85, Mul81, RM89]. Rectangular [Ben82a, EHH86, AO80a]. Recursive [Bla82, BGR86, CC88, Hat84, PJ86]. Reduced [BS89, HK85a]. Reducible [KV84]. Reducing [Bas86]. Reduction [Bur87a, BBKR89]. Redundant [SD82]. refereeing [MP85]. Referencing [Leu83b]. Referential [Bur88]. Refinements [vdW88]. Regions [Par89]. registration [Ano81i, Ano81h, Ano82b]. regression [Gom81, Per80]. Regular [Er88b]. rehashing [Mad80]. Related [BMSA88, Mur85, OW83b, BKK87, FL80]. Relation [Jaj86]. Relational [AK81, BMT85, Bra86c, Bra86a, Bra88, CS85, Fro82, Mac81, Ull88, CAC82, ET88, GE85, Tal84, TL85c, Uck85, WS86]. Relations [Sha81c]. Relationship [Bra88, Par82, DP88]. Relaxation [HM83]. Relevance [Mid87]. Relevant [Tul85]. Reliability [BL81a, Ben82f, ZE88, Fun85]. Reliable [FJA80, FJA80]. Relles [FSWL82]. Remote [FC87, Teu87]. removal [Tom82]. Rendering [Mor89]. Rendezvous [SSDS89]. reorganization [Fun84]. Replacement [Dob84, Jak85]. Replicated [LW85a, MP89]. Replicating [Thi87, Akm86]. Replication [Ben85, Cav83, FD88]. Report [Ano81n, Cho82, Mid87, Tre82]. represent [Ken85]. Representation [Er84c, Fro85, KVW89, PK86, Par89, SW83, Er82b, FL80, RB86].

Representations [SD82]. Representing [BT81]. Requirements [Mum85, NH89]. Research [Ano80e, Har80, Hop80, EDS86, Fro86]. reshuffling [Er88c]. resistance [HN88]. Resolution [Par88, Stu84]. Resource [Lic89, Wan80]. Resources [GG86]. Response [JP86, PAW86, Per80]. Restructuring [WC85]. results [EDS86]. Retained [Hop81]. Retrieval [Har80, Mac81, Zve80, Van86, All82, EHW89, RSS87, SvR83, Ano80e]. Retroactive [Kol83]. Reuse [RL89]. Reve [RG86]. reversal [YY89]. Review [Add81, Ait83, Ale81, And81, Bak80, Bar80, BB80a, Bar83b, Bar82b, Bea81, Bel81b, Bel82, Ben81, Ben82b, Bly80, Bra80a, Bri81a, Bri81b, Bri81c, Bri82, Bro89a, Bur80a, Cha81b, Che82, Cla81, Clu81, Cor81, Day80, De 80, Den80a, Eas80, Edw81, Eva81, Fel80, Fis80, Fre81a, Fre81b, Geo81, Gil82, Gil83a, Gou81, Gra80, Hil82, Hol81, Hop84, Ing81, Jen84, Jon81a, Jon81b, Ken81, Kil81a, Kin81, Lan84, Lav81, Leh81, Lin81, Lov81, Mil82a, MH81, Mux81, Mux84, New80, New81, Pan81, Par83a, Par83b, Pau80a, Pot81, Pro80c, Pro80d, Qui80, Qui81, Rad81a, Rad81b, Ree80a, Ree81, Ril81, Ros81a, Ros81b, Sha81a, Sha81b, Sha82b, Sim81, Sma81, Tho80c, Tho80d, Tho81, Val80, Wal80, Wal81, Wel80b, Wel84a, Wel84b, Wil83a, Wil83b, Wis80, Wis81, Wis82, Woo81, Bur80b, Den80b, JC86, Kil81b]. review [Pau80b]. reviewed [Ano81d, Ano81e, Ano81f, Ano81g]. Reviews [Add80, For80, Jac80d, BL81a, CF81, Eme80, FSWL82, FE80, FJA80, JR81, NQ80, NPCA81, VP80, WD80]. Revisited [Bar84, DĎ88a, OW80]. Richard [Eme80]. Right [Vel82]. Right-to-Left [Vel82]. Rina [WD80]. Rings [Pac88]. ripple [Sal87]. Robert [BL81a, FE80]. role [EH88, MV88, SC87]. Rolland [Eme80]. ROMs [Cla88]. Root [JBJ84, PK83, Pro83, Wes80b]. Roots [Jam89]. rotating [Mar82]. Rotation [Pal88, vdW88, AO87]. rotations [Er89b]. Routines [Dix85, RD80, Teu87]. Routing [SK85, VT87]. RSA [Hun85, Wic87]. rule [BK87, MV88, PJ86, Mic82]. rule-based [MV88, Mic82]. Rules [HY81, HK80, Hou86, Lei83]. Run [PG82]. Rushby [FSWL82].

S [BL81a, NPCA81]. Safe [DB88, Coo89b]. Samhita [PSD^{+}81]. Sampling [EN82, Jak85]. Saving [RDT80, Zak84]. scale [Har88]. Scaling [Bro83, CG85, Wil86a, Wil86b]. Scan [Col87, Ste86]. Scanning [QY89, PW81b]. Scattered [Sch82a]. Scheduling [Deo83, Mit82]. Schema [FL85, Fro83, Sha81c, Dee80, Uck85]. SCHEMAL [Fro83]. Schemas [Jaj86, Oul87]. Schemata [PG81, SGW82, Wil82a, Wil83c]. Schematics [MTW88]. Scheme [CL86, Cha87, Jam87, LW83, Pag84, Pik81, BM89, BP81, Nat86]. Schemes [CCD86, GP88, RL82, Mäk89]. Science [Che86a, Che86b, FE80, GG86, Cap86, GGH^{+}87, BL81a, FJA80]. Scientific [Sta85, Smi86]. Scope [Hou86]. Screen [DMR88, Dix85, Dix85]. seamed [ST81]. Search [Bra85, Bra86b, FP82, Les83b, Mah86, MKB87, AO80b, CM89]. Searches [Lyo85]. Searching [CCL87, MKH86, SPC89, Sur86]. Second [Ano83c, BD81, RD80]. Secondary [BMT85, HK85b, SB85, Aki88]. secure [Mit89c]. Security [Bak82a, Mar80, Tho80b, Wil84]. Seek [Bas86]. Segments [Pag82b]. Segregation [WR86]. Selected [FSWL82]. Selection [AB84, BMT85, Dob84, HK85a, HK85b, Kie88]. Selective [PBH86]. Self [Akm86, Thi87, WL87, Hut86a]. self-augmenting [Hut86a]. Self-Replicating [Thi87, Akm86]. Self-test [WL87]. Semantic [ABD84, Fro85, Lo84, Pag82a, BD80, FW83]. Semantically [SEPK89]. Semantics [All83, BBK87, MA89, NPS89, Tak86, All85, Ser81a, Ser81b, Yel89]. Semaphores [FK89, KRR82]. Semi [LW83]. Semi-Markov [LW83]. Seminars [Ano81i]. Sensitive [AMN81]. Separated [PW81a]. Sequence [CCL87]. Sequential [KW84, MKH86, Zve80]. serial [Sow87]. Series [IS81]. Server [MHS85]. servers [Lip80]. Services [GK84]. Set [Bae84b, DAEH81, DMS^{+}89, Er88a, HK85a, HD82, KB87, Ree84a, WW82, Atk80, Bae84a, HTW89, O’K85]. Set-Oriented [KB87]. set-theoretic [HTW89]. Sets [BS89, Gri80]. SGML [Smi86]. shaded [PW81b]. shaded-picture [PW81b]. Shared [Bur87a]. Shared-Memory [Bur87a]. sharing [Lip80]. Shave [Bar82b]. Short [AB84, Akm86, Bae84b, Bra86a, BS84, BK83a, BKK84, CC88, Che86b, CG85, Dav84b, DMS^{+}89, EL89, Er87c, FP82, Fro86, GC87, Hut84, IH86b, Irv84, Jam87, Jam89, Lo84, Lu89, MD84, MT83, ME88a, Mum84a, NPS89, Oxl89, Pac88, Pal88, PK83, RvS88, Rey83, Sal87, SEPK89, SAM88, Tse87b, Wil86a]. Shortcut [PG82]. sided [BJ80]. significance [Tri83]. Simple [Bar85, Er88b, HM83, NS87, Sub85, Kec87, Leu82b]. Simplified [Bar82a]. Simplifying [Dix85]. SIMULA [Hel84, PGM84, IP86]. Simulation [CD81, DBH85, ER80, Gil83c, Sch82b, UPT89, Gom81, Har88, Har82b, PRO80a, SB81]. simulation/regression [Gom81]. Simulations [Hen88, CM89]. simultaneous [IZP81]. Singh [Den80b]. Single [HWW87, Lar85, Tse87b, Wri82]. Single-File [Lar85]. Singular [NS87]. Sint [Kil81b]. Situ [SW84]. Size [SPC89, Iga80]. Sized [MD84, Pag82b]. skeleton [Lei87]. Skies’ [GGH^{+}87]. Slices [Skl81]. slots [Hut89]. Slotted [Pac88]. Small [Fre82, Lic89, Kec87, Mul87]. Smalltalk [HW89]. Smalltalk-80 [HW89]. smooth [Er88c]. Smoothing [MZ86]. Social [Mum80, Nyg80]. Society [Bel82]. Soft [WW82]. Soft-edged [WW82]. Software [BB80a, Ben82g, Che86a, Che86b, Col80, Dea82, FW86, FJA80, GW88, Hop80, IW86, JK80, Laz81, LE86, LB86, MP85, Pra84, RL89, SWB87, ST89, Fal81, MBC^{+}86, BL81a, NQ80]. solid [HTW89]. Solution [BE82, BGR86, DAEH81, Lu89, Sch85, Sel88, Abb80, AO80a, PJ89]. Solutions [BD81, Bla82]. Solve [LG86]. Solving [BR86, BWW81, Col83, Jam87, NPCA81, UHS85, Er88c, Fun85, IZP81, KH81, RD80]. Some [Abe85, All88, ASG79, FR80, HK80, Hut86b, Ing82, Les83b, Lic89, Mid87, Pal88, Roh87, TL85b, VR81b, Wri83, EDS86, Fun85, Hil80, MR82, Rob80a, Sha82a, Tay80]. sort [EL82]. Sorter [WI84]. Sorting [CMS80, Dro87, Er83c, KV84, McC82, SW84, WI84, SAA88, TH82, VR81a]. sound [Bir81]. Source [Lle87]. Space [BD82, Col87, DW88, DD87, DD88b, FA88, Har81a, MCL89, KR89, Zak84]. Space-Based [DW88]. Space-Filling [Col87]. space-optimal [KR89]. sparsely [Dav82]. Spatial [BWW81, Bra86c, Bra86a, Mil81b, RB86]. Speaker [Ray80]. special [Mot80]. Specialisation [Clo86]. Specification [Dro87, DF87, Ing80, JK80, Mai85, Rob88, Sil82, Wic89b, Lei87, Pag82a, Shi89, TL85c, WS86]. Specifications [CCS87, Dix85, Les83a, Wil85b, FM85, GCS89]. Specifying [PPB84c]. Specimen [Cho82]. Speech [Flo80, Ray80]. speeding [Sow87]. Spiral [Mul85]. Spline [BT81]. Splines [MZ86]. Split [Sel88]. Splitting [PBH86]. Spss [Eva83]. SQL [NPS89]. Square [Jam89, PK83, Pro83]. Squares [McI82b]. stability [Car86]. Stable [DD87]. Stack [Goo80, Goo85a, Goo85b]. Stack-Length [Goo85b]. Stamp [Mit89a]. Standard [Clo86, Mil81a, VP80, BP81, MA89, Tri83]. Standardisation [NPCA81]. Standards [Bro89b]. State [Atk84, Lle83, Ree84b]. Statement [Ano84g]. statements [Pag82a, WC85]. States [McI82a]. Static [KS82, Mah84]. Statistical [FSWL82, Sch83, SLH83, CF81]. Statistics [Fre82]. STATLILB [FSWL82]. Steiner [Col82]. step [FW83]. Steps [Sha81c, Yel89]. stock [AO80a]. Storage [Bas86, Fro82, Har81a, Har82a, Leu82a, McI82a, Mul85, Ree84b, Sam81, Ste86, Aki88, All82, Gom81, Leu82b, TH82, Wil80]. Store [Ree83, Ree80b]. Storing [AF88]. straight [CP87]. Strategies [AW85, Cha81a, CD82, Bra80b, Noo84]. strategy [BWM88]. Stream [BS89]. Strict [WF89]. String [Lea86]. strings [Er83b]. strongly [DB88]. Structural [CGK80, Car86, CP87]. Structure [BK83a, Fur89, MT83, Woo82, Wri83, BK83b, Hut86a, PJ86, SP82, Stu84]. structure-directed [PJ86]. Structured [ADD86, Bro89b, Eme80, Fis83, Gol82, Hou86, Ing80, Jac80b, Jac80c, Lei83, Mil81a, MTW88, NPCA81, Tri80, BB80b, Lem80, ME80, Oul87]. Structures [BIY88, BHK85, Fro82, FK86, Gol85, IS81, SKW85, Sur86, Bir81, She85, Bar82b]. Structuring [ER86a, FM86b, Ran86, Wil85b]. Student [Jan88, Lem80]. Students [Mil82a]. Studies [GW88, Coo83]. Study [Goo85b, IP86, Laz81, Lle83, Mum85, Aki88, Lip80, Sha82a]. Style [Lee82, Pag81]. Subject [Sta85]. Sublinear [DD87]. subroutine [CL83b, CHPS85, IZP81]. Subset [SM88, Var80]. Subset-Generating [SM88]. Substitution [HM83]. subsystem [Bro82]. Subsystems [Kie88]. Subtypes [BL81b]. Success [Lic89, EDS86]. Suitable [CCL87, CF89]. Superposed [FS87]. supervision [Wan80]. Supervisor [Cav83]. Supplement [Ano81a, Ano84a, Ano84b, Ano87a, CL83a, Wil83e, Ano80a, Ano81b, Ano82a]. supplier [EH88]. Support [ST89]. Surface [FR80, Oxl85, Tom82]. Surfaces [McL85, NW87]. Survey [LCP81, Mur85, Nav87, Pau85, Sco82, Mur83]. Switching [Ram85]. Symbols [Mil81a, KGC85]. Symmetric [SC81, ZOG85, CL83b, CHPS85, RM89, SE82]. Symposium [Ano80e, Har80]. Synchronisation [CF88, Nat86, RM89]. Synchronizing [KRR82]. synchronous [MP89]. synonym [GF87]. Syntactic [Lo84, Pag81]. syntax [Ser81a, Ser81b]. Synthesis [CGK80, CD80]. synthetic [GP81]. System [AB84, APW86, BT81, Bas86, BMNS87, CD81, Cav83, DM83, DAODT85, DM85, Fro85, Gil82, Gil83b, HH80, Hop81, HS82, Ing80, JP86, Laz81, LE86, Mac81, MM84, McI82a, MT86, Mum85, Nav87, Par82, Par88, PAW86, PSD^{+}81, Rai87, Ran86, Rob88, Teu87, UPT89, WC86, BWM88, BM86, Cra86, DNV81, DTIR88, Fal81, Gom81, MBC^{+}86, MMB87, MP85, Per80, RSS87, RB85, Sha82a, Shi89, Sma88, SvR83, Eme80]. Systematic [ST82]. SYSTEMATICS [Ser81a, Ser81b]. Systems [Abe85, BB80a, BE82, Bar84, Bar85, BS87, Bow83, CC87, Col82, Coo89a, Dav81, Dea82, DM83, DBH85, FE80, FSW85, Geh83, GM83, Hai83, Ham89, HM84, Hul86, Hut86b, IP86, Jac80c, JS89, Lea80, Lom88, Mah84, MCL89, Mil81a, MK82, Mum84a, NF80, NH89, PB86, PBH86, Pot82, Rad82, Sch82b, Seg87, Sip88, TB82, TF87, VS87, Wil80, WHF82, WHF83, Zah81, ZE88, AFWH88, BM81, Ber87b, BP82, Coo83, ET88, Fro86, Gea80, Ham81b, HM80a, HN88, IZP81, JC86, Kec87, Lip80, MV88, Mar82, Mic80, Mum80, Mum84b, Ros82, SI80, Spu88, Tom82, Wat80, WN81, Bri82, Ham80a, LM88, BB80a, CF81, JR81, Den80b]. Systems/Virtual [Wil80]. Systolic [ME88a, ME88b, WI84].

T [BB80a, Kil81b]. Table [Lew84, Lyo85, Sam81, Fro81]. Tables [Sch85, Mad80, MV88]. Target [FL85]. task [BM80]. tasks [BM86]. Tattoo [MR85]. Taxonomic [Coo83]. Taxonomy [Fur89, WHF82]. TC8 [VS87]. Teach [HS82]. Teaching [WHF83, Cap86, GGH^{+}87, Spu88]. Tech [Lic89]. technical [KC88]. Technique [BB80a, Dee85, Mul81, CP87, RB86]. Techniques [Col87, DF87, ES81, HK88, Jac80c, JW82, MZ86, NPCA81, Sha82b, Woo86]. Technology [Bog83, Hop80, Cap86, Cla88, Ham80b, Mor89]. telecommunications [DTIR88]. Telematic [Bel82]. Teleprocessing [BB80a]. template [Coo89c, HW88]. Temporal [Bar87, Lom88]. ten [Kil81b]. Tennent [FSWL82]. Terminal [PKK87]. Termination [Ree88, Er86b, Hun81b]. tessellation [Wat81]. tessellations [Bow81]. Test [DW88, Inc87, RS88, WL87]. Test-Data [DW88]. Testable [Wey82]. Testing [IH86a, IH86b, WL87, Wey82]. Tests [BKK84]. Text [BS86, BHK85, Pik81, KR89, Thi81]. texts [KC88]. Textual [SPC89, Mul87]. Their [DM83, Sel88, Jam89, MA89, Ros82]. Themes [VS87]. theoretic [CCS87, HTW89]. Theories [Pra88, Tul85]. Theory [BK87, Cha87, Col83, Hor83, Lan85, Mic81, Pra84, Rad82, Rze85, Sta85, BWM88, CAC82, EL82, HN88, CF81]. theory-based [BWM88]. thetas [Cr081]. Threaded [BF81]. three [Erk84]. throughs [Lem80]. Time [Deo83, DD87, DD88b, ER86a, HY81, Jak85, JP86, Lip80, Lut80, Per80, Tom82, YY89, Ben82b]. Times [Deo83, Har84b, JP86, PAW86]. Timetable [SS80, Sel88]. Titli [Den80b]. Tolerance [Ben82e]. Tool [BS87, BS86, AFWH88, BWM88]. Tools [CD81, Hai83, IW86, Jac80c, PB86, GGS89, HKPW89]. Top [ABD84, HW88]. Top-Down [ABD84, HW88]. Topics [Ben82b]. Topological [Er83c, KV84, VR81a]. topology [BP82]. total [PJ86]. Tower [Er82b]. Towers [Er84a, Er84c, Er84d, GC87, Lu89, Roh87]. Trace [BBK87]. Traces [Foo86]. Tracking [Ear80]. Tracks [PW81a]. training [GGH^{+}87]. Trains [CMS80]. Transaction [MK82]. Transaction-Processing [MK82]. transcendental [JBJ84]. transferable [HWW87]. transform [Pie82]. Transformation [DY80, HK88, Oul87]. Transformations [BT81, Hen88, Mid87, YY89]. Transitions [Bro84a]. Translating [WC86]. Translation [Hil83, OM88, Hil82]. Transmission [Fen87]. Transparency [Bur88]. TRANSPLINE [BT81]. Transport [Mit82, NH89]. transputer [AO87]. Trap [Bra88]. Traversal [Bri85, vdW88, Hat84]. traversals [Er89a]. treatment [DPSW89]. Tree [BK83a, BHK85, Har84b, JJ86, SW83, Sur86, TB86, Woo82, ZOG85, vdW88, AO80b, BK83b, FL80, Hat84]. Tree-Like [Har84b]. Trees [BD84, BF81, Bri85, DY80, Er85b, Er87b, Er88b, EA83, FP82, GOW83, LG86, Mah86, Mul81, Pal86, Pal88, Sip88, Woo84, CM89, Er89a, Er89b, OW80, Zak84]. Trends [VS87]. Triangulation [Oxl85]. Triodic [Gol82]. Triple [Col82]. Trotter [VR81b]. Tuple [Dee85]. Tuples [BMSA88]. Turchin [FE80]. Tutorial [Lan82]. Two [DF87, EB89, ES81, JW82, KV84, Lyo85, Pac88, PW81a, PB86, Seg87, Wil85b, Mäk89, Pag82a, PRO80a, SPMD83]. Two-Level [Wil85b, EB89, Pag82a]. Two-Way [Seg87]. Type [ABM88, Cow86, Hen88, Les83a, Sho81, Coo89b, Hou83, Mor89, WS86]. type-safe [Coo89b]. typed [DB88]. Types [BK87, BL81b, Har84a, Mai85, YP86]. Typesetting [CW82].

UFL [YP86]. Ullman [JR81]. UMIST [GCS89]. Unbiased [EN82]. uncertain [Qui83]. Uncertainty [AB84]. Unconditional [Wig81b, Tho80a]. undefined [Cze88]. Underflow [Goo80, Goo85a]. Underlying [Rob88]. Understanding [Cap80, Cla88]. Unification [ME88b]. Unifying [Lo84]. union [Atk80]. Universal [NPS89, Zah81]. University [GG86, Muc82]. UNIX [MBC^{+}86, Hek89]. UNIX-based [MBC^{+}86]. Unranking [Pal86]. Unravelling [Oul82]. Unstable [DD88b]. Unstructured [CCD86, Oul82]. Unstructuredness [Tse87a]. Unstructuredness* [Tse87b]. updates [CM89]. Updating [JW82]. Upon [Cha87, YY89]. Usability [Sha86, Tyl88]. Usage [EM81b, LE86]. Use [Bra85, Bra86b, Gol82, Sur86, Ver87, Er86b, Gri80, Har88]. Useful [AC84]. User [Gil82, Gil83b, Har84a, NF80, PJ80, SLH83, HN88, JC86, Jam80, O’S88, RB85, Bak82b]. User-Defined [Har84a]. User-Extensible [NF80]. User-Friendly [SLH83]. Using [BT81, Ben82c, BGR86, BHK85, Col87, CH87, DE86, Fro85, IP86, JL80, Leu83a, LW85b, LG86, Lom88, MZ86, Mil81a, Pik81, Ree80c, Sch82b, Sho81, SW83, Thi86, UVS87, Wel80a, Wil89, ZBS81, vdW88, Er89b, HW89, KC88, ME88a, Mul87, Pag82a, PRO80a, RvS88]. utilizing [All82].

V [Eme80, FE80]. Validating [Fos86]. Validation [FS81, LCP81]. Value [NS87]. Values [Sho81]. Variable [Lle85, MT83, RM89, SP82]. Variable-length [Lle85]. variables [Cze88]. Variant [Thi86]. Various [Fro85]. VDM [Gib88]. Vector [Ben85, MF82, Par89]. Veen [Kil81b]. Vegas [EL89]. Verification [Gra87, LCP81]. Version [Lar85]. Versions [Roh87]. Versus [BD84, Wri82, DMR88]. Vertex [Sel88]. Vertices [Sel88]. Very [Tre82]. via [Ull88]. view [Mor89]. Viewdata [MM84]. Virtual [Cro80, Ver87, Wil80, Gom81]. Visibility [FA88]. VLSI [CF89, Lea86, Tre82]. Vol.1 [Wis82]. Volume [Bur80b]. Volumes [Oxl89, Rit89]. Voronoı [Wat81]. vote [HWW87]. Vs [MKH86]. VULCAN [MM84].

W [CF81, FE80, JR81, Kil81b, Mil82a, NPCA81]. W. [FSWL82]. walk [Lem80]. walk-throughs [Lem80]. Wasserman [VP80]. Waveguides [MCL89]. Way [Bar87, Seg87, Ken85]. Web [Thi86]. Weber [VP80]. Weighted [Ric88, Sch85, Wil86a, Wil86b]. Welcome [Gea80]. Well [CCD86, Er83b]. Well-formed [CCD86, Er83b]. Westreich [Bre81b]. which [PJ89]. While [BK87, Tak86]. While-rule [BK87]. Wilkes [Ano81o, Ano85m, HT81]. Wilkinson [Fox87]. William [Eme80]. Window [Hek89, PKK87, Mic82]. Winston [Bur80b]. Within [Cha81a]. without [Jak85]. Wombat [Cla88]. Word [McI82b, Thi81]. Workshop [Tre82, FJA80]. Workstation [SKW85, WSG^{+}84]. Worst [Lar82, Ric88, Erk84]. Worst-Case [Lar82]. Worst-Possible [Ric88]. Writers [KRR82]. Writing [Akm86, Cap80, HW89, Pag81]. written [BP81]. WSI [Lea86].

XENO [Smi83].

Yarmish [WD80]. years [SPMD83]. Yen [Pau80b].

zeros [Cr081]. Zipf [Pra88]. zones [HTW89].

References

Agmon:1984:SND

[AB84] T. Agmon and I. Borovits. Short notes: Dynamic and uncertainty considerations in the selection and the acquisition of an information system: an adoption of a break-even analysis to investment decisions. The Computer Journal, 27(1):87–89, February 1984. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

           URL http://comjnl.oxfordjournals.org/content/27/1/87.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/88.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/89.tif.

Abbott:1980:ACS

[Abb80] J. P. Abbott. Algorithm 110: Computing solution arcs of nonlinear equations with a parameter. The Computer Journal, 23 (1):85–89, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_23/Issue_01/tiff/85.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/86.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/88.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/89.tif.

Anderson:1983:ALL

[ABBS83] S. O. Anderson, R. C. Backhouse, E. H. Bugge, and C. P. Stirling. An assessment of locally least-cost error recovery. The Computer Journal, 26(1):15–24, February 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/26/1/15.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/15.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/16.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/17.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/19.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/20.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/21.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
22.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/23.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/24.tif.

Atkinson:1983:APP

[ABC^{+}83] M. P. Atkinson, P. J. Bailey, K. J. Chisholm, P. W. Cockshott, and R. Morrison. An approach to persistent programming. The Computer Journal, 26(4):360–365, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/360.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/360.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/361.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/362.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/363.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/364.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/365.

tif. Also published in/as: In “Readings in Object-Oriented Database Systems” edited by S. Zdonik and D. Maier, Morgan Kaufman, 1990.

Adorni:1984:TSA

[ABD84] G. Adorni, A. Boccalatte, and M. Di Manzo. Top-down semantic analysis. The Computer Journal, 27(3):233–237, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/233.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/233.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/234.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/235.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/236.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/237.tif.

Abel:1985:SEO

[Abe85] D. J. Abel. Some elemental operations on linear quadtrees for geographic information systems. The Computer Journal, 28(1):73–77, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/1/73.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/76.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/77.tif.

Atkinson:1988:BTC

[ABM88] M. Atkinson, P. Buneman, and R. Morrison. Binding and type checking in database programming languages. The Computer Journal, 31(2):99–109, April 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/31/2/99.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_31/Issue_02/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_31/Issue_02/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_31/
Issue_02/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/103.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/105.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/106.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/107.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/108.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/109.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/99.tif.

Arora:1984:NCU

[AC84] Adarsh K. Arora and C. R. Carlson. Normalization could be useful. The Computer Journal, 27(1):57–61, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/27/1/57.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/57.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/58.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/59.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/60.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/61.
tif.

Addyman:1980:BR

[Add80] A. M. Addyman. Book reviews. The Computer Journal, 23(2): 146, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/146.3.full.pdf+html.

Addyman:1981:BR

[Add81] A. M. Addyman. Book review. The Computer Journal, 24(4): 384, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/384.4.full.pdf+html.

Ancona:1986:PPS

[ADD86] M. Ancona, L. De Floriani, and J. S. Deogun. Path problems in structured graphs. The Computer Journal, 29(6):553–563, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/6/553.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/553.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/554.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/555.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/556.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/557.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/558.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/559.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/560.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/561.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/562.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/563.
tif.

Anedda:1988:CQI

[AF88] C. Anedda and L. Felician. P -compressed quadtress for image storing. The Computer Journal, 31(4):353–357, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/4/353.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/353.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/354.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/355.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/356.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/357.tif.

Avison:1988:ISD

[AFWH88] D. E. Avison, G. Fitzgerald, and A. T. Wood-Harper. Information systems development: a tool kit is not enough. The Computer Journal, 31(4):379–380, August 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/4/379.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/379.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/380.tif.

Aitchison:1983:BR

[Ait83] I. Aitchison. Book review. The Computer Journal, 26(2):190– 191, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/190.3.full.pdf+html.

Augustsson:1989:CLM

[AJ89] L. Augustsson and T. Johnsson. The Chalmers Lazy-ML compiler. The Computer Journal, 32(2):127–141, April 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/2/127.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/127.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/128.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/129.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/130.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/131.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/132.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/133.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/134.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/135.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/136.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/137.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/138.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/139.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/140.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/141.
tif.

Alagic:1981:RPD

[AK81] S. Alagić and A. Kulenović. Relational Pascal data base interface. The Computer Journal, 24(2):112–117, May 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/112.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/112.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/113.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/114.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/115.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/116.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/117.tif.

Akinyokun:1988:ESC

[Aki88] O. C. Akinyokun. The experimental study of CODASYL database/administration at the secondary storage level. The Computer Journal, 31(4):344–352, August 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/4/344.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/344.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/345.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/346.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/347.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/348.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/350.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/352.tif.

Akl:1987:AOP

[Akl87] S. G. Akl. Adaptive and optimal parallel algorithms for enumerating permutations and combinations. The Computer Journal, 30(5):433–436, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/5/433.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/433.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_05/tiff/434.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/435.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/436.tif.

Akman:1986:SNW

[Akm86] V. Akman. Short notes: Writing self-replicating code. The Computer Journal, 29(6):573–574, December 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/6/573.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/573.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/574.tif.

Aleksander:1981:BR

[Ale81] I. Aleksander. Book review. The Computer Journal, 24(2):106, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/106.1.full.pdf+html.

Allen:1982:FID

[All82] F. W. Allen. A file index for document storage and retrieval utilizing descriptor fragments. The Computer Journal, 25(1):2–6, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/25/1/2.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/2.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/3.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/4.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_25/Issue_01/tiff/5.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_25/
Issue_01/tiff/6.tif.

Allison:1983:PDS

[All83] Lloyd Allison. Programming denotational semantics. The Computer Journal, 26(2):164–174, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/2/164.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/164.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/165.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/

tiff/166.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/167.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/168.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/169.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/170.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/171.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/172.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/173.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/174.tif.

Allison:1985:PDS

[All85] Lloyd Allison. Programming denotational semantics. II. The Computer Journal, 28(5):480–486, November 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/5/480.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/480.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/481.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/482.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/483.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/484.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/485.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/486.tif.

Allison:1988:SAC

[All88] L. Allison. Some applications of continuations. The Computer Journal, 31(1):9–11, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/31/1/9.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_
01/tiff/10.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/9.tif.

Atkinson:1981:CSE

[AMN81] L. V. Atkinson, J. J. McGregor, and S. D. North. Context sensitive editing as an approach to incremental compilation. The Computer Journal, 24(3):222–229, August 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/3/222.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/222.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/223.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/224.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/225.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/226.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/227.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/228.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/229.tif.

Anderson:1981:BR

[And81] J. Anderson. Book review. The Computer Journal, 24(4):342, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/342.1.full.pdf+html.

Anonymous:1980:AS

[Ano80a] Anonymous. Algorithms supplement. The Computer Journal, 23 (1):85–89, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/1/85.full.pdf+html.

Anonymous:1980:CP

[Ano80b] Anonymous. Calls for papers. The Computer Journal, 23(1):33, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/33.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/33.tif.

Anonymous:1980:EEP

[Ano80c] Anonymous. Erratum: “An efficient predictor-corrector algorithm”. The Computer Journal, 23(4):383, November 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/383.tif. See [Wes80a, Sne80].

Anonymous:1980:ICD

[Ano80d] Anonymous. International conference on data bases. The Computer Journal, 23(1):36, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/1/36.full.pdf+html.

Anonymous:1980:JBA

[Ano80e] Anonymous. Joint BCS ACM Symposium on Research and Development in Information Retrieval. The Computer Journal, 23 (2):131, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/131.full.pdf+html.

Anonymous:1981:ASa

[Ano81a] Anonymous. Algorithms supplement. The Computer Journal, 24 (1):87–90, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/1/87.full.pdf+html.

Anonymous:1981:ASb

[Ano81b] Anonymous. Algorithms supplement. The Computer Journal, 24(4):383–384, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/24/4/383.full.pdf+html.

Anonymous:1981:A

[Ano81c] Anonymous. Announcement. The Computer Journal, 24(2):192, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/192.3.full.pdf+html.

Anonymous:1981:BRIa

[Ano81d] Anonymous. Books reviewed in this issue. The Computer Journal, 24(1):8, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/1/8.full.pdf+html.

Anonymous:1981:BRIb

[Ano81e] Anonymous. Books reviewed in this issue. The Computer Journal, 24(2):138, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/138.full.pdf+html.

Anonymous:1981:BRIc

[Ano81f] Anonymous. Books reviewed in this issue. The Computer Journal, 24(3):255, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/255.full.pdf+html.

Anonymous:1981:BRId

[Ano81g] Anonymous. Books reviewed in this issue. The Computer Journal, 24(4):382, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/382.full.pdf+html.

Anonymous:1981:CRF

[Ano81h] Anonymous. Call for registration: First British national conference on databases. The Computer Journal, 24(1):16, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/1/16.full.pdf+html.

Anonymous:1981:CRC

[Ano81i] Anonymous. Calls for registration: The Computer Journal Seminars. The Computer Journal, 24(2):111, ???? 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/24/2/111.
full.pdf+html.

Anonymous:1981:C

[Ano81j] Anonymous. Corrigenda. The Computer Journal, 24(2):155, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/155.full.pdf+html.

Anonymous:1981:FMa

[Ano81k] Anonymous. Forthcoming meeting. The Computer Journal, 24 (3):270, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/270.4.full.pdf+html.

Anonymous:1981:FMb

[Ano81l] Anonymous. Forthcoming meetings. The Computer Journal, 24 (4):373, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/373.full.pdf+html.

Anonymous:1981:IA

[Ano81m] Anonymous. Information for authors. The Computer Journal, 24 (1):2, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/2.tif.

Anonymous:1981:R

[Ano81n] Anonymous. Report. The Computer Journal, 24(4):377, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/4/377.5.full.pdf+html.

Anonymous:1981:WP

[Ano81o] Anonymous. The Wilkes Prize. The Computer Journal, 24(1): 1, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/1.tif.

Anonymous:1982:AS

[Ano82a] Anonymous. Algorithms supplement. The Computer Journal, 25(2):278–282, ???? 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/2/278.full.pdf+html.

Anonymous:1982:CR

[Ano82b] Anonymous. Call for registration. The Computer Journal, 25 (2):286, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/286.tif.

Anonymous:1982:FM

[Ano82c] Anonymous. Forthcoming meetings. The Computer Journal, 25(4): 485, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/485.full.pdf+html.

Anonymous:1982:IA

[Ano82d] Anonymous. Information for authors. The Computer Journal, 25 (2):288, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/288.tif.

Anonymous:1983:A

[Ano83a] Anonymous. Announcements. The Computer Journal, 26(4):384a– 384, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/384-a.full.pdf+html.

Anonymous:1983:C

[Ano83b] Anonymous. Conferences. The Computer Journal, 26(2):192, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/2/192.2.full.pdf+html.

Anonymous:1983:SIC

[Ano83c] Anonymous. Second international conference on databases. The Computer Journal, 26(2):192, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/2/192.1.full.pdf+html.

Anonymous:1984:ASa

[Ano84a] Anonymous. Algorithms supplement. The Computer Journal, 27(2):178–187, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/27/2/178.full.pdf+html.

Anonymous:1984:ASb

[Ano84b] Anonymous. Algorithms supplement. The Computer Journal, 27(4):368–372, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/27/4/368.full.pdf+html.

Anonymous:1984:A

[Ano84c] Anonymous. Announcement. The Computer Journal, 27(3):288, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/3/288.3.full.pdf+html.

Anonymous:1984:CP

[Ano84d] Anonymous. Call for papers. The Computer Journal, 27(3):288, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/3/288.2.full.pdf+html.

Anonymous:1984:C

[Ano84e] Anonymous. Corrigendum. The Computer Journal, 27(3):288, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/3/288.4.full.pdf+html.

Anonymous:1984:FE

[Ano84f] Anonymous. Forthcoming events. The Computer Journal, 27(3): 288, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/288.1.full.pdf+html.

Anonymous:1984:SP

[Ano84g] Anonymous. Statement of policy. The Computer Journal, 27(2): 187, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/2/187.full.pdf+html.

Anonymous:1985:Aa

[Ano85a] Anonymous. Announcement. The Computer Journal, 28(1):95, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/1/95.2.full.pdf+html.

Anonymous:1985:Af

[Ano85b] Anonymous. Announcement. The Computer Journal, 28(5):544, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/5/544.full.pdf+html.

Anonymous:1985:Ab

[Ano85c] Anonymous. Announcements. The Computer Journal, 28(2):104, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/2/104.1.full.pdf+html.

Anonymous:1985:Ac

[Ano85d] Anonymous. Announcements. The Computer Journal, 28(4):446, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/4/446.2.full.pdf+html.

Anonymous:1985:Ad

[Ano85e] Anonymous. Announcements. The Computer Journal, 28(5):473, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/5/473.2.full.pdf+html.

Anonymous:1985:Ae

[Ano85f] Anonymous. Announcements. The Computer Journal, 28(5):529, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/5/529.full.pdf+html.

Anonymous:1985:CPa

[Ano85g] Anonymous. Call for papers. The Computer Journal, 28(2):104, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/2/104.3.full.pdf+html.

Anonymous:1985:CPb

[Ano85h] Anonymous. Call for papers. The Computer Journal, 28(4):446, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/4/446.3.full.pdf+html.

Anonymous:1985:Ca

[Ano85i] Anonymous. Conference. The Computer Journal, 28(2):104, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/2/104.2.full.pdf+html.

Anonymous:1985:Cb

[Ano85j] Anonymous. Conference. The Computer Journal, 28(5):473, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/5/473.1.full.pdf+html.

Anonymous:1985:Cc

[Ano85k] Anonymous. Conference. The Computer Journal, 28(5):517, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/5/517.full.pdf+html.

Anonymous:1985:FE

[Ano85l] Anonymous. Forthcoming events. The Computer Journal, 28(4): 446, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/446.4.full.pdf+html.

Anonymous:1985:WA

[Ano85m] Anonymous. The Wilkes Award 1984. The Computer Journal, 28 (5):507, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/507.full.pdf+html.

Anonymous:1986:Ab

[Ano86a] Anonymous. Announcement. The Computer Journal, 29(4):347, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/4/347.full.pdf+html.

Anonymous:1986:Aa

[Ano86b] Anonymous. Announcements. The Computer Journal, 29(3):284, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/3/284.full.pdf+html.

Anonymous:1986:Ac

[Ano86c] Anonymous. Announcements. The Computer Journal, 29(4):360, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/4/360.full.pdf+html.

Anonymous:1986:N

[Ano86d] Anonymous. Notices. The Computer Journal, 29(3):286, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
29/3/286.full.pdf+html.

Anonymous:1987:AS

[Ano87a] Anonymous. Algorithm supplement. The Computer Journal, 30 (3):276, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/3/276.full.pdf+html.

Anonymous:1987:Ab

[Ano87b] Anonymous. Announcement. The Computer Journal, 30(2):109, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/2/109.full.pdf+html.

Anonymous:1987:Ad

[Ano87c] Anonymous. Announcement. The Computer Journal, 30(3):248, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/3/248.full.pdf+html.

Anonymous:1987:Ae

[Ano87d] Anonymous. Announcement. The Computer Journal, 30(3):253, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/3/253.full.pdf+html.

Anonymous:1987:Af

[Ano87e] Anonymous. Announcement. The Computer Journal, 30(3):267, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/3/267.full.pdf+html.

Anonymous:1987:Ag

[Ano87f] Anonymous. Announcement. The Computer Journal, 30(3):275, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/3/275.full.pdf+html.

Anonymous:1987:Ai

[Ano87g] Anonymous. Announcement. The Computer Journal, 30(5):392, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/5/392.full.pdf+html.

Anonymous:1987:Aa

[Ano87h] Anonymous. Announcements. The Computer Journal, 30(1):93, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/1/93.2.full.pdf+html.

Anonymous:1987:Ac

[Ano87i] Anonymous. Announcements. The Computer Journal, 30(2):118, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/2/118.full.pdf+html.

Anonymous:1987:Ah

[Ano87j] Anonymous. Announcements. The Computer Journal, 30(3):286, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/3/286.full.pdf+html.

Anonymous:1987:Aj

[Ano87k] Anonymous. Announcements. The Computer Journal, 30(5):477– 479, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/477.full.pdf+html.

Anonymous:1987:Ak

[Ano87l] Anonymous. Announcements. The Computer Journal, 30(6):573– 575, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/573.full.pdf+html.

Anonymous:1987:E

[Ano87m] Anonymous. Editorial. The Computer Journal, 30(5):385, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/5/385.full.pdf+html.

Anonymous:1988:Ab

[Ano88a] Anonymous. Announcement. The Computer Journal, 31(2):140, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/140.full.pdf+html.

Anonymous:1988:Ad

[Ano88b] Anonymous. Announcement. The Computer Journal, 31(2):164, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/164.full.pdf+html.

Anonymous:1988:Ag

[Ano88c] Anonymous. Announcement. The Computer Journal, 31(2):190, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/190.4.full.pdf+html.

Anonymous:1988:Ah

[Ano88d] Anonymous. Announcement. The Computer Journal, 31(3):200, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/3/200.full.pdf+html.

Anonymous:1988:Ar

[Ano88e] Anonymous. Announcement. The Computer Journal, 31(6):502, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/6/502.full.pdf+html.

Anonymous:1988:As

[Ano88f] Anonymous. Announcement. The Computer Journal, 31(6):516, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/6/516.full.pdf+html.

Anonymous:1988:Aa

[Ano88g] Anonymous. Announcements. The Computer Journal, 31(1):93– 96, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/1/93.full.pdf+html.

Anonymous:1988:Ac

[Ano88h] Anonymous. Announcements. The Computer Journal, 31(2):154, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/154.full.pdf+html.

Anonymous:1988:Ae

[Ano88i] Anonymous. Announcements. The Computer Journal, 31(2):174, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/174.full.pdf+html.

Anonymous:1988:Af

[Ano88j] Anonymous. Announcements. The Computer Journal, 31(2):189, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/189.full.pdf+html.

Anonymous:1988:Ai

[Ano88k] Anonymous. Announcements. The Computer Journal, 31(3):208, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/3/208.full.pdf+html.

Anonymous:1988:Aj

[Ano88l] Anonymous. Announcements. The Computer Journal, 31(3):242, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/3/242.full.pdf+html.

Anonymous:1988:Ak

[Ano88m] Anonymous. Announcements. The Computer Journal, 31(3):275, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/3/275.full.pdf+html.

Anonymous:1988:Al

[Ano88n] Anonymous. Announcements. The Computer Journal, 31(3):286– 287, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/286.2.full.pdf+html.

Anonymous:1988:Am

[Ano88o] Anonymous. Announcements. The Computer Journal, 31(4):312, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/4/312.full.pdf+html.

Anonymous:1988:An

[Ano88p] Anonymous. Announcements. The Computer Journal, 31(4):357, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/4/357.full.pdf+html.

Anonymous:1988:Ao

[Ano88q] Anonymous. Announcements. The Computer Journal, 31(4):363, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/4/363.full.pdf+html.

Anonymous:1988:Ap

[Ano88r] Anonymous. Announcements. The Computer Journal, 31(4):380a– 382, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/4/380-a.full.pdf+html.

Anonymous:1988:Aq

[Ano88s] Anonymous. Announcements. The Computer Journal, 31(5):476, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/5/476.full.pdf+html.

Anonymous:1988:N

[Ano88t] Anonymous. Notice. The Computer Journal, 31(1):96, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
31/1/96.full.pdf+html.

Anonymous:1989:Ab

[Ano89a] Anonymous. Announcement. The Computer Journal, 32(1):15, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/15.full.pdf+html.

Anonymous:1989:Ac

[Ano89b] Anonymous. Announcement. The Computer Journal, 32(1):23, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/23.full.pdf+html.

Anonymous:1989:Ad

[Ano89c] Anonymous. Announcement. The Computer Journal, 32(1):28, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/28.full.pdf+html.

Anonymous:1989:Ae

[Ano89d] Anonymous. Announcement. The Computer Journal, 32(1):54, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/54.full.pdf+html.

Anonymous:1989:Ag

[Ano89e] Anonymous. Announcement. The Computer Journal, 32(1):89, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/89.full.pdf+html.

Anonymous:1989:Ah

[Ano89f] Anonymous. Announcement. The Computer Journal, 32(1):94, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/94.full.pdf+html.

Anonymous:1989:Ai

[Ano89g] Anonymous. Announcement. The Computer Journal, 32(1):95, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/95.2.full.pdf+html.

Anonymous:1989:Am

[Ano89h] Anonymous. Announcement. The Computer Journal, 32(3):261, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/3/261.full.pdf+html.

Anonymous:1989:Ao

[Ano89i] Anonymous. Announcement. The Computer Journal, 32(4):304, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/4/304.full.pdf+html.

Anonymous:1989:Aq

[Ano89j] Anonymous. Announcement. The Computer Journal, 32(5):398, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/5/398.full.pdf+html.

Anonymous:1989:As

[Ano89k] Anonymous. Announcement. The Computer Journal, 32(5):436, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/5/436.full.pdf+html.

Anonymous:1989:At

[Ano89l] Anonymous. Announcement. The Computer Journal, 32(5):444, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/5/444.full.pdf+html.

Anonymous:1989:Au

[Ano89m] Anonymous. Announcement. The Computer Journal, 32(5):473, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/5/473.full.pdf+html.

Anonymous:1989:Av

[Ano89n] Anonymous. Announcement. The Computer Journal, 32(6):515, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/6/515.full.pdf+html.

Anonymous:1989:Aw

[Ano89o] Anonymous. Announcement. The Computer Journal, 32(6):540, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/6/540.full.pdf+html.

Anonymous:1989:Ax

[Ano89p] Anonymous. Announcement. The Computer Journal, 32(6):553, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/6/553.full.pdf+html.

Anonymous:1989:Aa

[Ano89q] Anonymous. Announcements. The Computer Journal, 32(1):12, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/12.full.pdf+html.

Anonymous:1989:Af

[Ano89r] Anonymous. Announcements. The Computer Journal, 32(1):85, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/85.full.pdf+html.

Anonymous:1989:Aj

[Ano89s] Anonymous. Announcements. The Computer Journal, 32(3):201, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/3/201.full.pdf+html.

Anonymous:1989:Ak

[Ano89t] Anonymous. Announcements. The Computer Journal, 32(3):211, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/3/211.full.pdf+html.

Anonymous:1989:Al

[Ano89u] Anonymous. Announcements. The Computer Journal, 32(3):219, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/3/219.full.pdf+html.

Anonymous:1989:An

[Ano89v] Anonymous. Announcements. The Computer Journal, 32(3):266, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/3/266.full.pdf+html.

Anonymous:1989:Ap

[Ano89w] Anonymous. Announcements. The Computer Journal, 32(4):381– 382, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/381.full.pdf+html.

Anonymous:1989:Ar

[Ano89x] Anonymous. Announcements. The Computer Journal, 32(5):412, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/5/412.full.pdf+html.

Anonymous:1989:Ay

[Ano89y] Anonymous. Announcements. The Computer Journal, 32(6):573, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/6/573.full.pdf+html.

Albano:1980:HSR

[AO80a] A. Albano and R. Orsini. A heuristic solution of the rectangular cutting stock problem. The Computer Journal, 23(4):338– 343, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/338.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/339.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/340.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/341.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/342.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/343.tif.

Albano:1980:TSA

[AO80b] A. Albano and R. Orsini. A tree search approach to the Mpartition and Knapsack problems. The Computer Journal, 23(3): 256–261, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/256.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/256.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/257.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/258.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/259.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/260.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/261.
tif.

Arabnia:1987:TNA

[AO87] H. R. Arabnia and M. A. Oliver. A transputer network for the arbitrary rotation of digitised images. The Computer Journal, 30(5):425–432, October 1987. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/5/425.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/425.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/426.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/427.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/428.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/429.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/430.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/431.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/432.tif.

Apostolatos:1989:C

[Apo89] N. Apostolatos. Correspondence. The Computer Journal, 32(3): 282, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/3/282.1.full.pdf+html.

Anderson:1986:PCS

[APW86] M. Anderson, R. D. Pose, and C. S. Wallace. A passwordcapability system. The Computer Journal, 29(1):1–8, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/1/1.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/1.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/2.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/3.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_29/Issue_01/tiff/4.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_29/
Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
7.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/8.tif.

Arblaster:1979:JSP

[ASG79] A. T. Arblaster, M. E. Sime, and T. R. G. Green. Jumping to some purpose. The Computer Journal, 22(2):105– 109, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/22/2/105.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_22/Issue_02/tiff/105.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_02/tiff/106.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_02/tiff/107.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/108.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/109.tif. See correspondence

[Hil80, Rob80a, Tay80, MR82].

Atkinson:1980:PGS

[Atk80] M. D. Atkinson. Permutation groups and set union algorithms. The Computer Journal, 23(2):187, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/23/2/187.2.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/187.tif.

Atkinson:1984:JAG

[Atk84] L. V. Atkinson. Jumping about and getting into a state. The Computer Journal, 27(1):42–46, February 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/27/1/42.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/46.
tif.

Abel:1985:REF

[AW85] D. J. Abel and W. T. Williams. A re-examination of four classificatory fusion strategies. The Computer Journal, 28(4):439– 443, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/439.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/439.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/440.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/441.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/

tiff/442.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/443.tif.

Bjorn-Andersen:1988:HFH

[BA88] N. Bjørn-Andersen. Are ‘human factors’ human? The Computer Journal, 31(5):386–390, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/5/386.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/386.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/387.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/388.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/389.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/390.
tif.

Baecker:1984:ISE

[Bae84a] H. D. Baecker. On instruction set encoding. The Computer Journal, 27(3):283, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_27/Issue_03/tiff/283.tif.

Baecker:1984:SNI

[Bae84b] H. D. Baecker. Short note: on instruction set encoding. The Computer Journal, 27(3):283, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/27/3/283.1.full.pdf+html.

Baer:1986:C

[Bae86] R. M. Baer. Correspondence. The Computer Journal, 29(4):381– 382, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/381.2.full.pdf+html.

Bailes:1988:APO

[Bai88] P. A. Bailes. An advanced problem-oriented introduction to imperative programming with Modula-2. The Computer Journal, 31(6):531–539, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/6/531.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/531.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_06/tiff/532.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/533.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/534.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/535.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/536.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/537.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/538.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/539.tif.

Baker:1980:BR

[Bak80] R. K. Baker. Book review. The Computer Journal, 23(1):45, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/45.1.full.pdf+html.

Baker:1982:DSI

[Bak82a] Geoffrey J. Baker. Database security and integrity. The Computer Journal, 25(3):400b–400, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/3/400-b.full.pdf+html.

Baker:1982:EUG

[Bak82b] Geoffrey J. Baker. An End-User’s guide to data base. The Computer Journal, 25(3):400–e–400, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/3/400-
e.full.pdf+html.

Barber:1980:BR

[Bar80] D. L. A. Barber. Book review. The Computer Journal, 23(2): 141, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/141.1.full.pdf+html.

Barnard:1982:SAC

[Bar82a] T. M. Barnard. Simplified accounting for the computer industry. The Computer Journal, 25(3):337, August 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/3/337.
full.pdf+html.

Barritt:1982:BRM

[Bar82b] M. M. Barritt. Book review: Michael Shave, Data Structures. The Computer Journal, 25(4):492b–492, November 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/4/492-
b.full.pdf+html.

Barber:1983:DPM

[Bar83a] B. Barber. Dynamic programming-models & applications. The Computer Journal, 26(4):381c–381, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/4/381-
c.full.pdf+html.

Barber:1983:BR

[Bar83b] Barry Barber. Book review. The Computer Journal, 26(2):191– 192, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/191.5.full.pdf+html.

Barton:1984:PAS

[Bar84] D. Barton. Problem with algebra systems — revisited. The Computer Journal, 27(2):159–164, May 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/2/159.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/159.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/160.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/161.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/162.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/163.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/164.tif.

Barton:1985:PSA

[Bar85] David Barton. Parallelism in simple algebra systems. The Computer Journal, 28(2):142–147, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/2/142.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/142.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_02/tiff/143.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/144.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/145.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/146.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/147.tif.

Barringer:1987:TW

[Bar87] H. Barringer. Up and down the temporal way. The Computer Journal, 30(2):134–148, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/2/134.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/134.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/135.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/136.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/137.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/138.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/139.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/140.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/141.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/142.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/143.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/144.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/145.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/147.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/148.
tif.

Bassiouni:1986:SPM

[Bas86] M. A. Bassiouni. System and program models of storage allocation for reducing seek delay. The Computer Journal, 29(1):47–51, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/1/47.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_

01/tiff/48.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/50.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/51.tif.

Barber:1980:BRB

[BB80a] D. L. A. Barber and J. M. Brady. Book review: Data Communications and Teleprocessing Systems, by T. Horsley, 1979; A Technique for Software and Systems Design, by R. J. Lano, 1979. The Computer Journal, 23(2):141, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/141.tif.

Bishop:1980:PCS

[BB80b] J. M. Bishop and D. W. Barron. Procedure calling and structured architecture. The Computer Journal, 23(2):115– 123, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/115.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/115.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/117.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/119.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/120.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/121.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/122.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/

tiff/123.tif. See correspondence [ME80].

Bishop:1981:PD

[BB81] J. M. Bishop and D. W. Barron. Principles of descriptors. The Computer Journal, 24(3):210–221, August 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/3/210.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/210.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/211.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/212.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/213.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/214.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/215.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/216.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/218.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/219.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/221.tif.

Bishop:1982:MD

[BB82] J. M. Bishop and D. W. Barron. Mu5 descriptors. The Computer Journal, 25(3):399a–399, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/3/399-a.full.pdf+html.

Brailsford:1989:EPJ

[BB89] D. F. Brailsford and R. J. Beach. Electronic publishing — a journal and its production. The Computer Journal, 32(6):482– 493, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/482.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/482.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/483.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/484.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/485.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/486.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/487.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/488.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/489.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/490.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/491.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/492.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/493.tif.

Baeten:1987:RTS

[BBK87] J. C. M. Baeten, J. A. Bergstra, and J. W. Klop. Readytrace semantics for concrete process algebra with the priority operator. The Computer Journal, 30(6):498–506, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/498.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/498.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/499.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/500.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/501.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/502.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/503.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/504.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/505.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/506.tif.

Bevan:1989:PDD

[BBKR89] D. I. Bevan, G. L. Burn, R. J. Karia, and J. D. Robson. Principles for the design of a distributed memory architecture for parallel graph reduction. The Computer Journal, 32(5):461– 469, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/5/461.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/461.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/462.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/463.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/464.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/465.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/466.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/467.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/468.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/469.tif.

Boccalatte:1980:ADS

[BD80] A. Boccalatte and M. Di Manzo. An approach to the detection of semantic errors. The Computer Journal, 23(4):317–323, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_04/tiff/317.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_04/tiff/318.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_04/tiff/319.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/320.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/321.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/322.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/323.
tif.

Bateman:1981:CAS

[BD81] S. O. Bateman, Jr. and S. D. Danielopoulos. Computerised analytic solutions of second order differential equations. The Computer Journal, 24(2):180–183, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/180.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/180.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/181.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/182.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/183.tif.

Bell:1982:KSC

[BD82] D. A. Bell and S. M. Deen. Key space compression and hashing in PRECI. The Computer Journal, 25(4):486–492, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/486.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/486.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/487.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/488.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/489.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/490.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/491.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/492.tif.

Bell:1984:HTV

[BD84] D. A. Bell and S. M. Deen. Hash trees versus B-trees. The Computer Journal, 27(3):218–224, August 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/3/218.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/218.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/219.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/220.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/221.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/222.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/224.tif.

Boccalatte:1982:ERA

[BDS82] A. Boccalatte, M. Di Manzo, and D. Sciarra. Error recovery with attribute grammars. The Computer Journal, 25(3):331– 337, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/331.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/331.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/332.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/333.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/334.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/335.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/336.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/337.tif.

Barlow:1982:PAI

[BE82] R. H. Barlow and D. J. Evans. Parallel algorithms for the iterative solution to linear systems. The Computer Journal, 25(1):56–60, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/25/1/56.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/56.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/57.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/58.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/59.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/60.tif.

Beauchamp:1981:BR

[Bea81] K. G. Beauchamp. Book review. The Computer Journal, 24 (1):51, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/1/51.full.pdf+html.

Bell:1981:C

[Bel81a] D. Bell. Correspondance. The Computer Journal, 24(2):191, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/191.1.full.pdf+html.

Bell:1981:BR

[Bel81b] D. M. R. Bell. Book review. The Computer Journal, 24(4):315, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/315.full.pdf+html.

Bell:1982:BRJ

[Bel82] A. G. Bell. Book review: James Martin, Telematic Society. The Computer Journal, 25(4):422a–422, November 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/4/422-
a.full.pdf+html.

Bell:1984:DDP

[Bel84] David A. Bell. Difficult data placement problems. The Computer Journal, 27(4):315–320, November 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/4/315.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/315.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/316.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/317.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/318.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/319.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/320.tif.

Bennett:1981:BR

[Ben81] P. A. Bennett. Book review. The Computer Journal, 24(4):366, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/366.2.full.pdf+html.

Bengtsson:1982:PRP

[Ben82a] Bengt Erik Bengtsson. Packing rectangular pieces — a heuristic approach. The Computer Journal, 25(3):353–357, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/353.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/353.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/354.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/355.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/356.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/357.tif.

Bennett:1982:BRC

[Ben82b] P. A. Bennett. Book review: Caxton C. Foster, Real-Time Programming — Neglected Topics. The Computer Journal, 25 (4):494b–494, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/4/494-b.full.pdf+html.

Bennett:1982:DCU

[Ben82c] P. A. Bennett. Digital control using microprocessors. The Computer Journal, 25(4):496a–496, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/4/496-
a.full.pdf+html.

Bennett:1982:ECP

[Ben82d] P. A. Bennett. Elements of computer process control. The Computer Journal, 25(3):396b–396, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/3/396-
b.full.pdf+html.

Bennett:1982:FTP

[Ben82e] P. A. Bennett. Fault tolerance: Principles and practice. The Computer Journal, 25(3):400–d–400, August 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/3/400-
d.full.pdf+html.

Bennett:1982:PRE

[Ben82f] P. A. Bennett. Practical reliability engineering. The Computer Journal, 25(3):337c–337, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/3/337-c.full.pdf+html.

Bennett:1982:SM

[Ben82g] P. A. Bennett. Software metrics. The Computer Journal, 25 (3):400a–400, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/3/400-a.full.pdf+html.

Benyon:1985:EVC

[Ben85] Peter R. Benyon. Exploiting vector computers by replication. The Computer Journal, 28(2):138–141, May 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/2/138.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/138.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/139.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/140.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/141.tif.

Berghel:1987:CCH

[Ber87a] H. Berghel. Crossword compilation with Horn clauses. The Computer Journal, 30(2):183–188, April 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/2/183.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/183.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/184.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/

tiff/185.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/186.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/187.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/188.tif.

Bertino:1987:EPI

[Ber87b] Elisa Bertino. An evaluation of precompilation and interpretation in distributed database management systems. The Computer Journal, 30(6):519–528, December 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/6/519.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/519.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/520.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/521.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/522.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/523.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/524.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/525.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/526.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/527.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/528.
tif.

Barlow:1983:PMA

[BES83] R. H. Barlow, D. J. Evans, and J. Shanehchi. Parallel multisection applied to the eigenvalue problem. The Computer Journal, 26(1): 6–9, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/1/6.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_26/Issue_01/tiff/6.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_26/
Issue_01/tiff/7.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
9.tif.

Brinck:1981:AAT

[BF81] K. Brinck and N. Y. Foo. Analysis of algorithms on threaded trees. The Computer Journal, 24(2):148–155, May 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/148.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/148.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/149.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/150.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/151.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/152.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/153.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/154.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/155.tif.

Boardman:1986:RAO

[BGR86] J. T. Boardman, C. Garrett, and G. C. A. Robson. A recursive algorithm for the optimal solution of a complex allocation problem using a dynamic programming formulation. The Computer Journal, 29(2):182–186, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/2/182.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/182.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/183.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/184.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/185.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/186.
tif.

Burton:1985:MGT

[BHK85] F. Warren Burton, Matthew M. Huntbach, and J. (Yiannis) G. Kollias. Multiple generation text files using overlapping tree structures. The Computer Journal, 28(4):414–416, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/414.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/414.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/415.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/416.tif.

Bloss:1989:OCM

[BHY89] Adrienne Bloss, P. R. Hudak, and J. H. Young. An optimising compiler for a modern functional language. The Computer Journal, 32(2):152–161, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/2/152.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/152.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/153.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/154.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/155.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/156.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/158.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/160.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/161.
tif.

Bird:1981:DCD

[Bir81] P. R. Bird. Discussion and correspondence: Data structures for catalogues of sound recordings. The Computer Journal, 24(1):92– 94, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/1/92.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_24/Issue_01/tiff/92.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_24/
Issue_01/tiff/93.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/94.tif.

Bird:1989:AIP

[Bir89] R. S. Bird. Algebraic identities for program calculation. The Computer Journal, 32(2):122–126, April 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/2/122.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/122.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/123.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/124.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_32/Issue_02/tiff/125.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/126.
tif.

Bastani:1988:CMD

[BIY88] F. B. Bastani, S. S. Iyengar, and I-Ling Yen. Concurrent maintenance of data structures in a distributed environment. The Computer Journal, 31(2):165–174, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/31/2/165.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/165.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/166.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/167.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/168.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/169.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/170.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/171.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/172.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/173.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/174.tif.

Bryn-Jones:1980:CPS

[BJ80] Mervyn Bryn-Jones. Correspondence: Points and n-sided irregular figures. The Computer Journal, 23(2):190, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/190.tif.

Burton:1983:SNC

[BK83a] F. W. Burton and J. G. Kollias. Short notes: Comment on ‘the explicit quad tree as a structure for computer graphics’. The Computer Journal, 26(2):188, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/2/188.full.pdf+html. See

[Woo82].

Burton:1983:CEQ

[BK83b] F. Warren Burton and J. G. Kollias. Comment on “The explicit quad tree as a structure for computer graphics”. The Computer Journal, 26(2):188, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_26/Issue_02/tiff/188.tif.

Backhouse:1987:WRM

[BK87] R. C. Backhouse and A. A. Khamiss. A while-rule in Martin-Löf’s theory of types. The Computer Journal, 30(1):27–36, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
30/1/27.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/27.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
28.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/30.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
31.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/32.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/33.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
34.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/35.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/36.tif.

Burton:1984:SNC

[BKK84] F. W. Burton, V. J. Kollias, and J. G. Kollias. Short notes: Consistency in point-in-polygon tests. The Computer Journal, 27(4):375– 376, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/375.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/375.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/376.tif.

Burton:1987:GPP

[BKK87] F. W. Burton, V. J. Kollias, and J. G. Kollias. A general PASCAL program for map overlay of quadtrees and related problems. The Computer Journal, 30(4):355–361, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/355.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/355.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

30/Issue_04/tiff/356.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/357.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/358.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/359.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/360.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/361.tif.

Bell:1981:BRB

[BL81a] A. G. Bell and R. Lovett. Book reviews: Software Reliability Guidebook, by Robert L. Glass, 1979; Foundations of Computer Science, by M. S. Carberry, H. M. Khalil and J. F. Leathrum and L. S. Levy, 1979. The Computer Journal, 24(1):91, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/91.tif.

Burton:1981:ADT

[BL81b] F. Warren Burton and Brian J. Lings. Abstract data types, subtypes and data independence. The Computer Journal, 24(4):308– 311, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/308.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/308.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/309.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/310.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/311.tif.

Blake:1982:FPS

[Bla82] A. Blake. Fixed point solutions of recursive operations on Boolean arrays. The Computer Journal, 25(2):231–234, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/25/2/231.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/231.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/232.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/233.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/234.tif.

Blythe:1980:BR

[Bly80] M. J. Blythe. Book review. The Computer Journal, 23(1):65, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/65.full.pdf+html.

Bratko:1980:APC

[BM80] I. Bratko and D. Michie. An advice program for a complex chess programming task. The Computer Journal, 23(4):353–359, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/353.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/355.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/356.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/357.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/358.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/359.tif.

Bell:1981:CES

[BM81] D. Bell and Donald Michie. Correspondence: Expert systems. The Computer Journal, 24(2):191, May 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/191.tif. See [Mic80].

Barlow:1983:C

[BM83] R. H. Barlow and S. MacQueen. Correspondence. The Computer Journal, 26(2):189, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/2/189.1.full.pdf+html.

Bunt:1984:MLB

[BM84] Richard B. Bunt and Jennifer M. Murphy. The measurement of locality and the behaviour of programs. The Computer Journal, 27(3):238–245, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/3/238.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

27/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/239.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/241.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/242.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/243.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/244.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/245.tif.

Barringer:1986:PSA

[BM86] H. Barringer and I. Mearns. A proof system for Ada tasks. The Computer Journal, 29(5):404–415, October 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/5/404.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/404.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/405.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/406.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/407.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/408.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/409.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/410.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/411.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/412.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/413.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/414.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/415.tif.

Benyon:1988:EAI

[BM88] D. Benyon and Dianne Murray. Experience with adaptive interfaces. The Computer Journal, 31(5):465–473, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/5/465.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/465.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/466.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_05/tiff/467.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/468.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/469.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/470.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/471.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/472.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/473.tif.

Bell:1989:NDD

[BM89] T. Bell and A. Moffat. A note on the DMC data compression scheme. The Computer Journal, 32(1):16–20, February 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/32/1/16.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/17.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/18.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/19.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
20.tif.

Blair:1987:KBO

[BMNS87] G. S. Blair, J. A. Mariani, J. R. Nicol, and D. Shepherd. A knowledge-based operating system. The Computer Journal, 30(3): 193–200, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/3/193.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/193.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/195.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/197.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/198.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/199.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/200.tif.

Bell:1988:CRT

[BMSA88] D. A. Bell, F. J. McErlean, P. M. Stewart, and W. Arbuckle. Clustering related tuples in databases. The Computer Journal, 31(3): 253–257, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/253.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/253.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/255.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/256.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/257.tif.

Bonanno:1985:AAS

[BMT85] R. Bonanno, D. Maio, and P. Tiberio. An approximation algorithm for secondary index selection in relational database physical design. The Computer Journal, 28(4):398–405, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/398.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/398.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/399.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/400.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/401.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/402.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/403.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/404.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/405.tif.

Bogod:1983:CCB

[Bog83] J. L. Bogod. CAD/CAM as a basis for the development of technology in developing nations. The Computer Journal, 26(4):383c– 383, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/383-c.full.pdf+html.

Bowyer:1981:CDT

[Bow81] A. Bowyer. Computing Dirichlet tessellations. The Computer Journal, 24(2):162–166, May 1981. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/162.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/165.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/166.
tif.

Bowles:1983:DAS

[Bow83] J. A. Bowles. Data acquisition systems. The Computer Journal, 26(4):382c–382, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/4/382-c.full.pdf+html.

Bowers:1984:DAA

[Bow84] David S. Bowers. A database architecture for aggregateincomplete data. The Computer Journal, 27(4):294–300, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/294.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/294.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/295.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/296.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/297.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/298.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/299.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/300.tif.

Bowker:1987:C

[Bow87] A. V. Bowker. Correspondence. The Computer Journal, 30(2): 191, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/2/191.1.full.pdf+html.

Butler:1981:FFD

[BP81] G. F. Butler and J. Pike. A free-format data input scheme written in standard Fortran (ansi 66). The Computer Journal, 24(3):256– 257, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/256.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/256.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/257.tif.

Bozyigit:1982:TRM

[BP82] M. Bozyigit and Y. Paker. A topology reconfiguration mechanism for distributed computer systems. The Computer Journal, 25(1): 87–92, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/87.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_25/Issue_01/tiff/87.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_25/
Issue_01/tiff/88.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
90.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/92.tif.

Basu:1987:OPL

[BPG87] Julie Basu, L. M. Patnaik, and A. K. Goswami. Ordered Ports — a language concept for high-level distributed programming. The Computer Journal, 30(6):487–497, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/487.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/487.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/488.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/489.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/490.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/491.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/492.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/493.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/494.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/495.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/496.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/497.
tif.

Boardman:1986:TPS

[BR86] J. T. Boardman and G. C. A. Robson. Towards a problemsolving methodology for coping with increasing complexity: an engineering approach. The Computer Journal, 29(2):161– 166, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/161.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/161.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/165.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/166.
tif.

Brady:1980:BR

[Bra80a] J. M. Brady. Book review. The Computer Journal, 23(2):141, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/141.2.full.pdf+html.

Bramer:1980:COS

[Bra80b] M. A. Bramer. Correct and optimal strategies in game playing programs. The Computer Journal, 23(4):347–352, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_04/tiff/347.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_04/tiff/348.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/350.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/352.tif.

Bradley:1985:UMD

[Bra85] James Bradley. Use of mean distance between overflow records to compute average search lengths in hash files with open addressing. The Computer Journal, ??(??):??, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Bradley:1986:SNJ

[Bra86a] J. Bradley. Short notes: Join dependencies in relational databases and the geometry of spatial Grids. The Computer Journal, 29(4): 378–379, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/378.full.pdf+html.

Bradley:1986:UMD

[Bra86b] J. Bradley. Use of mean distance between overflow records to compute average search lengths in hash files with open addressing. The Computer Journal, 29(2):167–170, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/29/2/167.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/167.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/169.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/170.tif.

Bradley:1986:JDR

[Bra86c] James Bradley. Join dependencies in relational databases and the geometry of spatial grids. The Computer Journal, 29(4):378– 380, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/378.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/379.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/380.tif.

Bradley:1988:PJD

[Bra88] J. Bradley. Polygon join dependencies, closed co-relationship chains and the connection trap in relational databases. The Computer Journal, 31(2):141–146, April 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/2/141.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/141.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/142.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/143.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/144.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/146.tif.

Breiner:1981:C

[Bre81a] C. G. Breiner. Correspondence. The Computer Journal, 24(3): 270, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/270.2.full.pdf+html.

Brent:1981:CCP

[Bre81b] R. P. Brent. Correspondence: Comments on papers by Maddison and Westreich. The Computer Journal, 24(1):95–96, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/95.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/96.

tif. See [Mad80, Wes80b].

Bresenham:1982:ILC

[Bre82] Jack E. Bresenham. Incremental line compaction. The Computer Journal, 25(1):116–120, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/1/116.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/117.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/119.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/120.
tif.

Brittain:1980:DCE

[Bri80a] J. N. G. Brittain. Design for a changing environment. The Computer Journal, 23(1):13–22, February 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/14.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/15.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/16.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
17.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/18.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/19.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
20.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/21.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/22.tif.

Brittan:1980:DCE

[Bri80b] J. N. G. Brittan. Design for a changing environment. The Computer Journal, 23(1):13–19, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/1/13.full.pdf+html.

Bridges:1981:BR

[Bri81a] A. H. Bridges. Book review. The Computer Journal, 24(2):117, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/117.3.full.pdf+html.

Brinson:1981:BR

[Bri81b] M. E. Brinson. Book review. The Computer Journal, 24(1):24, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/24.full.pdf+html.

Brittan:1981:BR

[Bri81c] J. N. G. Brittan. Book review. The Computer Journal, 24(2): 117, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/117.1.full.pdf+html.

Brittan:1982:BRD

[Bri82] J. Brittan. Book review: D. Millington, Systems Analysis and Design for Computer Applications. The Computer Journal, 25(4): 422, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/422.full.pdf+html.

Brinck:1985:EPT

[Bri85] Keith Brinck. The expected performance of traversal algorithms in binary trees. The Computer Journal, 28(4):426–432, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/426.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/426.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/427.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/428.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/429.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/430.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/431.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/432.tif.

Brooker:1982:DSB

[Bro82] R. A. Brooker. A ‘database’ subsystem for BCPL. The Computer Journal, 25(4):448–464, November 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/4/448.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/448.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/449.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/450.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/451.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/452.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/453.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/454.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/455.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/456.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/457.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/458.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/459.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/460.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/461.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/462.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/463.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/464.tif.

Broyden:1983:MSA

[Bro83] C. G. Broyden. A mark-scaling algorithm. The Computer Journal, 26(2):109–112, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/2/109.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/109.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/110.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/111.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/112.tif.

Brown:1984:TAI

[Bro84a] D. H. J. Brown. Transitions to Ada: an incremental approach. The Computer Journal, 27(1):37–41, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/27/1/37.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/37.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/39.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/40.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/41.
tif.

Brown:1984:EMP

[Bro84b] P. J. Brown. Embedded macro processors. The Computer Journal, 27(4):348–353, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/4/348.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/348.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/349.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/350.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/351.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/352.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/353.tif.

Brokate:1986:C

[Bro86] K. Brokate. Correspondence. The Computer Journal, 29(3):285, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/3/285.1.full.pdf+html.

Brokate:1988:Ca

[Bro88a] K. Brokate. Correspondence. The Computer Journal, 31(2):190, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/190.1.full.pdf+html.

Brokate:1988:Cb

[Bro88b] K. Brokate. Correspondence. The Computer Journal, 31(2):190, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/190.2.full.pdf+html.

Brown:1989:BR

[Bro89a] Heather Brown. Book review. The Computer Journal, 32(6): 532, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/532.full.pdf+html.

Brown:1989:SSD

[Bro89b] Heather Brown. Standards for structured documents. The Computer Journal, 32(6):505–514, December 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/6/505.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/505.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/506.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/507.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/508.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/509.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/510.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/511.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/512.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/513.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/514.
tif.

Buckley:1984:SNH

[BS84] G. N. Buckley and A. Silberschatz. Short notes: on the heterogeneous guard locking protocol. The Computer Journal, 27(1):86–87, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/1/86.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/87.tif.

Blair:1986:DTD

[BS86] A. Blair and I. Sommerville. DSA — a tool for descriptive text analysis. The Computer Journal, 29(5):460–466, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/460.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/460.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/461.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/462.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/463.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/464.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/465.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/466.tif.

Benyon:1987:TTK

[BS87] D. Benyon and S. Skidmore. Towards a tool kit for the systems analyst. The Computer Journal, 30(1):2–7, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/1/2.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_30/Issue_01/tiff/2.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_
01/tiff/3.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/4.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_30/Issue_01/tiff/5.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_30/
Issue_01/tiff/6.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/7.tif.

Bennett:1989:NRB

[BS89] J. P. Bennett and G. C. Smith. The need for reduced byte stream instruction sets. The Computer Journal, 32(4):370–373, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/370.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/370.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/371.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/372.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/373.tif.

Barsky:1981:TSR

[BT81] Brian A. Barsky and Spencer W. Thomas. TRANSPLINE — a system for representing curves using transformations among four spline formulations. The Computer Journal, 24(3):271– 277, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/271.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/271.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/272.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/273.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/274.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/275.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/276.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/277.tif.

Burton:1980:BR

[Bur80a] C. T. Burton. Book review. The Computer Journal, 23(2):106, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/106.full.pdf+html.

Burton:1980:BRB

[Bur80b] C. T. Burton. Book review: Artificial Intelligence, Volume 1 by P. H. Winston and R. H. Brown. The Computer Journal, 23(2):106, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/106.tif.

Burkimsher:1987:CRS

[Bur87a] P. C. Burkimsher. Combinator reduction in a shared-memory multiprocessor. The Computer Journal, 30(3):214–222, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/3/214.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/214.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/215.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/216.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/218.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/219.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/221.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/222.tif.

Burton:1987:FPC

[Bur87b] F. W. Burton. Functional programming for concurrent and distributed computing. The Computer Journal, 30(5):437–450, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/437.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/437.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/438.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/439.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/440.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/441.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/442.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/443.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/444.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/445.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/446.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/447.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/448.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_05/tiff/449.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/450.tif.

Burton:1988:NRT

[Bur88] F. W. Burton. Nondeterminism with referential transparency in functional programming languages. The Computer Journal, 31(3): 243–247, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/243.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/243.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/244.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/245.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/246.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/247.tif.

Barnard:1988:AMC

[BWM88] P. Barnard, M. Wilson, and A. Maclean. Approximate modelling of cognitive activity with an expert system: a theorybased strategy for developing an interactive design tool. The Computer Journal, 31(5):445–456, October 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/5/445.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/445.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/446.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/447.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/448.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/449.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/450.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/451.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/452.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/453.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/454.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/455.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/456.tif.

Bowyer:1981:MAS

[BWW81] Adrian Bowyer, Philip J. Willis, and John R. Woodwark. A multiprocessor architecture for solving spatial problems. The Computer Journal, 24(4):353–357, November 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/4/353.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/353.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/354.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/355.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/356.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/357.
tif.

Berghel:1989:CCC

[BY89] H. Berghel and C. Yi. Crossword compiler-compilation. The Computer Journal, 32(3):276–280, June 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/3/276.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/277.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/278.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/279.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/280.
tif.

Carlson:1982:AFD

[CAC82] C. Robert Carlson, Adarsh K. Arora, and Miroslava Milosavljevic Carlson. The application of functional dependency theory to relational databases. The Computer Journal, 25(1):68–73, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/1/68.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/68.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
69.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/70.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/71.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/

72.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/73.tif.

Cairns:1982:IP

[Cai82] David J. Cairns. Invitation to Pascal. The Computer Journal, 25(4):496, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/4/496.full.pdf+html.

Cairns:1983:C

[Cai83] D. J. Cairns. Correspondence. The Computer Journal, 26(2): 189–190, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/189.4.full.pdf+html.

Capon:1980:UWC

[Cap80] P. C. Capon. Understanding and writing compilers. The Computer Journal, 23(3):273, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/273.full.pdf+html.

Capper:1986:PTC

[Cap86] L. Capper. A philosophy for the teaching of computer science and information technology. The Computer Journal, 29(1):83–89, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/1/83.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/83.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/84.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/86.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/87.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/88.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
89.tif.

Carden:1986:SSC

[Car86] J. Carden. The structural stability of corporate and devolved databases. The Computer Journal, 29(4):361–367, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/361.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/361.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/362.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/363.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/364.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/365.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/367.tif.

Cavouras:1983:CSE

[Cav83] John C. Cavouras. Computer system evaluation through supervisor replication. The Computer Journal, 26(2):134–141, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/134.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/134.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/135.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/136.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/137.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/138.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/139.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/140.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/141.tif.

Chang:1987:PAG

[CC87] C. C. Chang and C. Y. Chen. Performance analysis of the generalised disc modulo allocation method for multiple key hashing files on multi-disc systems. The Computer Journal, 30(6):535– 540, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/535.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/535.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/536.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/537.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/538.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_30/Issue_06/tiff/539.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/540.
tif.

Chang:1988:SNR

[CC88] C. C. Chang and C. Y. Cheng. Short notes: a recursive performance formula of the disc modulo allocation method for binary Cartesian product files. The Computer Journal, 31(6):565–566, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/565.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/565.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/566.tif.

Cantone:1986:WFC

[CCD86] G. Cantone, A. Cimitile, and U. De Carlini. Well-formed conversion of unstructured One-in/one-out schemes for complexity measurement and program maintenance. The Computer Journal, 29(4):322–329, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/4/322.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/322.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/323.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/324.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/325.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/326.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/327.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/328.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/329.tif.

Coulson:1987:PNA

[CCL87] A. F. W. Coulson, J. F. Collins, and A. Lyall. Protein and nucleic acid sequence database searching: a suitable case for parallel processing. The Computer Journal, 30(5):420–424, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/420.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/420.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

30/Issue_05/tiff/421.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/422.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/423.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/424.tif.

Coulter:1987:ITC

[CCS87] N. S. Coulter, R. B. Cooper, and M. K. Solomon. Informationtheoretic complexity of program specifications. The Computer Journal, 30(3):223–227, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/3/223.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/224.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/226.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/227.
tif.

Clark:1980:ACT

[CD80] K. L. Clark and J. Darlington. Algorithm classification through synthesis. The Computer Journal, 23(1):61–65, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
23/1/61.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/61.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
62.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/63.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/64.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
65.tif.

Cavouras:1981:STC

[CD81] J. C. Cavouras and R. H. Davis. Simulation tools in computer system design methodologies. The Computer Journal, 24(1):25–28, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/25.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/26.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_24/Issue_01/tiff/27.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/28.
tif.

Cocu:1982:CAS

[CD82] J. J. Cocu and R. E. Devillers. On a class of allocation strategies inducing bounded delays only. The Computer Journal, 25(1):52–55, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/52.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/52.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/53.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/54.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/55.tif.

Crookes:1985:AOP

[CD85] D. Crookes and C. J. Devlin. An algorithm for optimal procedural nesting. The Computer Journal, 28(1):17–21, February 1985. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/28/1/17.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/17.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/18.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
21.tif.

Cairns:1981:BRB

[CF81] D. J. Cairns and P. R. Freeman. Book reviews: Theory and Design of Digital Computer Systems, 2nd ed., by Douglas Lewin, 1980; Statistical Computing, by W. J. Kennedy and J. E. Gentle, 1980. The Computer Journal, 24(1):96, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/96.1.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_24/Issue_01/tiff/96.tif.

See [Mad80, Wes80b].

Cole:1988:ECS

[CF88] R. Cole and Clare Foxcroft. An experiment in clock synchronisation. The Computer Journal, 31(6):496–502, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/496.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/496.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/497.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/498.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/499.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/500.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/501.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/502.tif.

Cosnard:1989:GPV

[CF89] M. Cosnard and A. G. Ferreira. Generating permutations on a VLSI suitable linear network. The Computer Journal, 32(6):571– 573, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/571.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/571.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/572.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/573.tif.

Corsini:1984:IAO

[CFL84] P. Corsini, G. Frosini, and L. Lopriore. The implementation of abstract objects in a capability based addressing architecture. The Computer Journal, 27(2):127–134, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/27/2/127.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/127.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/128.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/129.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/130.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/131.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/132.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/133.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/134.tif.

Corsini:1989:PIL

[CFS89] P. Corsini, G. Frosini, and G. Speranza. The parallel interpretation of logic programs in distributed architectures. The Computer Journal, 32(1):29–35, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/29.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/30.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/31.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/32.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/33.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/34.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/35.tif.

Colwell:1985:SNN

[CG85] D. J. Colwell and J. R. Gillett. Short notes: a note on Broyden’s mark scaling algorithm. The Computer Journal, 28(2):111, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/2/111.full.pdf+html.

Cowell:1980:SSA

[CGK80] D. F. Cowell, D. F. Gillies, and A. A. Kaposi. Synthesis and structural analysis of abstract programs. The Computer Journal, 23(3): 243–247, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/243.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/243.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/244.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/245.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/246.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/247.tif.

Cormack:1987:DCU

[CH87] G. V. Cormack and R. N. S. Horspool. Data compression using dynamic Markov modelling. The Computer Journal, 30(6):541– 550, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/541.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/541.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/542.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/543.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/544.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/545.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/546.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/547.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/548.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/549.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/550.tif.

Chambers:1981:CSC

[Cha81a] A. D. Chambers. Current strategies for computer auditing within an organisation. The Computer Journal, 24(4):290–294, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/290.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/290.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/291.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/292.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/293.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/294.tif.

Chang:1981:BR

[Cha81b] C. C. Chang. Book review. The Computer Journal, 24(4):384, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/384.3.full.pdf+html.

Chang:1982:DBA

[Cha82] C. C. Chang. Data base administration. The Computer Journal, 25(1):160, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/1/160.full.pdf+html.

Chang:1987:IPS

[Cha87] C. C. Chang. An information protection scheme based upon number theory. The Computer Journal, 30(3):249–253, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/3/249.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/249.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/251.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/253.tif.

Cheng:1982:BRL

[Che82] B. K. F. Cheng. Book review: L. L. Lapin, Quantitative Methods for Business Decisions (2nd Edn). The Computer Journal, 25(4):494a–494, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/4/494-a.full.pdf+html.

Chen:1986:DAT

[Che86a] T. Y. Chen. Dynamic approach towards software science. The Computer Journal, 29(2):189, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
02/tiff/189.tif.

Chen:1986:SND

[Che86b] T. Y. Chen. Short note: a dynamic approach towards software science. The Computer Journal, 29(2):189, ???? 1986. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/29/2/189.
full.pdf+html.

Cormack:1985:PPH

[CHK85] G. V. Cormack, R. N. S. Horspool, and M. Kaiserswerth. Practical perfect hashing. The Computer Journal, 28(1):54–58, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
28/1/54.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/54.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
55.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/56.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/57.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
58.tif.

Cho:1982:ARF

[Cho82] J. R. Lin Sin Cho. Automatic report formatting from a report specimen. The Computer Journal, 25(2):242–247, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/25/2/242.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/242.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/243.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/244.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/246.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/247.
tif.

Clint:1985:ADF

[CHPS85] M. Clint, C. Holt, R. H. Perrott, and A. Stewart. Algorithm 120: A DAP FORTRAN subroutine for the eigensolution of real symmetric matrices. The Computer Journal, 28(3):340– 342, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/340.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/342.tif.

Chan:1983:AS

[CL83a] N. N. Chan and K-H. Li. Algorithms supplement. The Computer Journal, 26(2):184–186, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/2/184.full.pdf+html.

Chan:1983:AFS

[CL83b] N. N. Chan and Kim-Hung Li. Algorithm 115: a FORTRAN subroutine for finding a real symmetric matrix with prescribed diagonal elements and eigenvalues. The Computer Journal, 26(2): 184–186, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/184.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/185.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/186.tif.

Chang:1986:LOM

[CL86] C. C. Chang and R. C. T. Lee. A letter-oriented minimal perfect hashing scheme. The Computer Journal, 29(3):277– 281, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/277.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/277.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/278.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/279.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/280.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/281.tif.

Clarke:1981:BR

[Cla81] P. A. Clarke. Book review. The Computer Journal, 24(3):234, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/234.2.full.pdf+html.

Clarke:1988:JUI

[Cla88] R. A. Clarke. Judicial understanding of information technology: the case of the Wombat ROMs. The Computer Journal, 31(1):25– 33, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/1/25.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_31/Issue_01/tiff/25.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_31/
Issue_01/tiff/26.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/27.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
28.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

31/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/30.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
31.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/32.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/33.tif.

Clocksin:1986:ASS

[Clo86] W. F. Clocksin. Automatic specialisation of standard designs. The Computer Journal, 29(6):495–499, December 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/6/495.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/495.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/496.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/497.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/498.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/499.
tif.

Clocksin:1987:PDP

[Clo87] W. F. Clocksin. Principles of the Delphi Parallel Inference Machine. The Computer Journal, 30(5):386–392, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/386.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/386.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/387.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/388.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/389.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/390.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/391.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/392.tif.

Cluley:1980:DCB

[Clu80] J. C. Cluley. Digital circuits for binary arithmetic. The Computer Journal, 23(3):269, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/269.full.pdf+html.

Cluley:1981:BR

[Clu81] J. C. Cluley. Book review. The Computer Journal, 24(4):330, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/330.1.full.pdf+html.

Culberson:1989:EBB

[CM89] J. Culberson and J. I. Munro. Explaining the behaviour of binary search trees under prolonged updates: a model and simulations. The Computer Journal, 32(1):68–75, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/68.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/68.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/69.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/70.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/71.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/72.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/73.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/74.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
75.tif.

Colin:1980:ST

[CMS80] A. J. T. Colin, A. D. McGettrick, and P. D. Smith. Sorting trains. The Computer Journal, 23(3):270–273, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/23/3/270.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/270.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/271.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/272.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/273.tif.

Chan:1982:ECA

[CN82] Arvola Chan and Bahram Niamir. On estimating the cost of accessing records in blocked database organizations. The Computer Journal, 25(3):368–374, August 1982. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/3/368.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/368.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/369.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/370.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/371.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/373.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/374.tif.

Coates:1983:C

[Coa83] P. J. Coates. Correspondence. The Computer Journal, 26(1):95, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/1/95.2.full.pdf+html.

Colemen:1980:SE

[Col80] Bernard Colemen. Software engineering. The Computer Journal, 23(3):200a–200, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/200-a.full.pdf+html.

Colbourn:1982:CCI

[Col82] Charles J. Colbourn. Computing the chromatic index of Steiner triple systems. The Computer Journal, 25(3):338–339, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/338.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/338.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/339.tif.

Coleman:1983:QTP

[Col83] Rodney Coleman. Queueing theory: a problem solving approach. The Computer Journal, 26(4):383, November 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/4/383.
full.pdf+html.

Cole:1987:CTR

[Col87] A. J. Cole. Compaction techniques for raster scan graphics using space-filling curves. The Computer Journal, 30(1):87–92, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
30/1/87.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/87.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
88.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/89.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/90.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
91.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/92.tif.

Cole:1989:EEP

[Col89] Fred Cole. Editorial — electronic publishing. The Computer Journal, 32(6):481, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/6/481.full.pdf+html.

Compton:1985:AED

[Com85] A. J. Compton. An algorithm for the even distribution of entities in one dimension. The Computer Journal, 28(5):530–537, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/530.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/530.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/531.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/532.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/533.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/534.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/535.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/536.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/537.tif.

Cookson:1983:TSC

[Coo83] M. J. Cookson. Taxonomic studies on current approaches to systems analysis. The Computer Journal, 26(3):283–284, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/283.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/283.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/284.tif.

Cook:1989:EOO

[Coo89a] S. Cook. Editorial — object-oriented systems. The Computer Journal, 32(4):289, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/4/289.full.pdf+html.

Cook:1989:PME

[Coo89b] W. R. Cook. A proposal for making Eiffel type-safe. The Computer Journal, 32(4):305–311, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/4/305.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/305.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/306.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/307.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/308.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/309.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/310.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/311.tif.

Cooper:1989:FHO

[Coo89c] M. C. Cooper. Formal hierarchical object models for fast template matching. The Computer Journal, 32(4):351–361, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/351.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/351.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/352.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/353.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/355.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/356.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

32/Issue_04/tiff/357.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/358.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/359.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/360.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/361.
tif.

Corner:1981:BR

[Cor81] L. Corner. Book review. The Computer Journal, 24(4):330, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/330.2.full.pdf+html.

Cowling:1986:TCP

[Cow86] A. J. Cowling. Type checking in polymorphic languages. The Computer Journal, 29(6):538–544, December 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/6/538.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/538.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/539.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/540.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/541.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/542.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/543.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/544.tif.

Castle:1986:C

[CP86] C. M. Castle and M. L. V. Pitteway. Correspondence. The Computer Journal, 29(6):574, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/29/6/574.full.pdf+html.

Castle:1987:EST

[CP87] C. M. A. Castle and M. L. V. Pitteway. An efficient structural technique for encoding ‘best-fit’ straight lines. The Computer Journal, 30(2):168–175, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/2/168.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

30/Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/169.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/170.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/171.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/172.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/173.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/174.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/175.tif.

Cr0xf0rd:1981:CZT

[Cr081] A. A. Cr0xf0rd. Correspondence: Os, zeros, thetas, and pis. The Computer Journal, 24(1):95, February 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_

24/Issue_01/tiff/95.tif. The use of the digit zero, rather than the letter oh, in the author’s name is intentional.

Clark:1983:GIP

[CR83] B. E. J. Clark and S. K. Robinson. A graphically interacting program monitor. The Computer Journal, 26(3):235–238, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/235.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/235.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/236.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/237.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/238.tif.

Challab:1987:BA

[CR87] D. J. Challab and J. D. Roberts. Buddy algorithms. The Computer Journal, 30(4):308–315, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/4/308.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/308.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/309.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/310.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_30/Issue_04/tiff/311.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/312.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/314.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/315.tif.

Craig:1986:ABS

[Cra86] I. D. Craig. The Ariadne-1 blackboard system. The Computer Journal, 29(3):235–240, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/3/235.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/236.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/237.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/238.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/239.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/240.tif.

Crowe:1980:PVO

[Cro80] M. K. Crowe. Porting virtual object files. The Computer Journal, 23(3):287–288, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/23/3/287.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/287.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/288.tif.

Crookes:1984:C

[Cro84] J. G. Crookes. Correspondence. The Computer Journal, 27(3): 286, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/286.1.full.pdf+html.

Cesarini:1985:DIR

[CS85] F. Cesarini and G. Soda. Design and implementation of a relational DBMS for microcomputers. The Computer Journal, 28(4): 391–397, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/28/4/391.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/391.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/392.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/393.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/394.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/395.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/396.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/397.tif.

Chen:1986:DAC

[CV86] W. C. Chen and J. S. Vitter. Deletion algorithms for coalesced hashing. The Computer Journal, 29(5):436–450, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/436.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/436.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/437.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/438.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/439.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/440.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/441.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/442.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/443.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/444.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/445.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/446.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/447.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/448.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/449.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/450.tif.

Corbett:1982:IPD

[CW82] Chris Corbett and Ian H. Witten. On the inclusion and placement of documentation graphics in computer typesetting. The Computer Journal, 25(2):272–277, May 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/272.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/272.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/273.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/274.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/275.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/276.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/277.tif.

Caride:1982:PBC

[CZ82] Anibal O. Caride and Susana I. Zanette. On the process of bisection to calculate the eigenvalues of quindiagonal matrices. The Computer Journal, 25(1):44, February 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/1/44.full.pdf+html.

Czech:1988:EID

[Cze88] Z. J. Czech. Efficient implementation of detection of undefined variables. The Computer Journal, 31(6):545–549, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/545.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/545.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/546.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/547.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/548.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/549.tif.

Delves:1981:SMS

[DAEH81] L. M. Delves, L. F. Abd-Elal, and J. A. Hendry. Set of modules for the solution of integral equations. The Computer Journal, 24(2): 184–190, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/184.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/184.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/185.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_02/tiff/186.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/187.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/188.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/189.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/190.tif.

Deen:1985:AGD

[DAODT85] S. M. Deen, R. R. Amin, G. O. Ofori-Dwumfuo, and M. C. Taylor. Architecture of a generalised distributed database system — PRECI. The Computer Journal, 28(3):282–290, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/282.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/282.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/284.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/285.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/286.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/287.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/288.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/289.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/290.tif.

Deen:1987:IPP

[DAT87] S. M. Deen, R. R. Amin, and M. C. Taylor. Implementation of a prototype for PRECI*. The Computer Journal, 30(2):157– 162, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/2/157.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/157.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/158.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/159.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/160.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/161.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/162.
tif.

Davenport:1981:DDD

[Dav81] R. A. Davenport. Design of distributed data base systems. The Computer Journal, 24(1):31–41, February 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/24/1/31.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/31.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/32.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/33.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/34.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/35.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/37.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
38.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/40.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
41.tif.

Davies:1982:NSF

[Dav82] D. Julian M. Davies. A note on sparsely filled dynamically allocated memory. The Computer Journal, 25(1):159, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/159.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/159.
tif.

Davis:1984:NDF

[Dav84a] R. H. Davis. A note on dynamic file organization models. The Computer Journal, 27(3):284, August 1984. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/284.tif.

Davis:1984:SNN

[Dav84b] R. H. Davis. Short note: a note on dynamic file organization models. The Computer Journal, 27(3):284, ???? 1984. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/27/3/284.
full.pdf+html.

Day:1980:BR

[Day80] A. Colin Day. Book review. The Computer Journal, 23(2):164, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/164.2.full.pdf+html.

Dutton:1981:NGC

[DB81] R. D. Dutton and R. C. Brigham. A new graph colouring algorithm. The Computer Journal, 24(1):85–86, February 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/24/1/85.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/85.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/86.
tif.

Dearle:1988:SBS

[DB88] A. Dearle and A. L. Brown. Safe browsing in a strongly typed persistent environment. The Computer Journal, 31(6):540–544, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/540.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/540.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/541.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/542.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/543.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/544.tif.

Djordjevic:1985:PLN

[DBH85] Jovan Djordjević, Mario Barbacci, and Brad Hosler. A PMS level notation for the description and simulation of digital systems. The Computer Journal, 28(4):357–365, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/357.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/357.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/358.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_04/tiff/359.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/360.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/361.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/362.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/363.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/364.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/365.tif. PMS.

Davis:1984:DFO

[DC84] R. H. Davis and P. Coumpas. A dynamic file organization model. The Computer Journal, 27(2):143–150, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/27/2/143.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/143.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/144.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/145.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/147.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/150.tif.

Dvorak:1987:SLT

[DD87] S. Dvorák and B. Durian. Stable linear time sublinear space merging. The Computer Journal, 30(4):372–375, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/372.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/373.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/374.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/375.tif.

Dvorak:1988:MDR

[DĎ88a] S. Dvořák and B. Ďurian. Merging by decomposition revisited. The Computer Journal, 31(6):553–556, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/31/6/553.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/553.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/554.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/555.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/556.tif.

Dvorak:1988:ULT

[DD88b] S. Dvorák and B. Durian. Unstable linear time O(1) space merging. The Computer Journal, 31(3):279–282, June 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/3/279.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/279.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/280.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/281.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/282.tif.

DeMorgan:1980:BR

[De 80] R. M. De Morgan. Book review. The Computer Journal, 23 (1):52, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/1/52.full.pdf+html.

Dowsing:1986:PBB

[DE86] R. D. Dowsing and R. Elliott. Programming a bounded buffer using the object and path expression constructs of Path Pascal. The Computer Journal, 29(5):423–429, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/423.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/423.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/424.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/425.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/

tiff/426.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/427.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/428.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/429.tif.

Dearnley:1982:SDM

[Dea82] P. A. Dearnley. Software development for microcomputer data processing systems. The Computer Journal, 25(2):253– 256, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/2/253.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/253.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/255.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/256.tif.

Deen:1980:CSG

[Dee80] S. M. Deen. A canonical schema for a generalised data model with local interfaces. The Computer Journal, 23(3):201–206, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/201.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/201.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/202.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/203.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/205.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/206.
tif.

Deen:1985:TPT

[Dee85] S. Misbah Deen. A tuple placement technique for databases. The Computer Journal, 28(3):291–297, July 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/291.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/291.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/292.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/

tiff/293.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/294.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/295.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/296.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/297.tif.

Denham:1980:BR

[Den80a] M. J. Denham. Book review. The Computer Journal, 23(1):89, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/89.full.pdf+html.

Denham:1980:BRB

[Den80b] M. J. Denham. Book review: Systems Decomposition, Optimisation and Control, by M. G. Singh and A. Titli. The Computer Journal, 23(1):89, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_23/Issue_01/tiff/89.tif.

Deogun:1983:SRT

[Deo83] J. S. Deogun. On scheduling with ready times to minimize mean flow time. The Computer Journal, 26(4):320–328, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/320.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/320.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/321.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/322.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/323.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/324.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/325.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/326.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/327.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/328.tif.

Duce:1987:FSC

[DF87] D. A. Duce and E. V. C. Fielding. Formal specification — a comparison of two techniques. The Computer Journal, 30(4):316– 327, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/316.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/316.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/317.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/318.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/319.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/320.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/321.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/322.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/323.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/324.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/325.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/326.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/327.tif.

Dixon:1985:SSS

[Dix85] F. J. Dixon. Simplifying screen specifications — the ‘Full Screen Manager’ interface and ‘Screen Form’ generating routines. The Computer Journal, 28(2):117–127, May 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/2/117.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/117.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/118.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/120.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/121.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/122.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/123.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/124.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/125.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/126.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/127.tif.

Dearnley:1983:FSP

[DM83] P. A. Dearnley and P. J. Mayhew. In favour of system prototypes and their integration into the systems development cycle. The Computer Journal, 26(1):36–42, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/26/1/36.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/37.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/38.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/39.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/40.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/41.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/42.tif.

Delves:1985:DAD

[DM85] L. M. Delves and S. C. Mawdsley. DAP-Algol: a development system for parallel algorithms. The Computer Journal, 28(2): 148–153, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/2/148.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/150.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/151.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/152.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/153.
tif.

Dillon:1988:RPV

[DMR88] A. Dillon, C. McKnight, and J. Richardson. Reading from paper versus reading from screen. The Computer Journal, 31(5):457– 464, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/5/457.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/457.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/458.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/459.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/460.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/461.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/462.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/463.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/464.tif.

Djokic:1989:SNF

[DMS^{+}89] B. Djokić, M. Miyakawa, S. Sekiguchi, I. Semba, and I. Stojmenović. Short note: a fast iterative algorithm for generating set partitions. The Computer Journal, 32(3):281–282, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/3/281.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/281.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/282.tif.

Deen:1981:DCD

[DNV81] S. M. Deen, D. Nikodem, and A. Vashishta. The design of a canonical database system (preci). The Computer Journal, 24(3):200– 209, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/200.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/200.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/201.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/202.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/203.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/204.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/205.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/206.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/207.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/208.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/209.tif.

Dobosiewicz:1984:RSL

[Dob84] Wlodzimierz Dobosiewicz. Replacement selection in 3-level memories. The Computer Journal, 27(4):334–339, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/334.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/334.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/336.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/338.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/339.
tif.

Dawson:1988:ERD

[DP88] Kathryn S. Dawson and Lorraine M. Purgailis Parker. From entity-relationship diagrams to fourth normal form: a pictorial aid to analysis. The Computer Journal, 31(3):258–268, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/258.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/258.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/259.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/260.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/261.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/262.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/263.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/265.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/266.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/267.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/268.
tif.

Dixon:1989:TPO

[DPSW89] G. N. Dixon, G. D. Parrington, S. K. Shrivastava, and S. M. Wheater. The treatment of persistent objects in Arjuna. The Computer Journal, 32(4):323–332, August 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/4/323.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/323.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_32/Issue_04/tiff/324.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/325.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/326.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/327.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/329.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/330.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/331.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/332.
tif.

Dromey:1987:DSA

[Dro87] R. G. Dromey. Derivation of sorting algorithms from a specification. The Computer Journal, 30(6):512–518, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/512.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/512.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/513.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/514.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/515.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/516.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/517.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/518.tif.

deSmith:1980:PPA

[dS80] Michael de Smith. Points, polygons, and areas. The Computer Journal, 23(2):189, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_

02/tiff/189.tif. See correspondence [Kio81].

Deen:1988:DDD

[DTIR88] S. M. Deen, M. C. Taylor, P. A. Ingram, and K. W. Rayner. A distributed directory database system for telecommunications. The Computer Journal, 31(2):175–181, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.

org/content/31/2/175.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/175.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/176.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/177.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/178.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/179.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/180.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/181.tif.

DuFeu:1986:C

[Du 86] D. Du Feu. Correspondence. The Computer Journal, 29(5):478, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/5/478.2.full.pdf+html.

Dunham:1982:NPP

[Dun82] Charles B. Dunham. The necessity of publishing programs. The Computer Journal, 25(1):61–62, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/61.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/62.tif.

Davis:1988:MSB

[DW88] Martin Davis and Elaine Weyuker. Metric space-based test-data adequacy criteria. The Computer Journal, 31(1):17–24, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
31/1/17.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/17.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
18.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
21.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/23.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
24.tif.

Dasarathy:1980:TOT

[DY80] B. Dasarathy and Cheng Yang. A transformation on ordered trees. The Computer Journal, 23(2):161–164, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/23/2/161.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/161.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/164.tif.

Evans:1983:ECB

[EA83] D. J. Evans and F. Abdollahzadeh. Efficient construction of balanced binary trees. The Computer Journal, 26(3):193–195, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/193.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/193.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/195.tif.

Earnshaw:1980:LTI

[Ear80] R. A. Earnshaw. Line tracking for incremental plotters. The Computer Journal, 23(1):46–52, February 1980. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/23/1/46.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/46.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/48.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/50.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/51.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/52.tif.

Easteal:1980:BR

[Eas80] C. D. Easteal. Book review. The Computer Journal, 23(2):114, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/114.2.full.pdf+html.

Edupuganty:1989:TLG

[EB89] B. Edupuganty and B. R. Bryant. Two-level grammar as a functional programming language. The Computer Journal, 32(1):36– 44, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/1/36.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_32/Issue_01/tiff/36.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_32/
Issue_01/tiff/37.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/38.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
39.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/40.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/41.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
42.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/43.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/44.tif.

Ein-Dor:1986:AAS

[EDS86] P. Ein-Dor and E. Segev. Attitudes, association and success of MIS: some empirical results from research in the context of a business game. The Computer Journal, 29(3):212– 221, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/212.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/212.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/213.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/214.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/215.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/216.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/217.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/218.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/219.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/220.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/221.tif.

Edwards:1981:BR

[Edw81] D. B. G. Edwards. Book review. The Computer Journal, 24(3): 248, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/248.2.full.pdf+html.

Eason:1988:SRD

[EH88] K. D. Eason and Susan Harker. The supplier’s role in the design of products for organisations. The Computer Journal, 31(5):426– 430, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/5/426.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/426.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/427.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/428.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/429.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/430.tif.

Edelsbrunner:1986:RPL

[EHH86] H. Edelsbrunner, G. Haring, and David Hilbert. Rectangular point location in D dimensions with applications. The Computer Journal, 29(1):76–82, February 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/29/1/76.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/76.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/77.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/78.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/80.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/81.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/82.tif.

El-Hamdouchi:1989:CHA

[EHW89] A. El-Hamdouchi and P. Willett. Comparison of hierarchic agglomerative clustering methods for document retrieval. The Computer Journal, 32(3):220–227, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/3/220.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/221.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/222.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/223.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/224.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/225.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/226.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/227.tif.

Er:1982:TPC

[EL82] M. C. Er and B. G. T. Lowden. The theory and practice of constructing an optimal polyphase sort. The Computer Journal, 25(1):93–101, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/93.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_25/Issue_01/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_25/
Issue_01/tiff/93.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/94.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
95.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/96.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/97.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
98.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/99.tif.

Ellis:1989:SNV

[EL89] J. A. Ellis and P. M. Lepolesa. Short note: a Las Vegas graph colouring algorithm. The Computer Journal, 32(5):474– 476, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/5/474.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/474.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

32/Issue_05/tiff/475.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/476.tif.

Elston:1986:C

[Els86] R. G. Elston. Correspondence. The Computer Journal, 29(3): 285–286, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/285.2.full.pdf+html.

Ekanadham:1981:MN

[EM81a] K. Ekanadham and A. Mahjoub. Microcomputer networks. The Computer Journal, 24(1):17–24, February 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/24/1/17.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/17.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/19.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/21.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/23.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
24.tif.

Ewusi-Mensah:1981:CNU

[EM81b] K. Ewusi-Mensah. Computer network usage control through peak load pricing. The Computer Journal, 24(1):71–77, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/1/71.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/71.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
72.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/73.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/74.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
75.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/76.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/77.tif.

Emery:1980:BRB

[Eme80] G. Emery. Book reviews: Associative networks, edited by N. V. Findler, 1979; The MU5 Computer System, by Derrick Morris and Rolland N. Ibbett, 1979; Cobol: An Introduction to Structured Logic and Modular Program Design, by William Davis and Richard Fisher, 1979. The Computer Journal, 23(2):191, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/191.tif.

Ernvall:1982:AUR

[EN82] Jarmo Ernvall and Olli Nevalainen. An algorithm for unbiased random sampling. The Computer Journal, 25(1):45–47, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/1/45.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
46.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/47.tif.

English:1981:C

[Eng81] J. English. Correspondence. The Computer Journal, 24(4):377, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/377.4.full.pdf+html.

Eidelson:1980:IBI

[ER80] A. F. Eidelson and I. J. Robinson. Implementation of BEDSOCS: an interactive simulation language. The Computer Journal, 23(3): 233–236, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/233.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/233.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/234.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/235.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/

tiff/236.tif. See correspondence [SB81].

Er:1982:NCP

[Er82a] M. C. Er. A note on computing precedence functions. The Computer Journal, 25(3):397–398, August 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/3/397.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/397.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/398.tif.

Er:1982:RAT

[Er82b] M. C. Er. A representation approach to the Tower of Hanoi problem. The Computer Journal, 25(4):442–447, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/25/4/442.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/442.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/443.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/444.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/445.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/446.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/447.
tif.

Er:1983:FAC

[Er83a] M. C. Er. Fast algorithm for computing order-K Fibonacci numbers. The Computer Journal, 26(3):224–227, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/26/3/224.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/224.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/225.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/226.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/227.tif.

Er:1983:NGW

[Er83b] M. C. Er. A note on generating well-formed parenthesis strings lexicographically. The Computer Journal, 26(3):205– 207, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/205.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/205.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

26/Issue_03/tiff/206.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/207.tif.

Er:1983:PCA

[Er83c] M. C. Er. A parallel computation approach to topological sorting. The Computer Journal, 26(4):293–295, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/293.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/293.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/294.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/295.tif.

Er:1984:CTHa

[Er84a] M. C. Er. The colour Towers of Hanoi: a generalization. The Computer Journal, 27(1):80–82, February 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/27/1/80.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/82.tif.

Er:1984:C

[Er84b] M. C. Er. Correspondence. The Computer Journal, 27(3):285, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/3/285.1.full.pdf+html.

Er:1984:CTHb

[Er84c] M. C. Er. The cyclic Towers of Hanoi: a representation approach. The Computer Journal, 27(2):171–175, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/27/2/171.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/171.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/172.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/173.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/

tiff/174.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/175.tif.

Er:1984:GCT

[Er84d] M. C. Er. The generalized colour Towers of Hanoi: an iterative algorithm. The Computer Journal, 27(3):278–282, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/278.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/278.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/279.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/280.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/281.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/282.tif.

Er:1985:C

[Er85a] M. C. Er. Correspondence. The Computer Journal, 28(5):543– 544, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/543.2.full.pdf+html.

Er:1985:EOT

[Er85b] M. C. Er. Enumerating ordered trees lexicographically. The Computer Journal, 28(5):538–542, November 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/5/538.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/538.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/539.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/540.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/541.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/542.
tif.

Elloy:1986:ELC

[ER86a] J. P. Elloy and O. Roux. Electre: a language for control structuring in real time. The Computer Journal, 29(3):229– 234, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/229.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/229.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/231.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/232.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/233.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/234.
tif.

Er:1986:UTI

[Er86b] M. C. Er. The use of termination indicators in computer programming. The Computer Journal, 29(5):430–433, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/430.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/430.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/431.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/432.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/433.tif.

Er:1987:EIP

[Er87a] M. C. Er. An efficient implementation of permutation backtracking in lexicographic order. The Computer Journal, 30(3):282, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_03/tiff/282.tif.

Er:1987:LLR

[Er87b] M. C. Er. Lexicographic listing and ranking of T -ary trees. The Computer Journal, 30(6):569–572, December 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/6/569.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/569.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/570.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/571.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/572.tif.

Er:1987:SNE

[Er87c] M. C. Er. Short note: an efficient implementation of permutation backtracking in lexicographic order. The Computer Journal, 30 (3):282, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/3/282.full.pdf+html.

Er:1988:FAG

[Er88a] M. C. Er. A fast algorithm for generating set partitions. The Computer Journal, 31(3):283–284, June 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/3/283.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/284.tif.

Er:1988:SAG

[Er88b] M. C. Er. A simple algorithm for generating non-regular trees in lexicographic order. The Computer Journal, 31(1):61–64, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
31/1/61.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/61.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
62.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/63.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/64.tif.

Er:1988:SRA

[Er88c] M. C. Er. A smooth reshuffling algorithm for solving the Bulterman’s problem. The Computer Journal, 31(6):557–560, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/557.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/557.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/558.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/559.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/560.tif.

Er:1989:CAP

[Er89a] M. C. Er. Classes of admissible permutations that are generatable by depth-first traversals of ordered trees. The Computer Journal, 32(1):76–85, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/76.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/76.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/77.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/78.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/80.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/81.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/82.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
83.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/84.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/85.tif.

Er:1989:NAG

[Er89b] M. C. Er. A new algorithm for generating binary trees using rotations. The Computer Journal, 32(5):470–473, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/5/470.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/470.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/471.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/472.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/473.tif.

Erkio:1984:WCP

[Erk84] Hannu Erkiö. The worst case permutation for median-of-three quicksort. The Computer Journal, 27(3):276–277, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/276.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/276.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/277.tif.

Elliman:1981:PTM

[ES81] A. D. Elliman and A. R. Sudhindra. Performances of two microprocessor-based multiplexor techniques. The Computer Journal, 24(4):343–346, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/4/343.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/343.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/344.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/345.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/346.tif.

Egyhazy:1988:QPA

[ET88] C. Egyhazy and K. Triantis. A query processing algorithm for distributed relational database systems. The Computer Journal, 31(1):34–40, February 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/31/1/34.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/34.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_31/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_
01/tiff/36.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/38.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/40.tif.

Evans:1981:BR

[Eva81] R. E. Evans. Book review. The Computer Journal, 24(3):234, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/234.1.full.pdf+html.

Evans:1983:SIG

[Eva83] S. Evans. Spss introductory guide. The Computer Journal, 26(4): 382, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/382.full.pdf+html.

Evans:1980:ADP

[EW80] D. J. Evans and Shirley A. Williams. Analysis and detection of parallel processable code. The Computer Journal, 23(1):66–72, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/66.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/67.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/68.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/69.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/70.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/71.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
72.tif.

Episkopou:1986:TFC

[EWH86] D. M. Episkopou and A. T. Wood-Harper. Towards a framework to choose appropriate IS approaches. The Computer Journal, 29(3): 222–228, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/222.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/222.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/224.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/226.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/227.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/228.tif.

Franklin:1988:AGP

[FA88] W. R. Franklin and V. Akman. Adaptive grid for polyhedral visibility in object space: an implementation. The Computer Journal, 31(1):56–60, February 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/31/1/56.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/56.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_31/Issue_01/tiff/57.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_
01/tiff/58.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/59.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/60.
tif.

Falla:1981:GSE

[Fal81] M. E. Falla. The gamma software engineering system. The Computer Journal, 24(3):235–242, August 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/3/235.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/236.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/237.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/238.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/239.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/240.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/241.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/242.tif.

Foster:1987:RC

[FC87] J. M. Foster and I. F. Currie. Remote capabilities. The Computer Journal, 30(5):451–457, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/5/451.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/451.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/452.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/453.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/454.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/455.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/456.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/457.tif.

Fang:1988:CPC

[FD88] Z. Fang and J. S. Deogun. Closure properties of certain classes of languages under generalized morphic replication. The Computer Journal, 31(4):325–329, August 1988. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/4/325.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/325.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/326.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/327.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/328.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/329.
tif.

Fellget:1980:BRB

[FE80] P. B. Fellget and C. D. Easteal. Book reviews: The Phenomenon of Science: A Cybernetic Approach to Human Evolution, by V. F. Turchin, 1977; Business Systems Handbook, by Robert W. Gilmour, 1979. The Computer Journal, 23(2):114, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/114.tif.

Fellgett:1980:BR

[Fel80] P. B. Fellgett. Book review. The Computer Journal, 23(2):114, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/114.1.full.pdf+html.

Fellgett:1984:C

[Fel84] P. Fellgett. Correspondence. The Computer Journal, 27(4):378– 379, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/378.2.full.pdf+html.

Fenwick:1987:FCA

[Fen87] P. M. Fenwick. A fast-carry adder with CMOS transmission gates. The Computer Journal, 30(1):77–79, February 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/30/1/77.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_30/Issue_01/tiff/78.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_
01/tiff/79.tif.

Ferraby:1986:C

[Fer86] E. F. Ferraby. Correspondence. The Computer Journal, 29(2): 190, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/190.2.full.pdf+html.

Findlay:1983:C

[Fin83] Bill Findlay. Correspondence. The Computer Journal, 26(2):189, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/2/189.2.full.pdf+html.

Fisher:1980:BR

[Fis80] D. L. Fisher. Book review. The Computer Journal, 23(1):84, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/84.4.full.pdf+html.

Fisher:1983:SAF

[Fis83] D. L. Fisher. A structured approach to Fortran 77 programming. The Computer Journal, 26(4):381a–381, November 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/4/381-
a.full.pdf+html.

Ford:1980:BRB

[FJA80] Brian Ford, David Jacobs, and A. M. Addyman. Book reviews: Workshop on Reliable Software: Applied Computer Science, by Peter Raulefs, 1979; Reliable Software Through Composite Design, by G. J. Myers, 1975; Fortran, PL/1 and the Algols, by Brian Meek, 1978. The Computer Journal, 23(2):146, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/146.tif.

Fucik:1986:HPC

[FK86] J. Fucík and J. Král. The hierarchy of program control structures. The Computer Journal, 29(1):24–32, February 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/29/1/24.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/25.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/26.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/27.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/28.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/30.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
31.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/32.tif.

Freisleben:1989:PS

[FK89] B. Freisleben and J. L. Keedy. Priority semaphores. The Computer Journal, 32(1):24–28, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/24.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/26.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/27.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/28.
tif.

Fenner:1980:BTR

[FL80] T. I. Fenner and G. Loizou. A binary tree representation and related algorithms for generating integer partitions. The Computer Journal, 23(4):332–337, November 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_04/tiff/332.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_04/tiff/333.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_04/tiff/334.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/335.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/336.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/337.tif.

Flynn:1985:MCS

[FL85] D. J. Flynn and A. H. F. Laender. Mapping from a conceptual schema to a target internal schema. The Computer Journal, 28(5):508–517, November 1985. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/5/508.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/508.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/509.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/510.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/511.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/512.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/513.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/514.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/515.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/516.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/517.
tif.

Frost:1989:CNL

[FL89]  R. Frost and J. Launchbury.  Constructing natural lan-

guage interpreters in a lazy functional language. The Computer Journal, 32(2):108–121, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/2/108.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/108.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/109.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/110.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/111.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/112.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/113.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/114.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/115.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/116.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/117.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/118.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/119.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/120.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/121.tif.

Fletcher:1982:CMD

[Fle82] R. Fletcher. Combinatorial methods of discrete programming. The Computer Journal, 25(1):160b–160, February 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/1/160-
b.full.pdf+html.

Florentin:1980:SCC

[Flo80] J. J. Florentin. Speech communication with computers. The Computer Journal, 23(3):255b–255, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/3/255-
b.full.pdf+html.

Fraenkel:1983:CCP

[FM83] Aviezri S. Fraenkel and Moshe Mor. Combinatorial compression and partitioning of large dictionaries. The Computer Journal, 26(4):336–343, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/4/336.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/336.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/337.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/338.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/339.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/343.tif.

Furtado:1985:IAF

[FM85] A. L. Furtado and T. S. E. Maibaum. An informal approach to formal (algebraic) specifications. The Computer Journal, 28(1):59–67, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/1/59.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_01/tiff/59.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/60.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/61.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/62.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/63.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/64.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
65.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/66.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/67.tif.

Fabbrini:1986:AQ

[FM86a] F. Fabbrini and C. Montani. Autumnal quadtrees. The Computer Journal, 29(5):472–474, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/5/472.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/472.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/473.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/474.tif.

Feldman:1986:EMC

[FM86b] P. Feldman and D. Miller. Entity model clustering: Structuring a data model by abstraction. The Computer Journal, 29(4):348– 360, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/348.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/348.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/350.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/352.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/353.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/354.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/355.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/356.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/357.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/358.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/359.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/360.tif.

Foo:1986:DIP

[Foo86] N. Y. Foo. Dewey indexing of Prolog traces. The Computer Journal, 29(1):17–19, February 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/29/1/17.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/17.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/19.tif.

Ford:1980:BR

[For80] B. Ford. Book reviews. The Computer Journal, 23(2):146, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
23/2/146.1.full.pdf+html.

Foster:1986:VMA

[Fos86] J. M. Foster. Validating microcode algebraically. The Computer Journal, 29(5):416–422, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/5/416.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/416.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/417.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/418.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/419.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/420.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/421.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/422.tif.

Fox:1987:ODJ

[Fox87] Leslie Fox. Obituary: Dr. J. H. Wilkinson, F.R.S. The Computer Journal, 30(1):1, 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/1.tif.

Frost:1982:SNB

[FP82] R. A. Frost and M. M. Peterson. A short note on binary search trees. The Computer Journal, 25(1):158, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/158.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/158.
tif.

Feng:1980:SNS

[FR80] D. Y. Feng and R. F. Riesenfeld. Some new surface forms for computer aided geometric design. The Computer Journal, 23(4):324– 331, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/324.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/325.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/326.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/327.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/329.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/330.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/331.tif.

Faidhi:1987:PEL

[FR87] J. A. W. Faidhi and S. K. Robinson. Programmer experience level indicators. The Computer Journal, 30(1):52–62, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
30/1/52.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/52.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
53.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/55.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
56.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/57.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/58.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
59.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/60.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/61.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
62.tif.

Freeman:1981:BRa

[Fre81a] P. R. Freeman. Book review. The Computer Journal, 24(1):96, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/96.2.full.pdf+html.

Freeman:1981:BRb

[Fre81b] P. R. Freeman. Book review. The Computer Journal, 24(2):172, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/172.1.full.pdf+html.

Freeman:1982:BPS

[Fre82] P. R. Freeman. Basic-pack statistics programs for small computers. The Computer Journal, 25(3):396, August 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/3/396.
full.pdf+html.

Frost:1981:ADI

[Fro81] R. A. Frost. Algorithm 112: Dumping the index of a dynamic hash table. The Computer Journal, 24(4):383–384, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_24/Issue_04/tiff/383.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_24/Issue_04/tiff/384.tif.

Frost:1982:BRS

[Fro82] R. A. Frost. Binary-relational storage structures. The Computer Journal, 25(3):358–367, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/3/358.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/358.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/359.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/360.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/361.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/362.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/363.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/364.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/365.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/366.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/367.

tif. Discusses the dynamic hashing scheme used by ASDAS, under development at Strathclyde University.

Frost:1983:SYA

[Fro83] R. A. Frost. SCHEMAL: Yet another conceptual schema definition language. The Computer Journal, 26(3):228–234, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/228.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/228.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/229.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/230.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/231.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/232.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/233.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/234.tif.

Frost:1985:USC

[Fro85] R. A. Frost. Using semantic concepts to characterise various knowledge representation formalisms: a method of facilitating the interface of knowledge base system components. The Computer Journal, 28(2):112–116, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/2/112.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/112.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/113.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/114.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/115.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/116.
tif.

Frost:1986:SNI

[Fro86] R. A. Frost. Short notes: Improving output from research (in the domain of knowledge base systems). The Computer Journal, 29(6): 572–573, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/6/572.full.pdf+html.

Froggatt:1987:C

[Fro87] T. J. Froggatt. Correspondence. The Computer Journal, 30(3): 285–286, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/3/285.2.full.pdf+html.

Forrest:1987:INN

[FRS^{+}87] B. M. Forrest, D. Roweth, N. Stroud, D. J. Wallace, and G. V. Wilson. Implementing neural network models on parallel computers. The Computer Journal, 30(5):413–419, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/413.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/413.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/414.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/415.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/416.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/417.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/418.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/419.tif.

Foxley:1981:VAM

[FS81] E. Foxley and O. Salman. Validation of an analytic model of computer performance. The Computer Journal, 24(4):347–352, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/347.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/347.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/348.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/349.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/350.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/351.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/352.
tif.

Firby:1987:CMS

[FS87] P. A. Firby and D. J. Stone. Colour manipulation of superposed families of curves. The Computer Journal, 30(4):349– 354, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/349.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/349.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/350.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/351.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/352.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/353.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/354.
tif.

Fitzgerald:1985:FAC

[FSW85] G. Fitzgerald, N. Stokes, and J. R. G. Wood. Feature analysis of contemporary information systems methodologies. The Computer Journal, 28(3):223–230, July 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/223.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/224.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/226.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/227.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/228.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/229.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/230.tif.

Farlie:1982:BRW

[FSWL82] D. J. G. Farlie, Mildred L. G. Shaw, B. A. Wichmann, and S. H. Lavington. Book reviews: W. M. Brelsford and R. A. Relles, STATLILB: A Statistical Computing Laboratory, 1981; N. Rushby (ed.), Selected Readings in Computer-Based Learning, 1981; R. D. Tennent, Principles of Programming Languages, 1981; D. Hutchison, Fundamentals of Computer Logic, 1981. The Computer Journal, 25(2):287, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_25/Issue_02/tiff/287.tif.

Fung:1984:RMB

[Fun84] K. T. Fung. A reorganization model based on the database entropy concept. The Computer Journal, 27(1):67–71, February 1984. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/27/1/67.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/67.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/68.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/70.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
71.tif.

Fung:1985:EAS

[Fun85] K. T. Fung. An entropy approach to solving some network reliability problems. The Computer Journal, 28(4):353–356, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/353.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/353.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/354.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/355.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/356.tif.

Furuta:1989:OBT

[Fur89] R. Furuta. An object-based taxonomy for abstract structure in document models. The Computer Journal, 32(6):494–504, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/494.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/494.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/495.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/496.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/497.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/498.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/499.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/500.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/501.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/502.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/503.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/504.
tif.

Frost:1983:STA

[FW83] R. A. Frost and S. Whittaker. A step towards the automatic maintenance of the semantic integrity of databases. The Computer Journal, 26(2):124–133, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/2/124.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/124.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/125.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/127.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/128.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/129.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/130.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/131.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/132.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/133.
tif.

Fenton:1986:AAS

[FW86] N. E. Fenton and R. W. Whitty. Axiomatic approach to software metrication through program decomposition. The Computer Journal, 29(4):330–339, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/4/330.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/330.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/331.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/332.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/333.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/334.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/336.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/338.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/339.
tif.

Gallard:1987:EDP

[Gal87] R. H. Gallard. An extension in the definition of a Petri net execution. The Computer Journal, 30(1):16–19, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/30/1/16.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/16.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_30/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_
01/tiff/18.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/19.tif.

Gupta:1983:PGP

[GB83] Phalguni Gupta and G. P. Bhattacharjee. Parallel generation of permutations. The Computer Journal, 26(2):97–105, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
26/2/97.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/100.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/102.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/105.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/97.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/99.
tif.

Gault:1987:SNF

[GC87] D. Gault and M. Clint. Short notes: a fast algorithm for the Towers of Hanoi problem. The Computer Journal, 30(4):376– 378, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/376.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/376.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/377.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/378.tif.

Gallimore:1989:UOL

[GCS89] R. M. Gallimore, D. Coleman, and V. Stavridou. UMIST OBJ: a language for executable program specifications. The Computer Journal, 32(5):413–421, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/5/413.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/413.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/414.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/415.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/416.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/417.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/418.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/419.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/420.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/421.tif.

Gray:1985:PRI

[GE85] P. M. D. Gray and P. Esslemont. The performance of a relational interface to a Codasyl database. The Computer Journal, 28(5):501–507, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/5/501.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/501.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/502.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/503.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/504.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/505.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/506.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/507.tif.

Gearing:1980:NAS

[Gea80] H. W. G. Gearing. New approaches to systems analysis and design. welcome. The Computer Journal, 23(1):3, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/1/3.1.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/3.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_01/tiff/4.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/5.tif.

Gehani:1983:HLF

[Geh83] N. H. Gehani. High level form definition in office information systems. The Computer Journal, 26(1):52–59, February 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/1/52.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/53.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/55.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
56.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/57.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/58.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
59.tif.

George:1981:BR

[Geo81] F. H. George. Book review. The Computer Journal, 24(4):346, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/346.2.full.pdf+html.

Gwei:1987:FSI

[GF87] G. M. Gwei and E. Foxley. A flexible synonym interface with application examples in CAL and help environments. The Computer Journal, 30(6):551–557, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/6/551.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/551.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/552.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/553.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/554.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/555.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/556.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/557.tif.

Georgeff:1986:RNU

[GG86] M. P. Georgeff and G. K. Gupta. Resources needed for a university computer science department. The Computer Journal, 29(1):90– 95, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/1/90.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_29/Issue_01/tiff/90.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_29/
Issue_01/tiff/91.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/92.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
93.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/94.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/95.tif.

Godfrey:1987:PCS

[GGH^{+}87] D. Godfrey, S. Gong, R. Hart, N. Koorland, and S. Smit. Project ‘Clear Skies’: teaching computer science by computer based training and electronic messaging in China. The Computer Journal, 30(5):469–474, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/5/469.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/469.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/470.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/471.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/472.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/473.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/474.tif.

Godwin:1989:CJD

[GGS89] A. N. Godwin, M. B. Gore, and D. W. Salt. A comparison of JSD and DFD as descriptive tools. The Computer Journal, 32(3): 202–211, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/3/202.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/202.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/203.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/204.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/205.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/206.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/207.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/208.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/209.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/210.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/211.tif.

Gibbins:1988:VAP

[Gib88] P. F. Gibbins. VDM: axiomatising its propositional logic. The Computer Journal, 31(6):510–516, December 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/6/510.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/510.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/511.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/512.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/513.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/514.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/515.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/516.tif.

Giles:1982:BRP

[Gil82] P. Giles. Book review: P. Calingaert, Operating System Elements — a User Perspective. The Computer Journal, 25(4):422b–422, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/422-b.full.pdf+html.

Giles:1983:BR

[Gil83a] P. Giles. Book review. The Computer Journal, 26(2):190, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
26/2/190.2.full.pdf+html.

Giles:1983:OSE

[Gil83b] P. Giles. Operating system elements — a user perspective. The Computer Journal, 26(4):383–d–383, November 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/4/383-
d.full.pdf+html.

Giles:1983:CNS

[Gil83c] Philip Giles. Computer networks and simulation 11. The Computer Journal, 26(4):381–d–382, November 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/4/381-
d.full.pdf+html.

Gale:1984:PIC

[GK84] William A. Gale and Roger Koenker. Pricing interactive computer services. The Computer Journal, 27(1):8–17, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/1/8.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/13.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/14.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/15.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
16.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/17.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
9.tif.

Ghosh:1983:PCG

[GM83] P. K. Ghosh and S. P. Mudur. Parametric curves for graphic design systems. The Computer Journal, 26(4):312–319, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/312.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/312.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/314.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/315.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/316.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/317.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/318.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/319.tif.

Griswold:1983:PDC

[GN83] R. E. Griswold and Michael Novak. Programmer-defined control operations. The Computer Journal, 26(2):175–183, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/175.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/175.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/176.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/177.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/178.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/179.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/180.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/181.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/182.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/183.tif.

Goyal:1988:DMI

[GN88] P. Goyal and T. S. Narayanan. Dictionary machine with improved performance. The Computer Journal, 31(6):490–495, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/31/6/490.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/490.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/491.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/492.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/493.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/494.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/495.
tif.

Goller:1982:TLU

[Gol82] N. E. Goller. Triodic logic and its use in structured program design. The Computer Journal, 25(2):218–226, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/25/2/218.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/218.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/219.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/220.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/221.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/222.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/223.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/224.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/225.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/226.tif.

Goller:1985:HDS

[Gol85] N. E. Goller. Hybrid data structures defined by indirection. The Computer Journal, 28(1):44–53, February 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/28/1/44.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/46.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/48.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

28/Issue_01/tiff/49.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/50.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
51.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/52.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/53.tif.

Gomaa:1981:HSR

[Gom81] H. Gomaa. A hybrid simulation/regression model of a virtual storage system. The Computer Journal, 24(4):358–366, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/358.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/358.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/359.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/360.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/361.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/362.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/365.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/366.tif.

Goodwin:1980:PNU

[Goo80] D. T. Goodwin. Partial non-underflow and non-overflow of an arithmetic stack. The Computer Journal, 23(2):153–160, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/153.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/153.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/154.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/155.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/156.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/157.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/158.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/159.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/160.tif.

Goodwin:1985:CSN

[Goo85a] D. T. Goodwin. The computability of stack non-underflow. The Computer Journal, 28(3):257–263, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/257.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/257.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/258.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/259.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/260.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/261.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/262.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/263.tif.

Goodwin:1985:FSS

[Goo85b] D. T. Goodwin. Further study of a stack-length model. The Computer Journal, 28(1):9–16, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/28/1/9.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/10.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/11.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/12.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/13.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
14.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/15.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/16.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
9.tif.

Gough:1980:IEO

[Gou80] T. G. Gough. Introducing the electronic office. The Computer Journal, 23(3):200, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/200.full.pdf+html.

Gough:1981:BR

[Gou81] T. G. Gough. Book review. The Computer Journal, 24(1):28, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/28.full.pdf+html.

Gonnet:1983:HRB

[GOW83] Gaston H. Gonnet, Henk J. Olivié, and Derick Wood. Heightratio-balanced trees. The Computer Journal, 26(2):106–108, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/106.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/106.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/107.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/108.tif.

Garside:1980:API

[GP80] G. R. Garside and P. E. Pintelas. An ALGOL 68 package for implementing graph algorithms. The Computer Journal, 23(3):237– 242, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/237.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/237.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/239.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/241.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/242.
tif.

Graybeal:1981:DCE

[GP81] W. T. Graybeal and U. W. Pooch. The design of calibration experiments for synthetic jobs. The Computer Journal, 24(2): 132–138, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/132.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/132.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/133.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/134.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/135.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/136.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/137.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/138.tif.

Gopalakrishnan:1988:MAC

[GP88] M. Gopalakrishnan and L. M. Patnaik. Medium access control schemes for local area networks with multiple priority function. The Computer Journal, 31(3):209–219, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/31/3/209.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/209.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/210.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/211.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/212.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/213.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/214.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/215.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/216.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/218.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/219.
tif.

Gray:1980:BR

[Gra80] P. M. D. Gray. Book review. The Computer Journal, 23(1):60, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/60.1.full.pdf+html.

Gray:1987:PVC

[Gra87] D. Gray. A pedagogical verification condition generator. The Computer Journal, 30(3):239–248, June 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/3/239.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/239.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/240.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/241.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_30/Issue_03/tiff/242.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/243.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/244.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/245.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/246.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/247.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/248.
tif.

Griswold:1980:UCS

[Gri80] R. E. Griswold. The use of character sets and character mappings in Icon. The Computer Journal, 23(2):107– 114, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/107.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/107.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/108.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/109.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/110.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/111.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/112.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/113.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/114.tif.

Griswold:1988:PG

[Gri88] R. E. Griswold. Programming with generators. The Computer Journal, 31(3):220–228, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/3/220.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/221.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/222.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/223.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/224.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/225.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/226.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/227.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/228.tif.

Grover:1982:CP

[Gro82] Derrick Grover. Cryptography: a primer. The Computer Journal, 25(3):400c–400, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/3/400-c.full.pdf+html.

Grosch:1984:C

[Gro84] H. R. J. Grosch. Correspondence. The Computer Journal, 27 (4):377, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/377.1.full.pdf+html.

Guariso:1988:SBE

[GW88] G. Guariso and H. Werthner. A software base for environmental studies. The Computer Journal, 31(6):550–552, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/550.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/550.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/551.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/552.tif.

Haine:1983:ATI

[Hai83] Peter Haine. Automated tools for information systems design. The Computer Journal, 26(4):383b–383, November 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/4/383-
b.full.pdf+html.

Hammersley:1980:CSA

[Ham80a] P. Hammersley. Corporate Systems analysis — a personal editorial. The Computer Journal, 23(1):2, February 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/1/2.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/2.tif.

Hammersley:1980:ICT

[Ham80b] P. Hammersley. The impact of changing technology. The Computer Journal, 23(1):7–9, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/1/7.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/7.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/8.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_01/tiff/9.tif.

Hammersley:1981:E

[Ham81a] P. Hammersley. Editorial. The Computer Journal, 24(1):1, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/1.tif.

Hammersley:1981:IMS

[Ham81b] P. Hammersley. The impact of microcomputer systems on commercial data processing. The Computer Journal, 24(1):14–16, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/14.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/16.tif.

Hammersley:1982:E

[Ham82] P. Hammersley. Editorial. The Computer Journal, 25(1):1, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/25/1/1.full.pdf+html.

Hammersley:1984:E

[Ham84] P. Hammersley. Editorial. The Computer Journal, 27(1): 1–2, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_27/Issue_01/tiff/1.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_27/
Issue_01/tiff/2.tif.

Hammersley:1989:EII

[Ham89] P. Hammersley. Editorial — information and information systems. The Computer Journal, 32(3):193, ???? 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/32/3/193.
full.pdf+html.

Hart:1980:JBA

[Har80] Deborah Hart. Joint BCS ACM symposium on Research and Development in Information Retrieval . The Computer Journal, 23 (2):131, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/131.tif.

Harrington:1981:SEC

[Har81a] Steven J. Harrington. Space efficient copying storage recovery. The Computer Journal, 24(4):316–319, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/24/4/316.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/316.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/317.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/318.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/319.tif.

Harris:1981:C

[Har81b] D. Harris. Correspondence. The Computer Journal, 24(3):270, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/270.1.full.pdf+html.

Hartley:1981:C

[Har81c] D. F. Hartley. Correspondance. The Computer Journal, 24(2): 192, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/192.2.full.pdf+html.

Harrison:1982:ESM

[Har82a] P. G. Harrison. Efficient storage management for functional languages. The Computer Journal, 25(2):264–271, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/25/2/264.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/264.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/265.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/266.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/267.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/268.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/269.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/270.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/271.tif.

Hart:1982:IBI

[Har82b] Roger Hart. Implementation of BEDSOCS: an interactive simulation language. The Computer Journal, 25(3):399, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/3/399.full.pdf+html.

Harper:1983:C

[Har83] D. Harper. Correspondence. The Computer Journal, 26(1):95, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/1/95.1.full.pdf+html.

Harland:1984:UDT

[Har84a] David M. Harland. User-defined types in a polymorphic language. The Computer Journal, 27(1):47–56, February 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/27/1/47.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/48.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/49.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/51.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/52.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/53.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
54.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/55.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/56.tif.

Harrison:1984:DCT

[Har84b] P. G. Harrison. The distribution of cycle times in tree-like networks of queues. The Computer Journal, 27(1):27–36, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
27/1/27.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/27.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
28.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/30.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
31.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/32.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/33.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
34.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/35.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/36.tif.

Harker:1988:UPS

[Har88] Susan Harker. The use of prototyping and simulation in the development of large-scale applications. The Computer Journal, 31(5):420–425, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/5/420.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/420.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/421.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/422.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/423.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/424.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/425.tif.

Hatter:1984:AGN

[Hat84] D. J. Hatter. Algorithm 116: a generalized non-recursive binary tree traversal algorithm. The Computer Journal, 27(2): 178–179, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/179.tif.

Horspool:1982:AOS

[HD82] R. Nigel Horspool and Laurence W. Dunkelman. Analysis and optimization of set expressions. The Computer Journal, 25(3):340– 346, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/340.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/343.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/344.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/345.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/346.tif.

Healey:1981:PIM

[Hea81] M. Healey. The possible impact of minicomputers and microprocessors on mainframe computer manufacturers. The Computer Journal, 24(1):3–8, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/24/1/3.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/3.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/4.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_24/Issue_01/tiff/5.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_24/
Issue_01/tiff/6.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/

8.tif. See correspondence [Wat82].

Heard:1984:C

[Hea84] R. J. Heard. Correspondence. The Computer Journal, 27(1):90, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/1/90.2.full.pdf+html.

Heard:1985:C

[Hea85] R. J. Heard. Correspondence. The Computer Journal, 28(5):543, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/5/543.1.full.pdf+html.

Heard:1986:C

[Hea86] R. J. Heard. Correspondence. The Computer Journal, 29(4):381, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/4/381.1.full.pdf+html.

Hekmatpour:1989:WMU

[Hek89] S. Hekmatpour. A window manager for Unix. The Computer Journal, 32(1):21–23, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/21.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/21.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/23.tif.

Helsgaun:1984:BPS

[Hel84] Keld Helsgaun. Backtrack programming with SIMULA. The Computer Journal, 27(2):151–158, May 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/2/151.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/151.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/152.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/153.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/154.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/155.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/156.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/157.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/158.tif.

Henson:1988:HOT

[Hen88] M. C. Henson. Higher-order transformations and type simulations. The Computer Journal, 31(6):517–524, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/31/6/517.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/517.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/518.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/519.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/520.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/521.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/522.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/523.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/524.tif.

Haskell:1980:SCN

[HH80] R. Haskell and P. G. Harrison. System conventions for non procedural languages. The Computer Journal, 23(2):132–141, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/132.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/132.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/133.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/134.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/135.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/136.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/137.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/138.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/139.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/140.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/141.tif.

Hill:1980:CJS

[Hil80] I. D. Hill. Correspondence: Jumping to some purpose. The Computer Journal, 23(1):94, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_

01/tiff/94.tif. See [ASG79, Rob80a, Tay80, MR82].

Hill:1981:C

[Hil81] I. D. Hill. Correspondence. The Computer Journal, 24(4):377, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/377.3.full.pdf+html.

Hill:1982:BRR

[Hil82] I. D. Hill. Book review: R. E. Berry, Programming Language Translation. The Computer Journal, 25(4):492a–492, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/4/492-a.full.pdf+html.

Hill:1983:PLT

[Hil83] I. D. Hill. Programming language translation. The Computer Journal, 26(4):384, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/4/384.full.pdf+html.

Hatzopoulos:1980:SRI

[HK80] M. Hatzopoulos and J. G. Kollias. Some rules for introducing indexing paths in a primary file. The Computer Journal, 23(3):207– 211, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/207.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/207.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/208.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/209.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/210.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/211.tif.

Hatzopoulos:1982:DOD

[HK82] M. Hatzopoulos and J. G. Kollias. The determination of the optimum database maintenance points. The Computer Journal, 25(1):126–129, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/1/126.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

25/Issue_01/tiff/126.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/127.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/128.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/129.tif.

Hatzopoulos:1985:SRS

[HK85a] Michael Hatzopoulos and J. (Yiannis) Kollias. On the selection of a reduced set of indexes. The Computer Journal, 28(4):406– 408, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/406.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/406.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/407.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/408.tif.

Hatzopoulos:1985:TOS

[HK85b] Michalis Hatzopoulos and J. G. Kollias. Towards the optimal secondary index organisation and secondary index selection. The Computer Journal, 28(5):524–529, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/28/5/524.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/524.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/525.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/526.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/527.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/528.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/529.
tif.

Harrison:1988:ATT

[HK88] P. G. Harrison and H. Khoshnevisan. Algebraic transformation techniques for functional languages. The Computer Journal, 31(3): 229–242, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/229.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/229.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/231.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/232.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/233.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/234.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/236.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/237.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/238.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/239.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/240.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/241.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/242.tif.

Heuring:1989:CDF

[HKPW89] V. P. Heuring, U. Kastens, R. G. Plummer, and W. M. Waite. COMAR: a data format for integration of CFG tools. The Computer Journal, 32(5):445–452, October 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/5/445.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/445.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/446.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/447.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/448.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/449.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/450.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/451.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/452.tif.

Hansen:1980:CAA

[HM80a] J. V. Hansen and Lynn J. McKell. A computer aid for the analysis of complex systems. The Computer Journal, 23(2):98–106, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
23/2/98.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/100.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/

tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/102.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/105.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/106.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/99.
tif.

Horspool:1980:APD

[HM80b] R. N. Horspool and N. Marovac. An approach to the problem of detranslation of computer programs. The Computer Journal, 23(3): 223–229, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/223.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/223.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/224.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/225.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/226.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/227.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/228.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/229.tif.

Hunter:1983:ERA

[HM83] D. G. N. Hunter and A. R. McKenzie. Experiments with relaxation algorithms for breaking simple substitution ciphers. The Computer Journal, 26(1):68–71, February 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/26/1/68.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/68.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/69.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/70.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/71.tif.

Hull:1984:CDD

[HM84] M. Elizabeth C. Hull and R. M. McKeag. Concurrency in the design of data processing systems. The Computer Journal, 27(4):289– 293, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/289.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/289.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/290.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/291.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/293.tif.

Hirschheim:1988:ISU

[HN88] R. Hirschheim and M. Newman. Information systems and user resistance: theory and practice. The Computer Journal, 31(5):398– 408, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/5/398.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/398.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/399.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/400.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/401.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/402.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/403.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/404.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/405.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/406.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/407.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/408.
tif.

Holden:1981:BR

[Hol81] S. C. Holden. Book review. The Computer Journal, 24(3):221, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/221.1.full.pdf+html.

Holdsworth:1984:C

[Hol84] D. Holdsworth. Correspondence. The Computer Journal, 27(3): 285, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/285.3.full.pdf+html.

Hopgood:1980:RDS

[Hop80] F. R. A. Hopgood. Research directions in software technology. The Computer Journal, 23(3):222, August 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/3/222.
full.pdf+html.

Hopper:1981:ROO

[Hop81]  K. Hopper.  Retained objects and operating system in-

terfaces. The Computer Journal, 24(4):331–338, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/24/4/331.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/331.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/332.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/333.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/334.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/335.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/336.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/337.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/338.tif.

Hopgood:1984:BR

[Hop84] F. R. A. Hopgood. Book review. The Computer Journal, 27(3): 287, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/287.3.full.pdf+html.

Horwood:1983:TPM

[Hor83] M. Horwood. Theory and practice of microprocessor. The Computer Journal, 26(4):382b–382, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/4/382-
b.full.pdf+html.

House:1983:PEF

[Hou83] R. T. House. A proposal for an extended form of type checking of expressions. The Computer Journal, 26(4):366–374, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/366.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/367.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/368.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/369.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/370.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/371.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/372.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/373.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/374.tif.

House:1986:ASR

[Hou86] R. T. House. Alternative scope rules for block-structured languages. The Computer Journal, 29(3):253–260, June 1986. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/29/3/253.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/253.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/255.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/256.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/257.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/258.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/259.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/260.tif.

Henstridge:1984:ABO

[HPB84] J. D. Henstridge, R. W. Payne, and R. J. Baker. Algorithm 117: Buffered output in FORTRAN. The Computer Journal, 27(2): 179–184, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_

journal/hdb/Volume_27/Issue_02/tiff/179.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/180.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/181.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/182.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/183.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/184.tif. See note

[WH87].

Helary:1988:DAM

[HPR88] J.-M. Hélary, N. Plouzeau, and M. Raynal. A distributed algorithm for mutual exclusion in an arbitrary network. The Computer Journal, 31(4):289–295, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/4/289.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/289.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/290.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/291.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/292.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/293.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/294.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/295.tif.

Hudson:1982:DST

[HS82] R. W. A. Hudson and J. A. Self. A dialogue system to teach database concepts. The Computer Journal, 25(1):135–139, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/135.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/135.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/136.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/137.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/139.tif.

Hehner:1983:PGE

[HS83] Eric C. R. Hehner and Brad A. Silverberg. Programming with grammars: an exercise in methodology-directed language design. The Computer Journal, 26(3):277–281, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/277.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/277.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/278.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/279.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/280.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/281.tif.

Hartley:1981:CWA

[HT81] D. F. Hartley and The Editor. Correspondence: The Wilkes Award. The Computer Journal, 24(2):192, May 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/192.tif.

Halbert:1989:GAZ

[HTW89] A. R. Halbert, S. J. P. Todd, and J. R. Woodwark. Generalizing active zones for set-theoretic solid models. The Computer Journal, 32(1):86–89, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/86.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/87.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/88.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/89.tif.

Hubbard:1988:C

[Hub88] M. H. Hubbard. Correspondence. The Computer Journal, 31 (1):92, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/1/92.1.full.pdf+html.

Hubbard:1989:C

[Hub89] M. H. Hubbard. Correspondence. The Computer Journal, 32(4): 379, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/379.1.full.pdf+html.

Hughes:1989:WFP

[Hug89] J. Hughes. Why functional programming matters. The Computer Journal, 32(2):98–107, April 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/2/98.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_02/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_32/Issue_02/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_32/
Issue_02/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/103.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/105.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/106.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/107.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/98.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/
99.tif.

Hull:1986:ICN

[Hul86] M. Elizabeth C. Hull. Implementations of the CSP notation for concurrent systems. The Computer Journal, 29(6):500–505, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/6/500.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/500.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/501.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/502.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/503.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/504.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/505.
tif.

Hunt:1981:C

[Hun81a] J. G. Hunt. Correspondence. The Computer Journal, 24(4):377, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/377.1.full.pdf+html.

Hunt:1981:CPT

[Hun81b] J. G. Hunt. Correspondence: Premature termination. The Computer Journal, 24(1):95, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/95.tif. See [Zar80].

Hunter:1985:ARK

[Hun85] D. G. N. Hunter. Algorithm 121: RSA key calculation in Ada. The Computer Journal, 28(3):343–348, July 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/343.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/343.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/344.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/345.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/346.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/347.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/348.tif. See note [Wic87].

Hutchinson:1984:SNH

[Hut84] Alan Hutchinson. Short notes: the halting problem does not matter. The Computer Journal, 27(4):376, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/27/4/376.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/376.
tif.

Hutchinson:1986:DSA

[Hut86a] A. Hutchinson. A data structure and algorithm for a selfaugmenting heuristic program. The Computer Journal, 29(2): 135–150, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/135.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/135.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/136.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/137.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_29/Issue_02/tiff/139.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/140.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/141.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/142.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/143.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/144.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/146.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/147.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/148.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/149.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/150.
tif.

Hutchinson:1986:SPP

[Hut86b] A. Hutchinson. Some practical principles for design of maintainable systems. The Computer Journal, 29(1):20–23, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
29/1/20.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
21.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/23.tif.

Hutchinson:1989:IKS

[Hut89] A. Hutchinson. Inheritance and kinds of slots. The Computer Journal, 32(1):63–67, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/63.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/63.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/65.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/67.
tif.

Henschen:1988:TIP

[HW88] L. J. Henschen and B. P. Weems. A top-down inference procedure for template dependencies. The Computer Journal, 31(1):51–55, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/1/51.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_31/Issue_01/tiff/51.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_
01/tiff/52.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/54.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/55.tif.

Hopkins:1989:WCO

[HW89] T. P. Hopkins and M. I. Wolczko. Writing concurrent object-oriented programs using Smalltalk-80. The Computer Journal, 32(4):341–350, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/4/341.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/343.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/344.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/345.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/346.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/347.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/348.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/349.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/350.
tif.

Hill:1987:AST

[HWW87] I. D. Hill, B. A. Wichmann, and D. R. Woodall. Algorithm 123: Single transferable vote by Meek’s method. The Computer Journal, 30(3):277–281, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/3/277.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

30/Issue_03/tiff/277.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/278.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/279.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/280.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/281.
tif.

Harrison:1981:ENR

[HY81] M. A. Harrison and A. Yehudai. Eliminating null rules in linear time. The Computer Journal, 24(2):156–161, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/24/2/156.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/156.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/158.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/160.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/161.
tif.

Ibbett:1981:C

[Ibb81] R. N. Ibbett. Correspondence. The Computer Journal, 24(4): 377, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/377.2.full.pdf+html.

Ibbett:1985:CHP

[IEH^{+}85] R. N. Ibbett, D. A. Edwards, T. P. Hopkins, C. K. Cadogan, and D. A. Train. Centrenet — a high performance local area network. The Computer Journal, 28(3):231–242, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/231.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/231.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/232.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/233.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/234.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/235.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/236.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/237.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/238.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/239.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/240.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/241.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/242.tif.

Igarashi:1980:SAP

[Iga80] Y. Igarashi. The size of arrays for a prime implicant generating algorithm. The Computer Journal, 23(1):73–77, February 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/1/73.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/73.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/74.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/75.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/76.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
77.tif.

Ince:1986:EER

[IH86a] D. C. Ince and S. Hekmatpour. Empirical evaluation of random testing. The Computer Journal, 29(4):380, August 1986. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/380.tif.

Ince:1986:SNE

[IH86b] D. C. Ince and S. Hekmatpour. Short notes: an empirical evaluation of random testing. The Computer Journal, 29(4):380, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/4/380.full.pdf+html.

Ince:1987:AGT

[Inc87] D. C. Ince. The automatic generation of test data. The Computer Journal, 30(1):63–69, February 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/30/1/63.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/63.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_30/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_
01/tiff/65.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/67.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/68.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/69.tif.

Inglis:1980:LPS

[Ing80] Jim Inglis. Learning to program in structured cobol; structured analysis and system specification. The Computer Journal, 23(3):285a–285, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/285-a.full.pdf+html.

Inglis:1981:BR

[Ing81] J. Inglis. Book review. The Computer Journal, 24(2):172, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/2/172.3.full.pdf+html.

Inglis:1982:JSP

[Ing82] Jim Inglis. Jumping to some purpose. The Computer Journal, 25(4):495a–495, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/4/495-a.full.pdf+html.

Inglis:1983:AC

[Ing83] J. Inglis. Advanced cobol. The Computer Journal, 26(4):381b– 381, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/381-b.full.pdf+html.

Iyer:1986:PSC

[IP86] K. Viswanathan Iyer and L. M. Patnaik. Performance study of a centralized concurrency control algorithm for distributed database systems using Simula. The Computer Journal, 29(2): 118–126, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/118.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/118.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/119.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/120.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/121.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/122.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/123.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/124.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/125.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/126.tif.

Irving:1984:SNP

[Irv84] Robert W. Irving. Short notes: Permutation backtracking in lexicographic order. The Computer Journal, 27(4):373–375, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/373.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/373.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/374.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/375.tif.

Izatt:1981:DSD

[IS81] W. T. Izatt and E. A. Schmitz. Data structures and descriptors in the ICL 2900 series and beyond. The Computer Journal, 24(4):301–307, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/4/301.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/301.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/302.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/303.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/304.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/305.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/306.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/307.tif.

Ince:1986:RGC

[IW86] D. Ince and M. Woodman. The rapid generation of a class of software tools. The Computer Journal, 29(2):151–160, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/151.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/151.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/152.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/153.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/154.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/155.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/156.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/158.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/160.tif.

Incerti:1981:AFS

[IZP81] S. Incerti, F. Zirilli, and V. Parisi. Algorithm 111: a FOR-TRAN subroutine for solving systems of nonlinear simultaneous equations. The Computer Journal, 24(1):87–91, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/88.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/89.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/90.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
91.tif.

Jackson:1980:MCG

[Jac80a] M. A. Jackson. Methodology in computer graphics. The Computer Journal, 23(3):200b–200, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/200-b.full.pdf+html.

Jackson:1980:SP

[Jac80b] M. A. Jackson. Structured programming. The Computer Journal, 23(3):255, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/255.full.pdf+html.

Jackson:1980:SSA

[Jac80c] M. A. Jackson. Structured systems analysis: Tools and techniques. The Computer Journal, 23(3):255a–255, August 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/3/255-
a.full.pdf+html.

Jacobs:1980:BR

[Jac80d] D. Jacobs. Book reviews. The Computer Journal, 23(2):146, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/146.2.full.pdf+html.

Jajodia:1986:RMD

[Jaj86] S. Jajodia. Recognizing multivalued dependencies in relation schemas. The Computer Journal, 29(5):458–459, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/458.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/458.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/459.tif.

Jakobsson:1985:SRL

[Jak85] Matti Jakobsson. Sampling without replacement in linear time. The Computer Journal, 28(4):412–413, August 1985. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/28/4/412.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/412.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/413.
tif.

James:1980:UI

[Jam80] E. B. James. The user interface. The Computer Journal, 23(1):25– 28, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/1/25.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_01/tiff/25.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_01/tiff/26.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/28.tif.

James:1981:MIC

[Jam81] E. B. James. Microelectronics: the impact on computer centres. The Computer Journal, 24(1):9–13, February 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/24/1/9.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/10.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/11.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/12.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/13.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
9.tif.

Jamieson:1987:SNN

[Jam87] M. J. Jamieson. Short note: a note on the convergence of an iterative scheme for solving a quadratic equation. The Computer Journal, 30(2):189–190, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/2/189.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/189.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/190.tif.

Jamieson:1988:C

[Jam88] M. J. Jamieson. Correspondence. The Computer Journal, 31(3): 286, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/286.1.full.pdf+html.

Jamieson:1989:SNR

[Jam89] M. J. Jamieson. Short notes: Rapidly converging iterative formulae for finding square roots and their computational efficiencies. The Computer Journal, 32(1):93–94, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/93.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/93.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/94.tif.

Jankowitz:1988:DPS

[Jan88] H. T. Jankowitz. Detecting plagiarism in student Pascal programs. The Computer Journal, 31(1):1–8, February 1988. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/31/1/1.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/1.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_31/Issue_01/tiff/2.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_31/
Issue_01/tiff/3.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
5.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/6.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
8.tif.

Jones:1984:ARI

[JBJ84] B. Jones, M. Banerjee, and L. Jones. Algorithm 118: Root isolation for transcendental equations. The Computer Journal, 27(2): 184–187, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/184.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/185.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/186.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/187.
tif.

Jagodzinski:1986:RMM

[JC86] A. P. Jagodzinski and D. D. Clarke. A review of methods for measuring and describing users’ attitudes as an essential constituent of systems analysis and design. The Computer Journal, 29(2):97–102, April 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/29/2/97.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_02/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_29/Issue_02/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_29/
Issue_02/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/97.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/
98.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/99.tif.

Jagodzinski:1988:MAM

[JC88] A. P. Jagodzinski and D. D. Clarke. A multidimensional approach to the measurement of human-computer performance. The Computer Journal, 31(5):409–419, October 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/5/409.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/409.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/410.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/411.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/412.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/413.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/414.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/415.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/416.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/417.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/418.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/419.tif.

Jenkins:1984:BR

[Jen84] J. O. Jenkins. Book review. The Computer Journal, 27(3):286– 287, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/286.4.full.pdf+html.

Jarosz:1986:CTP

[JJ86] J. Jarosz and J. R. Jaworowski. Computer tree — the power of parallel computations. The Computer Journal, 29(2):103– 108, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/103.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/105.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/106.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/107.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/108.
tif.

Jones:1980:ASD

[JK80] W. T. Jones and S. A. Kirk. APL as a software design specification language. The Computer Journal, 23(3):230–232, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/230.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/230.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/231.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/232.tif.

Johnson:1980:CUD

[JL80] S. N. Johnson and Jay Louise. Correspondence: Using data abstraction for logical design. The Computer Journal, 23(4):380, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/380.tif. See [Wel80a,

Ree80c].

Jones:1981:BR

[Jon81a] J. R. Jones. Book review. The Computer Journal, 24(4):307, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/307.full.pdf+html.

Jones:1981:BRb

[Jon81b] S. Jones. Book review. The Computer Journal, 24(2):190, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/2/190.1.full.pdf+html.

Jones:1982:EDB

[Jon82] S. Jones. Effective data base design. The Computer Journal, 25(1): 83, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/83.full.pdf+html.

Joseph:1986:FRT

[JP86] M. Joseph and P. Pandya. Finding response times in a real-time system. The Computer Journal, 29(5):390–395, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/390.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/390.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/391.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/392.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/393.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/394.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/395.
tif.

Jones:1981:BRaB

[JR81] S. Jones and C. M. Reeves. Book reviews: Principles of Database Systems, by J. D. Ullman, 1980; Mathematical Methods in Computer Graphics and Design, edited by K. W. Brodie, 1980. The Computer Journal, 24(2):190, May 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/190.tif.

Jones:1989:FPO

[JS89] S. B. Jones and A. F. Sinclair. Functional programming and operating systems. The Computer Journal, 32(2):162– 174, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/2/162.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/162.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/163.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/164.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/165.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/166.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/167.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/169.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/

tiff/170.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/171.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/172.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/173.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/174.tif.

Johnson:1982:UII

[JW82] J. S. Johnson and D. B. Webster. Updating an inverted index — a performance comparison of two techniques. The Computer Journal, 25(2):169–175, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/169.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/169.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/170.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/171.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/172.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/173.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/174.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/175.tif.

Kulkarni:1986:EEF

[KA86] K. G. Kulkarni and M. P. Atkinson. EFDM: Extended functional data model. The Computer Journal, 29(1):38–46, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
29/1/38.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/38.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
39.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/40.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/41.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
42.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/43.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/44.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
45.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/46.tif.

Kacsuk:1987:DPS

[KB87] P. Kacsuk and A. Bale. DAP Prolog: a set-oriented approach to Prolog. The Computer Journal, 30(5):393–403, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/393.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/393.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/394.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/395.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/396.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/397.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/398.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/399.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/400.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/401.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/402.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/403.
tif.

Kontos:1988:KAT

[KC88] John Kontos and J. C. Cavouras. Knowledge acquisition from technical texts using attribute grammars. The Computer Journal, 31(6):525–530, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/6/525.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/525.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/526.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/527.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/528.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/529.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/530.tif.

Kecskes:1987:SMB

[Kec87] C. S. Kecskes. A simple method for building up large computer systems from small computers. The Computer Journal, 30(4):295– 297, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/295.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/295.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/296.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/297.tif.

Kent:1980:MBH

[Ken80] S. A. Kent. A minicomputer based high level network control language. The Computer Journal, 23(2):124–131, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/23/2/124.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/124.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/125.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/126.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/127.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/128.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/129.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/131.tif.

Kent:1981:BR

[Ken81] P. Kent. Book review. The Computer Journal, 24(3):221, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/3/221.2.full.pdf+html.

Kent:1985:ENW

[Ken85] Philippe Kent. An efficient new way to represent multidimensional data. The Computer Journal, 28(2):184–190, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/2/184.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/184.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/185.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/186.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/187.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_28/Issue_02/tiff/188.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/189.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/190.tif.

Kimura:1985:DDM

[KGC85] T. D. Kimura, W. D. Gillett, and J. R. Cox, Jr. A design of a data model based on abstraction of symbols. The Computer Journal, 28(3):298–308, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/298.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/298.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/299.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/301.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/302.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/303.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/304.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/305.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/306.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/307.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/308.tif.

Kollias:1981:CAS

[KH81] J. G. Kollias and M. Hatzopoulos. Criteria to aid in solving the problem of allocating copies of a file in a computer network. The Computer Journal, 24(1):29–30, February 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/24/1/29.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/30.tif.

Klein:1987:CFD

[KH87] H. K. Klein and R. A. Hirschheim. A comparative framework of data modelling paradigms and approaches. The Computer Journal, 30(1):8–15, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/30/1/8.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_
01/tiff/10.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/12.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/13.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/14.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
15.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/8.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/9.tif.

Khabaza:1980:MI

[Kha80] I. M. Khabaza. Machine intelligence. The Computer Journal, 23 (3):229, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/229.full.pdf+html.

Kiessling:1988:APS

[Kie88] W. Kiessling. Access path selection in databases with intelligent disc subsystems. The Computer Journal, 31(1):41–50, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
31/1/41.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/41.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
42.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/43.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/44.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
45.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/46.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/47.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
48.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/49.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/50.tif.

Kilgour:1981:BR

[Kil81a] A. C. Kilgour. Book review. The Computer Journal, 24(3):257, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/257.full.pdf+html.

Kilgour:1981:BRB

[Kil81b] A. C. Kilgour. Book review: ILP: Intermediate Language for Pictures, by P. J. W. ten Hagen, T. Hagen, P. Klint, H. Noot, N. J. Sint, and A. H. Veen, 1981. The Computer Journal, 24(3):257, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/257.tif.

Kilgour:1984:C

[Kil84] A. C. Kilgour. Correspondence. The Computer Journal, 27(4): 377, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/377.2.full.pdf+html.

King:1981:BR

[Kin81] P. King. Book review. The Computer Journal, 24(2):117, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/2/117.2.full.pdf+html.

Kiossev:1981:CPP

[Kio81] Nickolia Kiossev. Correspondence: Points and polygons. The Computer Journal, 24(1):94–95, February 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/94.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/95.

tif. See [dS80].

Kafura:1989:IAB

[KL89] D. G. Kafura and K. H. Lee. Inheritance in actor based concurrent object-oriented languages. The Computer Journal, 32(4): 297–304, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/297.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/297.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/298.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/299.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/

tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/301.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/302.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/304.tif.

Knuth:1984:LP

[Knu84] Donald E. Knuth. Literate programming. The Computer Journal, 27(2):97–111, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/27/2/97.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
02/tiff/100.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/101.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/102.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/103.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/104.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/105.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/106.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/107.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/108.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/109.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/110.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/111.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/97.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/98.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/99.tif.

Kollias:1983:EMR

[Kol83] J. G. Kollias. The effect of missing, retroactive, erroneous and/ or late data in designing accounting applications. The Computer Journal, 26(2):117–123, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/2/117.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/117.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/118.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/120.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/121.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/122.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/123.tif.

Katajainen:1989:AAS

[KR89] J. Katajainen and T. Raita. An approximation algorithm for space-optimal encoding of a text. The Computer Journal, 32(3): 228–237, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/3/228.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/228.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/229.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/230.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/231.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/232.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/233.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/234.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/235.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/236.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/237.tif.

Keedy:1982:SRW

[KRR82] J. L. Keedy, J. Rosenberg, and K. Ramamohanarao. On synchronizing readers and writers with semaphores. The Computer Journal, 25(1):121–125, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/1/121.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/121.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/122.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/123.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/124.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/125.
tif.

Kieburtz:1982:SAC

[KS82] Richard Kieburtz and Abraham Silberschatz. On the static access-control mechanism in Concurrent Pascal. The Computer Journal, 25(1):17–21, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/17.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/17.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/18.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/19.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/20.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/21.
tif.

Kant:1985:EPR

[KS85] K. Kant and A. Silberschatz. Error propagation and recovery in concurrent environments. The Computer Journal, 28(5):466– 473, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/466.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/466.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/467.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/468.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/469.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/470.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/471.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/472.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/473.tif.

Kowalski:1987:C

[KS87] R. Kowalski and M. Sergot. Correspondence. The Computer Journal, 30(3):285, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/30/3/285.1.full.pdf+html.

Kalvin:1984:TPR

[KV84] Alan D. Kalvin and Yaakov L. Varol. On two problems reducible to topological sorting. The Computer Journal, 27(2): 176–177, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/2/176.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/176.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/177.tif.

Kirby:1989:CGR

[KVW89] G. H. Kirby, M. Visvalingam, and P. Wade. Computer graphics: Recognition and representation of a hierarchy of polygons with holes. The Computer Journal, 32(6):554–562, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/554.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/554.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/555.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/556.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/557.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/558.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/559.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/560.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/561.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/562.tif.

Kouvatsos:1984:OBS

[KW84] D. D. Kouvatsos and S. K. Wong. On optimal blocking of sequential files. The Computer Journal, 27(4):321–327, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/321.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/321.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/322.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/323.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/324.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/325.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/326.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/327.tif.

Land:1980:CI

[Lan80] F. F. Land. Chairman’s introduction. The Computer Journal, 23 (1):3–5, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/1/3.2.full.pdf+html.

Land:1982:T

[Lan82] F. F. Land. Tutorial. The Computer Journal, 25(2):283–285, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_02/tiff/283.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_25/Issue_02/tiff/284.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_25/
Issue_02/tiff/285.tif.

Lance:1984:BR

[Lan84] G. N. Lance. Book review. The Computer Journal, 27(3):287, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/3/287.1.full.pdf+html.

Land:1985:ITE

[Lan85] Frank Land. Is an information theory enough? The Computer Journal, 28(3):211–215, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/211.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/211.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/212.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/213.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/214.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/215.
tif.

Larson:1982:EWC

[Lar82] Per-Åke Larson. Expected worst-case performance of hash files. The Computer Journal, 25(3):347–352, August 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/3/347.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/347.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/348.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/350.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/352.tif.

Larson:1985:PAS

[Lar85] Per-Åke Larson. Performance analysis of a single-file version of linear hashing. The Computer Journal, 28(3):319– 329, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/319.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/319.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/320.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/321.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/322.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/323.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/324.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/325.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/326.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/327.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/328.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/329.
tif.

Lavington:1981:BR

[Lav81] S. H. Lavington. Book review. The Computer Journal, 24(2): 142, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/142.2.full.pdf+html.

Laver:1983:LCP

[Lav83] Murray Laver. Language as a cognitive process. The Computer Journal, 26(4):381, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/4/381.full.pdf+html.

Laver:1986:C

[Lav86] M. Laver. Correspondence. The Computer Journal, 29(2):190, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/2/190.1.full.pdf+html.

Layzell:1985:HMP

[Lay85] P. J. Layzell. The history of macro processors in programming language extensibility. The Computer Journal, 28(1):29–33, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
28/1/29.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/29.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
30.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/31.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/32.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
33.tif.

Lazos:1981:ICS

[Laz81] C. Lazos. Increasing computer system productivity software and hardware methods: a comparative study. The Computer Journal, 24(3):230–234, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/3/230.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/231.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/232.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/233.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/234.
tif.

Luker:1986:PGG

[LB86] P. A. Luker and A. Burns. Program generators and generation software. The Computer Journal, 29(4):315–321, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/315.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/315.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

29/Issue_04/tiff/316.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/317.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/318.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/319.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/320.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/321.tif.

Lucido:1981:SMV

[LCP81] Tony P. Lucido, Rahul Chattergy, and Udo W. Pooch. Survey of microprogram verification and validation methods. The Computer Journal, 24(2):139–142, May 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/139.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/139.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/140.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/141.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/142.tif.

LeQuesne:1988:IOF

[Le 88] P. N. Le Quesne. Individual and organisational factors and the design of IPSEs. The Computer Journal, 31(5):391–397, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/5/391.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/391.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/392.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/393.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/394.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/395.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/396.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/397.tif.

Levary:1986:AIA

[LE86] R. R. Levary and W. D. Edwards. Analysing the impact of adding a new software system on main memory usage. The Computer Journal, 29(6):522–526, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/29/6/522.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/522.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/523.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/524.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/525.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/526.tif.

Lea:1980:DNC

[Lea80] R. M. Lea. Digital networks and computer systems 2nd edition. The Computer Journal, 23(3):285b–285, August 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/3/285-
b.full.pdf+html.

Lea:1986:VWA

[Lea86] R. M. Lea. VLSI and WSI associative string processors for costeffective parallel processing. The Computer Journal, 29(6):486– 494, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/6/486.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/486.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/487.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/488.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/489.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/490.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/491.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/492.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/493.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/494.tif.

Lee:1982:DNS

[Lee82] Graham Lee. DOPLs: a new style of programming. The Computer Journal, 25(2):176–182, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/176.full.pdf+

html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/176.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/177.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/178.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/179.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/180.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/181.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/182.tif.

Lehman:1981:BR

[Leh81] M. Lehman. Book review. The Computer Journal, 24(4):346, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/346.1.full.pdf+html.

Leith:1983:HSP

[Lei83] Philip Leith. Hierarchically structured production rules. The Computer Journal, 26(1):1–5, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/1/1.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/1.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/2.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_26/Issue_01/tiff/3.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_26/
Issue_01/tiff/4.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/5.tif.

Leith:1986:FEL

[Lei86] P. Leith. Fundamental errors in legal logic programming. The Computer Journal, 29(6):545–552, December 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/6/545.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/545.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/546.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/547.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/548.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/549.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

29/Issue_06/tiff/550.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/551.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/552.tif.

Leith:1987:PSF

[Lei87] P. Leith. A programmed, skeleton formal specification method: the OUFDM. The Computer Journal, 30(4):337–342, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/337.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/337.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/338.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/339.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/340.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/341.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/342.
tif.

Leith:1988:C

[Lei88] P. Leith. Correspondence. The Computer Journal, 31(1):92–93, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/1/92.2.full.pdf+html.

Lemos:1980:DCM

[Lem80] R. S. Lemos. Discussion and correspondence: Measuring students’ attitudes towards structured walk-throughs. The Computer Journal, 23(4):377–380, November 1980. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
04/tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/378.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/379.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/380.tif.

Lescanne:1983:BCA

[Les83a] Pierre Lescanne. Behavioural categoricity of abstract data type specifications. The Computer Journal, 26(4):289–292, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/289.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/289.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/290.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/291.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/292.tif.

Lesuisse:1983:SLD

[Les83b] R. Lesuisse. Some lessons drawn from the history of the binary search algorithm. The Computer Journal, 26(2):154– 163, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/154.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/154.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/155.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/156.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/157.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/158.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/159.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/160.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/161.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/162.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/163.tif.

Leung:1982:IOF

[Leu82a] C. H. C. Leung. An improved optimal-fit procedure for dynamic storage allocation. The Computer Journal, 25(2):199– 206, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/2/199.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/199.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/201.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/203.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/204.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/205.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/206.tif.

Leung:1982:SMP

[Leu82b] C. H. C. Leung. A simple model for the performance analysis of disc storage fragmentation. The Computer Journal, 25(2):193– 198, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/2/193.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/193.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/195.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/197.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/198.
tif.

Leung:1983:ADF

[Leu83a] C. H. C. Leung. Analysis of disc fragmentation using Markov chains. The Computer Journal, 26(2):113–116, May 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/2/113.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/113.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/114.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/115.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/116.tif.

Leung:1983:MDL

[Leu83b] C. H. C. Leung. A model for disc locality referencing. The Computer Journal, 26(3):196–198, August 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/3/196.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/196.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/197.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/198.tif.

Lew:1984:PCD

[Lew84] Art Lew. Proof of correctness of decision table programs. The Computer Journal, 27(3):230–232, August 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/3/230.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/231.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/232.tif.

Libera:1986:UBT

[LG86] F. D. Libera and F. Gosen. Using B-trees to solve geographic range queries. The Computer Journal, 29(2):176– 181, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/176.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/176.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/177.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/180.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/181.
tif.

Licker:1989:CMH

[Lic89] P. S. Licker. Computer management: Human resource management and business success in small, hi-tech firms: Some lessons for data processing. The Computer Journal, 32(6):533– 540, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/533.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/533.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/534.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/535.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/536.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/537.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/538.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

32/Issue_06/tiff/539.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/540.tif.

Lindley:1981:BR

[Lin81] J. Lindley. Book review. The Computer Journal, 24(1):47, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/1/47.2.full.pdf+html.

Lipsky:1980:STS

[Lip80] L. Lipsky. A study of time sharing systems considered as queueing networks of exponential servers. The Computer Journal, 23(4):290–297, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
04/tiff/290.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/291.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/292.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/293.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/294.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/295.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/296.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/297.
tif.

List:1984:C

[Lis84] W. List. Correspondence. The Computer Journal, 27(3):285, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/3/285.2.full.pdf+html.

Llewellyn:1983:FSM

[Lle83] John A. Llewellyn. Finite state models in the study of comma-free codes. The Computer Journal, 26(4):303–311, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/303.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/303.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/304.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/305.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/306.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/307.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/308.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/309.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/310.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/311.tif.

Llewellyn:1985:CVL

[Lle85] John A. Llewellyn. The construction of variable-length commafree codes. The Computer Journal, 28(4):379–390, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/28/4/379.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/379.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/380.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/381.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/382.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/383.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/384.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/385.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/386.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/387.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/388.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/389.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/390.tif.

Llewellyn:1987:DCS

[Lle87] J. A. Llewellyn. Data compression for a source with Markov characteristics. The Computer Journal, 30(2):149–156, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/2/149.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/149.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/150.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/151.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/152.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/153.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/154.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/155.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/156.tif.

Lucas:1988:CAM

[LM88] C. Lucas and G. Mitra. Computer-Assisted Mathematical Programming (Modelling) Systems: CAMPS. The Computer Journal, 31(4):364–375, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/4/364.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/365.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/366.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/367.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/368.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/369.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/370.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/371.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/372.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/373.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/374.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/375.tif.

Lo:1984:SNU

[Lo84] C. J. Lo. Short notes: Unifying multiple semantic intentions for a syntactic construct. The Computer Journal, 27(1):83–84, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
27/1/83.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/83.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
84.tif.

Lombardi:1988:ACB

[Lom88] F. Lombardi. Analysis of comparison-based diagnosable systems using temporal criteria. The Computer Journal, 31(3):201– 208, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/201.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/201.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/202.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/203.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/205.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/206.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/207.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/208.tif.

Lovett:1981:BR

[Lov81] R. Lovett. Book review. The Computer Journal, 24(1):41, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/1/41.1.full.pdf+html.

Laenens:1989:BCO

[LSV89] E. Laenens, F. Staes, and D. Vermeir. Browsing à la carte in object-oriented databases. The Computer Journal, 32(4):333– 340, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/333.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/333.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/334.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/335.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/336.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/337.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/338.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/339.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/340.tif.

Lu:1989:SNI

[Lu89] Xue-Miao M. Lu. Short note: an iterative solution for the 4-peg Towers of Hanoi. The Computer Journal, 32(2):187– 189, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/2/187.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/187.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/188.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/189.tif.

Lunnon:1986:C

[Lun86] W. F. Lunnon. Correspondence. The Computer Journal, 29(5): 478, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/478.1.full.pdf+html.

Lutz:1980:ART

[Lut80] R. K. Lutz. An algorithm for the real time analysis of digitised images. The Computer Journal, 23(3):262–269, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/262.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/262.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/263.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/264.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/265.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/266.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/267.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/268.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/269.tif.

Leung:1983:DDE

[LW83] C. H. C. Leung and K. Wolfenden. Disc database efficiency: a scheme for detailed assessment based on semi-Markov models. The Computer Journal, 26(1):10–14, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/26/1/10.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/10.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/11.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_26/Issue_01/tiff/13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/14.
tif.

Leung:1985:AOD

[LW85a] C. H. C. Leung and K. Wolfenden. Analysis and optimisation of data currency and consistency in replicated distributed databases. The Computer Journal, 28(5):518–523, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/518.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/518.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/519.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/520.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/521.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/522.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/523.
tif.

Leung:1985:EDA

[LW85b] C. H. C. Leung and K. Wolfenden. Estimating disc access patterns using diffusion models. The Computer Journal, 28(4):409– 411, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/409.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/409.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/410.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/411.tif.

Leung:1985:MMF

[LW85c] C. H. C. Leung and K. Wolfenden. Mathematical models of file growth. The Computer Journal, 28(2):179–183, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/2/179.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/179.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/180.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/181.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/

tiff/182.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/183.tif.

Lyon:1985:AHT

[Lyo85] Gordon Lyon. Achieving hash table searches in one or two bucket probes. The Computer Journal, 28(3):313–318, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/313.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/313.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/314.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/315.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/316.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/317.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/318.
tif.

McDonald:1989:DSC

[MA89] C. S. McDonald and L. Allison. Denotational semantics of a command interpreter and their implementation in Standard ML. The Computer Journal, 32(5):422–431, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/32/5/422.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/422.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/423.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/424.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/425.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/426.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/427.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/428.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/429.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/430.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/431.tif.

Macleod:1981:RMB

[Mac81] Ian A. Macleod. Relational model as a basis for document retrieval system design. The Computer Journal, 24(4):312–315, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/312.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/312.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/314.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/315.tif.

MacCallum:1983:C

[Mac83] I. R. MacCallum. Correspondence. The Computer Journal, 26 (2):189, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/189.3.full.pdf+html.

Maclaren:1985:C

[Mac85] N. M. Maclaren. Correspondence. The Computer Journal, 28 (4):446, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/446.1.full.pdf+html.

Maddison:1980:FLH

[Mad80] J. A. T. Maddison. Fast lookup in hash tables with direct rehashing. The Computer Journal, 23(2):188–189, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/23/2/188.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/188.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

23/Issue_02/tiff/189.tif. See correspondence [Bre81b, CF81].

Mahjoub:1984:SED

[Mah84] A. Mahjoub. On the static evaluation of distributed systems performance. The Computer Journal, 27(3):201–208, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/201.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/201.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/202.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/203.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/

tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/205.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/206.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/207.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/208.tif.

Mahmoud:1986:EDD

[Mah86] H. M. Mahmoud. The expected distribution of degrees in random binary search trees. The Computer Journal, 29(1):36–37, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
29/1/36.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/36.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
37.tif.

Maibaum:1985:DIA

[Mai85] T. S. E. Maibaum. Database instances, abstract data types and database specification. The Computer Journal, 28(2):154– 161, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/2/154.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/154.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/155.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/156.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/157.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/158.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/159.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/160.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/161.tif.

Makinen:1989:ITA

[Mäk89] E. Mäkinen. On implementing two adaptive data-compression schemes. The Computer Journal, 32(3):238–240, June 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/32/3/238.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/238.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/239.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/240.tif.

Marshall:1980:CS

[Mar80] G. Marshall. Computer security. The Computer Journal, 23(3): 261, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/261.full.pdf+html.

Marovac:1982:RBB

[Mar82] Nenad Marovac. The rotating bus as a basis for interprocess communication in distributed systems. The Computer Journal, 25(1):22–31, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/22.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/24.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/28.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
29.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/30.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/31.tif.

Mackay:1986:UBS

[MBC^{+}86] D. Mackay, G. Ball, M. Crowe, M. Hughes, D. Jenkins, and C. Nicol. A UNIX-based system for software configuration management. The Computer Journal, 29(6):527–530, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/6/527.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/527.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/528.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/529.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/530.tif.

McCulloch:1982:QDS

[McC82] C. M. McCulloch. Quickshunt — a distributive sorting algorithm. The Computer Journal, 25(1):102–104, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/25/1/102.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/102.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/103.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/104.tif.

McIlroy:1982:NSD

[McI82a] M. D. McIlroy. The number of states of a dynamic storage allocation system. The Computer Journal, 25(3):388–392, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/388.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/388.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/389.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/390.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/391.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/392.tif.

McIlroy:1982:PGW

[McI82b] M. D. McIlroy. A program for generating word squares. The Computer Journal, 25(1):44a–44, February 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/1/44-a.full.pdf+html.

McLean:1985:MGS

[McL85] David McLean. A method of generating surfaces as a composite of cyclide patches. The Computer Journal, 28(4):433– 438, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/433.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/433.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/434.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/435.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/

tiff/436.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/437.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/438.
tif.

Melhem:1989:SMW

[MCL89] R. G. Melhem, D. M. Chiarulli, and S. P. Levitan. Space multiplexing of waveguides in optically interconnected multiprocessor systems. The Computer Journal, 32(4):362–369, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/362.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/362.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/363.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/364.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/365.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/366.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/367.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/368.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/369.tif.

Macmillan:1984:SNF

[MD84] R. J. Macmillan and J. H. Davenport. Short notes: Factoring medium-sized integers. The Computer Journal, 27(1):84–85, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/1/84.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/84.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/85.tif.

Mayhew:1987:APC

[MD87] P. J. Mayhew and P. A. Dearnley. An alternative prototyping classification. The Computer Journal, 30(6):481–484, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/481.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/481.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/482.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/483.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/484.tif.

Messham:1980:CPC

[ME80] D. K. Messham and A. Elliott. Correspondence: Procedure calling and structured architecture. The Computer Journal, 23(4):382– 383, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/382.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/383.tif. See [BB80b].

Megson:1988:SNI

[ME88a] G. M. Megson and D. J. Evans. Short notes: Improved matrix product computation using double-pipeline systolic arrays. The Computer Journal, 31(6):567–569, December 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/6/567.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/567.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/568.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/569.tif.

Megson:1988:USD

[ME88b] G. M. Megson and D. J. Evans. The unification of systolic differencing algorithms. The Computer Journal, 31(1):83–91, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
31/1/83.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/83.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
84.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/85.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/86.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
87.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
90.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/91.tif.

Mearns:1983:C

[Mea83] I. Mearns. Correspondence. The Computer Journal, 26(2):190, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/2/190.1.full.pdf+html.

Mepham:1984:C

[Mep84] M. J. Mepham. Correspondence. The Computer Journal, 27(4): 377–378, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/377.3.full.pdf+html.

Mor:1982:PGV

[MF82] M. Mor and A. S. Fraenkel. Permutation generation on vector processors. The Computer Journal, 25(4):423–428, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/423.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/423.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/424.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/425.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/426.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/427.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/428.
tif.

Miller:1988:MDD

[MF88] Mirka Miller and I. Fris. Minimum diameter of diregular digraphs of degree 2. The Computer Journal, 31(1):71–75, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
31/1/71.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/71.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
72.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/73.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/74.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
75.tif.

Maddison:1985:DDL

[MG85] R. N. Maddison and A. J. Gawronski. A data dictionary for learning data analysis. The Computer Journal, 28(3):270– 281, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/270.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/270.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/271.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/272.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/273.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/274.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/275.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/277.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/278.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/279.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/280.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/281.tif.

Moffat:1988:PPP

[MG88] D. S. Moffat and P. M. D. Gray. Perlog: a Prolog with persistence and modules. The Computer Journal, 31(2):110– 115, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/2/110.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/110.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/111.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/112.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/113.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/114.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/115.
tif.

Moelwyn-Hughes:1981:BR

[MH81] A. A. Moelwyn-Hughes. Book review. The Computer Journal, 24 (1):77, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/1/77.1.full.pdf+html.

Muir:1985:ALN

[MHS85] S. Muir, D. Hutchison, and D. Shepherd. ARCA: a local network file server. The Computer Journal, 28(3):243– 249, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/243.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/243.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/244.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/245.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/246.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/247.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/248.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/249.tif.

Michie:1980:ES

[Mic80] D. Michie. Expert systems. The Computer Journal, 23(4):369– 376, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/369.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/370.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/371.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/373.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/374.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/375.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/376.tif. See correspondence

[BM81].

Michie:1981:TEC

[Mic81] Donald Michie. A theory of evaluative comments in chess with a note on minimaxing. The Computer Journal, 24(3):278–286, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/278.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/278.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/279.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/280.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/281.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/282.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/283.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/284.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/285.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/286.tif.

Michie:1982:EMG

[Mic82] Donald Michie. Experiments on the mechanization of gamelearning. II. Rule-based learning and the human window. The Computer Journal, 25(1):105–113, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/1/105.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/105.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/106.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/107.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/108.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/109.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/110.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/111.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/113.tif.

Middleton:1987:STR

[Mid87] A. G. Middleton. Some transformations of relevance to report generation. The Computer Journal, 30(1):37–42, February 1987. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/30/1/37.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/38.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_01/tiff/40.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
41.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/42.tif.

Millington:1981:SSA

[Mil81a] Donald Millington. Structured systems analysis and design using standard flowcharting symbols. The Computer Journal, 24(4):295– 300, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/295.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/295.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/296.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/297.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/298.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/299.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/300.
tif.

Milne:1981:FIS

[Mil81b] Wendy J. Milne. A framework for the investigation of a spatial data base. The Computer Journal, 24(1):52–55, February 1981. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/24/1/52.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/53.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/55.tif.

Miller:1982:BRW

[Mil82a] C. D. F. Miller. Book review: E. W. Martin and W. C. Perkins, Fortran for Business Students: a Programmed Instruction Approach. The Computer Journal, 25(4):492, November 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/4/492.
full.pdf+html.

Miller:1982:PPP

[Mil82b] C. D. F. Miller. Programming in Pascal with Pascal/1000. The Computer Journal, 25(1):160a–160, February 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/1/160-
a.full.pdf+html.

Mingers:1988:CCM

[Min88] J. Mingers. Comparing conceptual models and data flow diagrams. The Computer Journal, 31(4):376–379, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/31/4/376.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/376.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/377.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/378.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/379.tif.

Mitra:1982:CSP

[Mit82] G. Mitra. Computer scheduling of public transport. The Computer Journal, 25(3):396a–396, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/3/396-a.full.pdf+html.

Mitchell:1989:APS

[Mit89a] C. Mitchell. Another postage stamp problem. The Computer Journal, 32(4):374–376, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/4/374.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/374.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/375.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/376.tif.

Mitchell:1989:Ca

[Mit89b] C. Mitchell. Correspondence. The Computer Journal, 32(4):379– 380, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/379.2.full.pdf+html.

Mitchell:1989:MDS

[Mit89c] C. Mitchell. Multi-destination secure electronic mail. The Computer Journal, 32(1):13–15, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/13.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/13.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/15.tif.

Morris:1982:PCA

[MK82] R. J. T. Morris and J. S. Kaufman. Performance comparison of I/O access disciplines for transaction-processing systems. The Computer Journal, 25(1):74–83, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/74.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/74.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/75.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/76.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/78.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/79.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/80.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
81.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/83.tif.

Markusz:1985:CCL

[MK85] Z. Markusz and A. A. Kaposi. Complexity control in logicbased programming. The Computer Journal, 28(5):487–495, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/487.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/487.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/488.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/489.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/490.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/491.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/492.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/493.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_05/tiff/494.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/495.tif.

Manolopoulos:1987:BIS

[MKB87] Y. P. Manolopoulos, J. (Y.) G. Kollias, and F. Warren Burton. Batched interpolation search. The Computer Journal, 30(6):565– 568, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/565.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/565.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/566.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/567.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/568.tif.

Manolopoulos:1986:SVB

[MKH86] Y. Manolopoulos, J. G. Kollias, and M. Hatzopoulos. Sequential vs. binary batched searching. The Computer Journal, 29(4):368– 372, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/368.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/368.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/369.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/370.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/371.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/372.tif.

Manolopoulos:1987:MIF

[MKP87] Y. Manolopoulos, D. Kleftouris, and L. Petrou. A model for an ISAM file with multiple overflow chains. The Computer Journal, 30(6):529–534, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/6/529.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/529.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/530.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/531.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/532.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/533.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/534.tif.

Mariani:1984:VMV

[MM84] John A. Mariani and Douglas R. McGregor. VULCAN: a modified viewdata system with improved indexing and formal hierarchy definition language. The Computer Journal, 27(2):121– 126, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/2/121.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/121.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/122.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/123.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/124.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/125.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/126.
tif.

Mitchell:1989:Cb

[MM89] R. J. Mitchell and P. R. Minchinton. Correspondence. The Computer Journal, 32(4):380, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/4/380.1.full.pdf+html.

Mackenzie:1987:MMS

[MMB87] L. M. Mackenzie, A. M. Macleod, and D. J. Berry. A multiple microprocessor system for CPU-bound calculations. The Computer Journal, 30(2):110–118, April 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/2/110.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/110.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/111.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/113.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/114.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/115.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/116.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/

tiff/117.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/118.tif.

McGregor:1987:AAP

[MMH87] D. McGregor, S. McInnes, and M. Henning. An architecture for associative processing of large knowledge bases (LKBs). The Computer Journal, 30(5):404–412, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/5/404.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/404.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/405.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/406.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/407.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/408.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/409.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/410.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/411.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/412.tif.

Morris:1989:RDT

[Mor89] R. A. Morris. Rendering digital type: a historical and economic view of technology. The Computer Journal, 32(6):524– 532, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/524.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/524.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/525.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/526.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/527.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/528.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/529.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/530.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/531.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/532.tif.

Mossige:1986:C

[Mos86] S. Mossige. Correspondence. The Computer Journal, 29(2):190, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/2/190.3.full.pdf+html.

Mottl:1980:DPN

[Mot80] J. Mottl. Description of a program for nonlinear programming: special problem of bivalent nonlinear programming. The Computer Journal, 23(4):360–368, November 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_04/tiff/360.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_04/tiff/361.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_04/tiff/362.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/363.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/364.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/365.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/367.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/368.tif.

Mottl:1985:DPN

[Mot85] J. Mottl. Description of a program for nonlinear programming: the centroid program. The Computer Journal, 28(1):89–94, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
28/1/89.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
90.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/92.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
93.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/94.tif.

Maude:1985:SRR

[MP85] T. I. Maude and D. J. Pullinger. Software for reading, refereeing and browsing in the BLEND system. The Computer Journal, 28(1):1–4, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/28/1/1.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/1.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/2.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_28/Issue_01/tiff/3.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_28/
Issue_01/tiff/4.tif.

Mancini:1989:PCP

[MP89] L. V. Mancini and G. P. Pappalardo. Proving correctness properties of a replicated synchronous program. The Computer Journal, 32(5):437–444, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/5/437.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/437.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/438.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/439.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/440.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/441.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/442.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/443.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/444.tif.

Missala:1982:CJS

[MR82] Marek Missala and Piotr Rudnicki. Correspondence: Jumping to some purpose. The Computer Journal, 25(2):286, May 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_

25/Issue_02/tiff/286.tif. See [ASG79, Hil80, Rob80a, Tay80].

Marwick:1985:TMC

[MR85] D. H. Marwick and C. R. G. B. Rinaldi. The Tattoo on a micro-computer. The Computer Journal, 28(3):216–222, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/216.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/216.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/217.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/218.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/219.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/220.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/221.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/222.tif.

McKenzie:1983:SNC

[MT83] B. J. McKenzie and T. Takaoka. Short note: a control structure for a variable number of nested loops. The Computer Journal, 26(3): 282–283, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/282.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/282.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/283.tif.

Mullender:1986:DCB

[MT86] S. J. Mullender and A. S. Tanenbaum. The design of a capabilitybased distributed operating system. The Computer Journal, 29(4): 289–299, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/289.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/289.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/290.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/291.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/293.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/294.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/295.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/296.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/297.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/298.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/299.
tif.

Morris:1988:HCI

[MTPL88] D. Morris, C. J. Theaker, R. Phillips, and W. Love. Humancomputer interface recording. The Computer Journal, 31(5):437– 444, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/5/437.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/437.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/438.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/439.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/440.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/441.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/442.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/443.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/444.tif.

Morris:1988:SAS

[MTW88] D. Morris, C. J. Theaker, and P. B. Whitehead. Structured abstract schematics. The Computer Journal, 31(3):193– 200, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/193.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/193.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/195.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/197.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/198.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/199.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/200.tif.

Muchsel:1982:PMP

[Muc82] Reginald Muchsel. Performance measurements — a practical example from a German University Computing Centre. The Computer Journal, 25(2):188–192, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/188.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

25/Issue_02/tiff/188.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/189.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/190.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/191.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/192.
tif.

Mullin:1981:CAB

[Mul81] James K. Mullin. Change area B-trees: a technique to aid error recovery. The Computer Journal, 24(4):367–373, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/367.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/367.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/368.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/369.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/370.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/371.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/373.tif.

Mullin:1985:SSE

[Mul85] James K. Mullin. Spiral storage: Efficient dynamic hashing with constant performance. The Computer Journal, 28(3):330– 334, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/330.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/330.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/331.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/332.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/333.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/334.tif.

Mullin:1987:ATD

[Mul87] J. K. Mullin. Accessing textual documents using compressed indexes of arrays of small Bloom filters. The Computer Journal, 30(4):343–348, August 1987. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/4/343.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/343.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/344.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/345.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/346.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/347.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/348.tif.

Mumford:1980:SAS

[Mum80] Enid Mumford. Social aspects of systems analysis. The Computer Journal, 23(1):5–7, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/1/5.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/5.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/6.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_01/tiff/7.tif.

Mumford:1984:SNP

[Mum84a] E. Mumford. Short note: Participative systems design. The Computer Journal, 27(3):283, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/27/3/283.2.full.pdf+html.

Mumford:1984:PSD

[Mum84b] Enid Mumford. Participative systems design. The Computer Journal, 27(3):283, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_27/Issue_03/tiff/283.tif.

Mumford:1985:DSR

[Mum85] Enid Mumford. Defining system requirements to meet business needs: a case study example. The Computer Journal, 28(2):97– 104, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/2/97.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_02/tiff/100.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
02/tiff/101.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_28/Issue_02/tiff/102.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/97.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/
98.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/99.tif.

Murtagh:1983:SRA

[Mur83] F. Murtagh. A survey of recent advances in hierarchical clustering algorithms. The Computer Journal, 26(4):354–359, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/354.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/354.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/355.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/356.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/357.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/358.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/359.
tif.

Murtagh:1985:SAC

[Mur85] F. Murtagh. A survey of algorithms for contiguity-constrained clustering and related problems. The Computer Journal, 28(1):82– 88, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/1/82.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_28/Issue_01/tiff/82.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_28/
Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/84.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
85.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/87.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
88.tif.

Muxworthy:1981:BR

[Mux81] D. T. Muxworthy. Book review. The Computer Journal, 24(1):77, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/77.3.full.pdf+html.

Muxworthy:1984:BR

[Mux84] D. Muxworthy. Book review. The Computer Journal, 27(3):287, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/3/287.2.full.pdf+html.

Maes:1988:RAI

[MV88] R. Maes and J. E. M. Van Dijk. On the role of ambiguity and incompleteness in the design of decision tables and rulebased systems. The Computer Journal, 31(6):481–489, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/481.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/481.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/482.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/483.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/484.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/485.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/486.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/487.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/488.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/489.tif.

MacCallum:1986:CST

[MZ86] K. J. MacCallum and J.-M. Zhang. Curve-smoothing techniques using B-splines. The Computer Journal, 29(6):564–571, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/6/564.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/564.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/565.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/566.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/567.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_29/Issue_06/tiff/568.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/569.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/570.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/571.tif.

Natarajan:1986:DSS

[Nat86] N. Natarajan. A distributed synchronisation scheme for communicating processes. The Computer Journal, 29(2):109–117, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/109.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/109.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/110.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/111.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/113.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/114.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/115.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/116.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/117.tif.

Navlakha:1987:SSC

[Nav87] J. K. Navlakha. A survey of system complexity metrics. The Computer Journal, 30(3):233–238, June 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/3/233.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/233.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/234.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/235.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/236.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/237.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/238.tif.

Newton:1980:BR

[New80] R. Newton. Book review. The Computer Journal, 23(2):160, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/160.1.full.pdf+html.

Newton:1981:BR

[New81] R. W. Newton. Book review. The Computer Journal, 24(4):384, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/384.1.full.pdf+html.

Napper:1980:RUE

[NF80] R. B. E. Napper and R. N. Fisher. RCC — a user-extensible systems implementation language. The Computer Journal, 23(3): 212–222, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/212.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/212.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/213.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/214.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/215.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/216.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/217.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/218.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/219.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/220.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/221.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/222.
tif.

Ngoh:1989:TPR

[NH89] L. H. Ngoh and T. P. Hopkins. Transport protocol requirements for distributed multimedia information systems. The Computer Journal, 32(3):252–261, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/3/252.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/252.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/253.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/254.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_32/Issue_03/tiff/255.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/256.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/257.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/258.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/259.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/260.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/261.
tif.

Noor:1984:FCM

[Noo84] A. El Sayed Noor. A framework for the creation and management of national computing strategies in developing countries. The Computer Journal, 27(3):193–200, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/27/3/193.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/193.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/195.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/197.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/198.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/199.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/200.tif.

Newton:1981:BRB

[NPCA81] R. W. Newton, F. Panzieri, C. C. Chang, and A. M. Addyman. Book reviews: Programming Language Standardisation, edited by I. D. Hill and B. L. Meek, 1980; Problem Solving and Structured Programming, by E. B. Koffmann, 1981; Data Base: Structured Techniques for Design, Performance and Management, by S. Atre, 1980; Pascal — The Language and its Implementation, edited by D. W. Barron, 1981. The Computer Journal, 24(4):384, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/384.tif.

Negri:1989:SNS

[NPS89] M. Negri, G. Pelagatti, and L. Sbattella. Short notes: Semantics and problems of universal quantification in SQL. The Computer Journal, 32(1):90–91, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/90.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/90.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/91.tif.

Newton:1980:BRB

[NQ80] Roy Newton and Harley Quilliam. Book reviews: Software Portability: an Advanced Course, edited by P. J. Brown, 1977; The Challenge of Microprocessors, by M. G. Hartley and Anne Buckley, 1979. The Computer Journal, 23(2):160, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/160.tif.

Nash:1987:SAP

[NS87] J. C. Nash and S. Shlien. Simple algorithms for the partial singular value decomposition. The Computer Journal, 30(3):268– 275, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/3/268.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/268.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/269.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/270.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/271.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/272.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/273.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/274.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/275.tif.

Newman:1987:HMP

[NSW87] I. A. Newman, R. P. Stallard, and M. C. Woodward. A hybrid multiple processor garbage collection algorithm. The Computer Journal, 30(2):119–127, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/2/119.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/119.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_02/tiff/120.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/121.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/122.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/123.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/124.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/125.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/127.tif.

Nedunuri:1987:DRS

[NW87] S. Nedunuri and N. E. Wiseman. Displaying random surfaces. The Computer Journal, 30(2):163–167, April 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/2/163.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/163.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/164.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/165.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/166.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/167.
tif.

Nygaard:1980:ISM

[Nyg80] K. Nygaard. The impact of social movements. The Computer Journal, 23(1):19–22, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/1/19.full.pdf+html.

OKeefe:1985:CCA

[O’K85] R. M. O’Keefe. Comment on ‘Complexity analyses of event set algorithms’. The Computer Journal, 28(5):496–497, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/496.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/496.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/497.tif. See [Ree84a].

OwrangO:1988:QTB

[OM88] M. M. Owrang O. and L. L. Miller. Query translation based on hypergraph models. The Computer Journal, 31(2):155– 164, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/2/155.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/155.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/156.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/157.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/158.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/159.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/160.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/161.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/162.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/163.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/164.tif.

Orman:1988:FDC

[Orm88] L. Orman. Functional database constraints. The Computer Journal, 31(4):336–343, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/4/336.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/336.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/337.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/338.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/339.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/343.tif.

OShea:1988:CDU

[O’S88] G. O’Shea. Controlling the dependency of user access control mechanisms on correctness of user identification. The Computer Journal, 31(6):503–509, December 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/6/503.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/503.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/504.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/
tiff/505.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/506.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/507.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/508.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/509.tif.

Oulsnam:1982:UUP

[Oul82] G. Oulsnam. Unravelling unstructured programs. The Computer Journal, 25(3):379–387, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/3/379.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/379.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/380.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/381.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/382.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/383.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/384.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/385.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/386.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/387.tif.

Oulsnam:1987:ATS

[Oul87] G. Oulsnam. The algorithmic transformation of schemas to structured form. The Computer Journal, 30(1):43–51, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
30/1/43.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/43.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
44.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/45.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/46.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/

47.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/48.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/49.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
50.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/51.tif.

Ottmann:1980:BTA

[OW80] T. Ottmann and D. Wood. 1 − 2 brother trees or AVL trees revisited. The Computer Journal, 23(3):248–255, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/248.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/248.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/249.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/250.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/251.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/252.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/253.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/255.tif.

Oliver:1983:OQE

[OW83a] M. A. Oliver and N. E. Wiseman. Operations on quadtree encoded images. The Computer Journal, 26(1):83–91, February 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/1/83.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/83.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/84.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/85.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/86.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
87.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
90.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/91.tif.

Oliver:1983:OQL

[OW83b] M. A. Oliver and N. E. Wiseman. Operations on quadtree leaves and related image areas. The Computer Journal, 26(4):375–380, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/375.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/375.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/376.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/377.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/378.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/379.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/380.
tif.

Oxley:1985:SFT

[Oxl85] A. Oxley. Surface fitting by triangulation. The Computer Journal, 28(3):335–339, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/335.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/336.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/338.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/339.
tif.

Oxley:1989:SNC

[Oxl89] A. Oxley. Short note: Calculating earthwords volumes — an exercise in geometric algorithms. The Computer Journal, 32(4):377– 379, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/377.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/377.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/378.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/379.tif.

Pachl:1988:SNT

[Pac88] J. Pachl. Short note: Two livelock-free protocols for slotted rings. The Computer Journal, 31(5):474–476, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/31/5/474.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/474.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/475.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/476.tif.

Pagan:1981:SWS

[Pag81] F. G. Pagan. Style for writing the syntactic portions of complete definitions of programming languages. The Computer Journal, 24(2):143–147, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/143.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/143.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/144.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/145.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/146.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/147.
tif.

Pagan:1982:SSU

[Pag82a] Frank G. Pagan. Semantic specification using two-level grammars: labels and GOTO statements. The Computer Journal, 25(1):63– 67, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/63.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_25/Issue_01/tiff/63.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_25/
Issue_01/tiff/64.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/65.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
66.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/67.tif.

Page:1982:OFA

[Pag82b] I. P. Page. Optimal fit of arbitrary sized segments. The Computer Journal, 25(1):32–33, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/32.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_25/Issue_01/tiff/32.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/33.tif.

Page:1984:ACP

[Pag84] I. P. Page. Analysis of a cyclic placement scheme. The Computer Journal, 27(1):18–26, February 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/27/1/18.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/18.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/19.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/20.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/22.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/23.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/24.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
25.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/26.tif.

Pallo:1986:ERU

[Pal86] J. M. Pallo. Enumerating, ranking and unranking binary trees. The Computer Journal, 29(2):171–175, April 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/2/171.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/171.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/172.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/173.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/174.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/175.
tif.

Pallo:1988:SNS

[Pal88] J. Pallo. Short notes: Some properties of the rotation lattice of binary trees. The Computer Journal, 31(6):564–565, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/564.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/564.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/565.tif.

Panzieri:1981:BR

[Pan81] F. Panzieri. Book review. The Computer Journal, 24(4):384, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/384.2.full.pdf+html.

Papakonstantinou:1982:OEQ

[Pap82] G. Papakonstantinou. Optimal evaluation of queries. The Computer Journal, 25(2):239–241, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/239.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/239.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/240.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/241.tif.

Parkin:1982:DAS

[Par82] A. Parkin. Data analysis and system design by entityrelationship modelling — a practical example. The Computer Journal, 25(4):401–409, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/4/401.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/401.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/402.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/403.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/404.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/405.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/406.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/407.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/408.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/409.tif.

Parkin:1983:BRa

[Par83a] J. E. Parkin. Book review. The Computer Journal, 26(2):191, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/2/191.2.full.pdf+html.

Parkin:1983:BRb

[Par83b] J. E. Parkin. Book review. The Computer Journal, 26(2):191, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/2/191.3.full.pdf+html.

Parr:1988:EAR

[Par88] G. P. Parr. Enhanced address resolution in a multi-LAN Ethernet communications system. The Computer Journal, 31(4):296– 303, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/4/296.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/296.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/297.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/298.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/299.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/300.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/301.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/302.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/303.tif.

Parker:1989:CGC

[Par89] J. R. Parker. Computer graphics: on the conversion of filled polygonal regions from raster to vector representation. The Computer Journal, 32(6):549–553, December 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/6/549.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/549.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/550.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/551.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/552.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/553.
tif.

Paulson:1980:BR

[Pau80a] R. W. Paulson. Book review. The Computer Journal, 23(2):123, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/123.full.pdf+html.

Paulson:1980:BRB

[Pau80b] R. W. Paulson. Book review: An Introduction to Computational Fluid Mechanics, by Chuen-Yen Chow, 1979. The Computer Journal, 23(2):123, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_23/Issue_02/tiff/123.tif.

Paulson:1985:LLL

[Pau85] Lawrence C. Paulson. Lessons learned from LCF: a survey of natural deduction proofs. The Computer Journal, 28(5):474–479, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/474.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/474.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/475.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/476.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/477.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/478.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/479.
tif.

Penny:1986:DRM

[PAW86] J. P. Penny, P. J. Ashton, and A. L. Wilkinson. Data recording and monitoring for analysis of system response times. The Computer Journal, 29(5):396–403, October 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/5/396.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/396.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/397.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/398.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/399.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/400.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

29/Issue_05/tiff/401.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/402.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/403.tif.

Patnaik:1986:TTI

[PB86] L. M. Patnaik and Julie Basu. Two tools for interprocess communication in distributed data-flow systems. The Computer Journal, 29(6):506–521, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/6/506.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/506.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/507.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/508.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/509.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/510.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/511.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/512.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/513.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/514.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/515.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/516.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/517.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/518.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/519.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/520.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/521.tif.

Pitman:1986:BSS

[PBH86] N. M. Pitman, F. W. Burton, and E. W. Haddon. Buddy systems with selective splitting. The Computer Journal, 29(2):127– 134, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/2/127.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/127.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/128.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_29/Issue_02/tiff/129.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_02/
tiff/130.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_02/tiff/131.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_02/tiff/132.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_02/tiff/133.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/134.tif.

Perros:1980:RMP

[Per80] H. G. Perros. A regression model for predicting the response time of a disc I/O system. The Computer Journal, 23(1):34–36, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/34.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/34.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/35.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/36.tif.

PeytonJones:1989:PIF

[Pey89] Simon L. Peyton Jones. Parallel implementations of functional programming languages. The Computer Journal, 32(2):175– 186, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/2/175.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/175.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/176.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/177.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/178.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/179.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/180.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/181.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/182.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/183.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/184.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/185.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/186.tif.

Pramanik:1986:IDM

[PF86] S. Pramanik and F. Fotouhi. Index database machine. The Computer Journal, 29(5):451–457, October 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/5/451.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/451.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/452.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/453.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/454.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/455.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/456.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/457.tif.

Pflug:1984:DMA

[Pfl84] G. Ch. Pflug. Dynamic memory allocation — a Markovian analysis. The Computer Journal, 27(4):328–333, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/328.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/328.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/329.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/330.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/331.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/332.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/333.
tif.

Prather:1981:DFS

[PG81] Ronald E. Prather and Shirla G. Giulieri. Decomposition of flowchart schemata. The Computer Journal, 24(3):258–262, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/258.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/258.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/259.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/260.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/

tiff/261.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/262.tif. See comment [Wil82a].

Pitteway:1982:BAR

[PG82] M. L. V. Pitteway and A. J. R. Green. Bresenham’s algorithm with run line coding shortcut. The Computer Journal, 25(1):114– 115, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/114.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/114.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/115.tif.

Papazoglou:1984:ACP

[PGM84] M. P. Papazoglou, P. I. Georgiadis, and D. G. Maritsas. Architectural considerations of the parallel SIMULA machine. The Computer Journal, 27(3):254–259, August 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/3/254.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/255.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/256.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/257.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/258.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/259.tif.

Phillpotts:1984:C

[Phi84] G. Phillpotts. Correspondence. The Computer Journal, 27(1):90, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/1/90.1.full.pdf+html.

Piessens:1982:AIL

[Pie82] Robert Piessens. Algorithm 113: Inversion of the Laplace transform. The Computer Journal, 25(2):278–282, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_02/tiff/278.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_25/Issue_02/tiff/279.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_25/

Issue_02/tiff/280.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/281.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/282.tif.

Pike:1981:TCU

[Pik81] J. Pike. Text compression using a 4-bit coding scheme. The Computer Journal, 24(4):324–330, November 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/4/324.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/324.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/325.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/326.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/327.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/328.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/329.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/330.tif.

Prowse:1980:NLD

[PJ80] P. H. Prowse and R. G. Johnson. A natural language data base interface to the user. The Computer Journal, 23(1):22–25, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
23/1/22.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/22.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
23.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/25.tif.

Pandya:1986:SDT

[PJ86] P. Pandya and M. Joseph. A structure-directed total correctness proof rule for recursive procedure calls. The Computer Journal, 29(6):531–537, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/6/531.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/531.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/532.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/

tiff/533.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/534.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/535.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/536.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/537.tif.

Page:1989:SME

[PJ89] I. P. Page and R. T. Jacob. The solution of mutual exclusion problems which can be described graphically. The Computer Journal, 32(1):45–54, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/45.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/45.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/46.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/47.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/48.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/49.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/51.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
52.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/54.tif.

Prosser:1983:SNN

[PK83] C. J. Prosser and A. C. Kilgour. Short notes: a note on computing the square root of an integer. The Computer Journal, 26 (2):187, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/187.full.pdf+html.

Pramanik:1985:HPM

[PK85] Sakti Pramanik and Chung-Ta King. A hardware pattern matching algorithm on a dataflow. The Computer Journal, 28(3):264– 269, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/264.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/264.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

28/Issue_03/tiff/265.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/266.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/267.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/268.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/269.
tif.

Papakonstantinou:1986:KRA

[PK86] George K. Papakonstantinou and John Kontos. Knowledge representation with attribute grammars. The Computer Journal, 29(3): 241–245, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/241.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/241.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/242.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/243.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/244.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/245.tif.

Parker:1987:WT

[PKK87] J. Parker, A. Kennard, and D. King. The ‘window’ terminal. The Computer Journal, 30(6):558–564, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/6/558.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/558.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/559.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/560.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/561.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/562.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/563.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/564.tif.

Potter:1981:BR

[Pot81] C. Potter. Book review. The Computer Journal, 24(3):199, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/199.3.full.pdf+html.

Potter:1982:FSA

[Pot82] C. Potter. Fundamentals of systems analysis. The Computer Journal, 25(1):83a–83, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/1/83-a.full.pdf+html.

Prakash:1983:DDF

[PPB83] Naveen Prakash, Parimala N., and N. Bolloju. Data definition facilities in Admin. The Computer Journal, 26(4):329–335, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/329.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/329.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/330.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/331.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/332.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/333.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/334.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/335.tif.

Parimala:1984:NCN

[PPB84a] N. Parimala, Naveen Prakash, and N. Bolloju. New corecs and new cosets in Admin. The Computer Journal, 27(4):310–314, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/310.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/310.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/311.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/312.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/313.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/314.tif.

ParimalaN:1984:CCA

[PPB84b] Parimala N., Naveen Prakash, and N. Bolloju. Concurrency control in Admin. The Computer Journal, 27(1):62–66, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
27/1/62.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_27/Issue_01/tiff/62.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
63.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/64.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/65.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/
66.tif.

Prakash:1984:SIC

[PPB84c] Naveen Prakash, N. Parimala, and N. Bolloju. Specifying integrity constraints in a network DBMS. The Computer Journal, 27(3): 209–217, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/209.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/209.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/210.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/211.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/212.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/213.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/214.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/215.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/216.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/217.tif.

Parimala:1989:QFN

[PPRB89] N. Parimala, N. Prakash, B. L. N. Rao, and N. Bolloju. A query facility to a network DBMS. The Computer Journal, 32(1):55–62, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/55.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/56.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/57.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/58.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/59.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/60.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/

61.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/62.tif.

Prather:1984:ATS

[Pra84] Ronald E. Prather. An axiomatic theory of software complexity measure. The Computer Journal, 27(4):340–347, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/340.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/343.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/344.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/345.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/346.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/347.tif.

Prather:1988:CET

[Pra88] R. E. Prather. Comparison and extension of theories of Zipf and Halstead. The Computer Journal, 31(3):248–252, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/248.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/248.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/249.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/250.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/251.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/252.tif.

Pringle:1981:JFH

[Pri81] Alison M. Pringle. Justification with fewer hyphens. The Computer Journal, 24(4):320–323, November 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/4/320.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/320.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/321.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/322.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/323.tif.

Perrott:1980:SEU

[PRO80a] R. H. Perrott, A. K. Raja, and P. C. O’Kane. A simulation experiment using two languages. The Computer Journal, 23(2): 142–146, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/142.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/142.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/143.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/144.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/145.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/146.tif.

Prowse:1980:DBA

[Pro80b] P. Prowse. The data base approach. The Computer Journal, 23(1): 9–12, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/1/9.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_01/tiff/10.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_01/tiff/11.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/12.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
9.tif.

Prowse:1980:BRa

[Pro80c] P. H. Prowse. Book review. The Computer Journal, 23(2):152, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/152.2.full.pdf+html.

Prowse:1980:BRb

[Pro80d] R. W. Prowse. Book review. The Computer Journal, 23(2):190, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/190.full.pdf+html.

Prosser:1983:NCS

[Pro83] C. J. Prosser. A note on computing the square root of an integer. The Computer Journal, 26(2):187–188, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_02/tiff/187.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_26/Issue_02/tiff/188.tif.

Prakash:1981:SDB

[PSD^{+}81] Naveen Prakash, P. Sadanandan, S. M. Desai, K. Rama Rao, and P. N. Gejji. Samhita — a data base management system. The Computer Journal, 24(1):48–51, February 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/24/1/48.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/48.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/50.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/51.tif.

Page:1981:EAM

[PW81a] I. P. Page and R. T. Wood. Empirical analysis of a moving head disc model with two heads separated by a fixed number of tracks. The Computer Journal, 24(4):339–342, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/339.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/339.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/340.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/341.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/342.tif.

Page:1981:SPS

[PW81b] I. P. Page and R. T. Wood. A shaded-picture scanning attachment for incremental plotters. The Computer Journal, 24(3): 287, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/287.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/287.
tif.

Quilliam:1980:BR

[Qui80] H. Quilliam. Book review. The Computer Journal, 23(2):160, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/160.2.full.pdf+html.

Quilliam:1981:BR

[Qui81] H. Quilliam. Book review. The Computer Journal, 24(2):142, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/142.1.full.pdf+html.

Quinlan:1983:ICA

[Qui83] J. R. Quinlan. INFERNO: a cautious approach to uncertain interference. The Computer Journal, 26(3):255–269, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/255.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/255.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/256.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/257.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/258.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/259.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/260.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/261.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/262.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/263.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/264.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/265.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/266.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/267.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/268.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/269.tif.

Quin:1989:CGD

[QY89] A. Quin and Y. Yanagisawa. Computer graphics: on data compaction of scanning curves. The Computer Journal, 32(6):563– 566, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/563.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/563.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/564.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/565.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/566.tif.

Radford:1981:BR

[Rad81a] A. S. Radford. Book review. The Computer Journal, 24(4):366, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/366.1.full.pdf+html.

Radley:1981:BR

[Rad81b] G. Radley. Book review. The Computer Journal, 24(3):248, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/248.1.full.pdf+html.

Radford:1982:PNT

[Rad82] Peter Radford. Petri net theory and the modeling of systems. The Computer Journal, 25(1):129, February 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/1/129.
full.pdf+html.

Raita:1987:ASF

[Rai87] T. Raita. An automatic system for file compression. The Computer Journal, 30(1):80–86, February 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/30/1/80.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_30/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_
01/tiff/82.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/83.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/84.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

30/Issue_01/tiff/85.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/86.tif.

Ramsay:1985:ECS

[Ram85] Allan Ramsay. On efficient context switching. The Computer Journal, 28(4):375–378, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/4/375.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/375.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/376.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/378.tif.

Randell:1986:SDS

[Ran86] B. Randell. System design and structuring. The Computer Journal, 29(4):300–306, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/4/300.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/300.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/301.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/302.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/303.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/304.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/305.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/306.tif.

Raymont:1980:ASS

[Ray80] P. G. Raymont. Automatic speech and speaker recognition. The Computer Journal, 23(3):242, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/3/242.
full.pdf+html.

Raymont:1982:A

[Ray82] P. G. Raymont. Apl80. The Computer Journal, 25(3):337b–337, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/25/3/337-b.full.pdf+html.

Robinson:1985:DDS

[RB85] J. Robinson and A. Burns. A dialogue development system for the design and implementation of user interfaces in Ada. The Computer Journal, 28(1):22–28, February 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/28/1/22.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/24.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/28.tif.

Rom:1986:NTA

[RB86] M. Rom and S. Bergman. A new technique for automatic contouring and contour representation from machine-readable spatial data. The Computer Journal, 29(5):467–471, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/467.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/467.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/468.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/469.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/470.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/471.tif.

Riddell:1980:CRS

[RD80] I. J. Riddell and L. M. Delves. The comparison of routines for solving Fredholm integral equations of the second kind. The Computer Journal, 23(3):274–285, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/23/3/274.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/274.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/275.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/276.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/277.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/278.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/279.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/280.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/281.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/282.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/283.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/284.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/285.tif.

Rabel:1980:SNN

[RDT80] M. Rabel, D. Dubus, and A. Tosser. Saving of NAND/NOR gates by inhibition or d-inhibition. The Computer Journal, 23(2):165– 185, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/165.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/165.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/166.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/167.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/168.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/169.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/170.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/171.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/172.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/173.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/174.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/175.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/176.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/177.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/178.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/179.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/180.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/181.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_23/Issue_02/tiff/182.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/183.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/184.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/185.
tif.

Rechenberg:1984:C

[Rec84] P. Rechenberg. Correspondence. The Computer Journal, 27(4): 378, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/378.1.full.pdf+html.

Rees:1980:BR

[Ree80a] D. H. Rees. Book review. The Computer Journal, 23(1):84, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/84.1.full.pdf+html.

Reeves:1980:FSD

[Ree80b] C. M. Reeves. Free store distribution under random fit allocation: Part 2. The Computer Journal, 23(4):298–306, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_04/tiff/298.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_04/tiff/299.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_04/tiff/300.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/301.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/302.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/303.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/304.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/305.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/306.tif.

Reeves:1980:CUD

[Ree80c] R. Reeves. Correspondence: Using data abstraction for logical design. The Computer Journal, 23(4):381–382, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_04/tiff/381.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_23/Issue_04/tiff/382.tif.

See [Wel80a, JL80].

Reeves:1981:BR

[Ree81] C. M. Reeves. Book review. The Computer Journal, 24(2):190, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/190.2.full.pdf+html.

Reeves:1983:FSD

[Ree83] C. M. Reeves. Free store distribution under random fit allocation: Part 3. The Computer Journal, 26(1):25–35, February 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/1/25.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/25.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/28.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
29.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/30.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/31.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
32.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/34.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
35.tif.

Reeves:1984:CAE

[Ree84a] C. M. Reeves. Complexity analyses of event set algorithms. The Computer Journal, 27(1):72–79, February 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/27/1/72.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/76.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_01/tiff/78.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_27/Issue_01/tiff/

79.tif. See comment [O’K85].

Reeves:1984:LSM

[Ree84b] C. M. Reeves. A lumped-state model of clustering in dynamic storage allocation. The Computer Journal, 27(2):135– 142, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/2/135.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/135.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/136.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/137.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/139.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/140.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/141.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/142.tif.

Reeves:1984:C

[Ree84c] J. Reeves. Correspondence. The Computer Journal, 27(1):90–91, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/1/90.3.full.pdf+html.

Reeves:1987:CMF

[Ree87] C. M. Reeves. Chainmail: a model of first-fit memory allocation. The Computer Journal, 30(3):201–213, June 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/3/201.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/201.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/202.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/203.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/204.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/205.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

30/Issue_03/tiff/206.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/207.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/208.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/209.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/210.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/211.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/212.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/213.tif.

Rees:1988:TIP

[Ree88] R. T. O. Rees. Termination indicators in programming. The Computer Journal, 31(4):380, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/31/4/380.full.pdf+html.

Reingold:1981:CEP

[Rei81] Edward M. Reingold. A comment on the evaluation of Polish postfix expressions. The Computer Journal, 24(3):288, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/288.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/288.
tif.

Reynolds:1983:SNH

[Rey83] R. A. Reynolds. Short note: Human performance in interactive graphics operations. The Computer Journal, 26(1):93–94, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/1/93.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/93.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/94.tif.

Rohl:1986:RP

[RG86] J. S. Rohl and T. D. Gedeon. The Reve’s puzzle. The Computer Journal, 29(2):187–188, April 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/2/187.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

29/Issue_02/tiff/187.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_02/tiff/188.tif.

Robison:1989:CGE

[RHS89] A. D. Robison, B. J. Hafner, and S. S. Skiena. Computer games: Eight pieces cannot cover a chess board. The Computer Journal, 32(6):567–570, December 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/6/567.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/567.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/568.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/569.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/570.tif.

Richards:1988:WPA

[Ric88] D. Richards. On the worst-possible analysis of weighted comparison-based algorithms. The Computer Journal, 31(3):276– 278, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/276.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/276.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/277.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/278.tif.

Riley:1981:BR

[Ril81] N. Riley. Book review. The Computer Journal, 24(1):41, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/1/41.2.full.pdf+html.

Ritchie:1989:HMT

[Rit89] I. Ritchie. Hypertext — moving towards large volumes. The Computer Journal, 32(6):516–523, December 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/6/516.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/516.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/517.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/518.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_32/Issue_06/tiff/519.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/520.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/521.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/522.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/523.tif.

Ramamohanarao:1982:DHS

[RL82] K. Ramamohanarao and John W. Lloyd. Dynamic hashing schemes. The Computer Journal, 25(4):478–485, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/478.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/478.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/479.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/480.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/481.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/482.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/483.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/484.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/485.tif.

Raj:1989:CMS

[RL89] R. K. Raj and H. M. Levy. A compositional model for software reuse. The Computer Journal, 32(4):312–322, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/312.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/312.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/314.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/315.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/316.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/317.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/318.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/319.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/

tiff/320.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/321.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/322.
tif.

Rahman:1989:EBS

[RM89] M. Rahman and S. Misbahuddin. Effect of a binary symmetric channel on the synchronisation recovery of variable length codes. The Computer Journal, 32(3):246–251, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/32/3/246.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/246.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/247.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/248.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/249.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/250.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/251.
tif.

Robinson:1980:CJS

[Rob80a] G. L. Robinson. Correspondence: Jumping to some purpose. The Computer Journal, 23(3):288, August 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_

23/Issue_03/tiff/288.tif. See [ASG79, Hil80, Tay80, MR82].

Robinson:1980:PMC

[Rob80b] S. K. Robinson. A programming metholodology in compiler construction Part I: Concepts. The Computer Journal, 23(3):211, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/3/211.full.pdf+html.

Robinson:1983:C

[Rob83] G. L. Robinson. Correspondance. The Computer Journal, 26 (1):95, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/1/95.3.full.pdf+html.

Roberts:1988:FSQ

[Rob88] W. T. Roberts. A formal specification of the QMC message system: the underlying abstract model. The Computer Journal, 31(4):313– 324, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/4/313.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/313.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/314.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/315.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/316.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/317.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/318.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/319.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/320.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/321.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/322.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/323.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/324.tif.

Rohl:1987:THD

[Roh87] J. S. Rohl. Towers of Hanoi: the derivation of some iterative versions. The Computer Journal, 30(1):70–76, February 1987. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/30/1/70.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_01/tiff/70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/71.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/73.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
74.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/75.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/76.tif.

Rohl:1988:C

[Roh88] J. S. Rohl. Correspondence. The Computer Journal, 31(2):190, ???? 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/31/2/190.3.full.pdf+html.

Ross:1981:BRa

[Ros81a] J. M. Ross. Book review. The Computer Journal, 24(1):70, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/70.full.pdf+html.

Ross:1981:BRb

[Ros81b] J. M. Ross. Book review. The Computer Journal, 24(3):199, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/3/199.2.full.pdf+html.

Rosenquist:1982:ELC

[Ros82] C. J. Rosenquist. Entity life cycle models and their applicability to information systems development life cycles, a framework for information systems design and implementation. The Computer Journal, 25(3):307–315, August 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/3/307.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/307.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/308.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/309.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/310.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/311.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/312.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/313.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/314.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/315.tif.

Rosenne:1987:C

[Ros87] J. Rosenne. Correspondence. The Computer Journal, 30(1):93, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/30/1/93.1.full.pdf+html.

Redish:1987:EMP

[RS87] K. A. Redish and W. F. Smyth. Evaluating measures of program quality. The Computer Journal, 30(3):228–232, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/3/228.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/228.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/229.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/230.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/231.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/232.tif.

Razaz:1988:TPM

[RS88] M. Razaz and J. L. Schonfelder. Test procedures for measurement of floating-point characteristics of computing environments. The Computer Journal, 31(1):12–16, February 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/31/1/12.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/12.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_31/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_
01/tiff/14.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/16.
tif.

Ramshaw:1987:IRP

[RSS87] M. Ramshaw, A. P. Steward, and D. Stoves. An intelligent retrieval and presentation system for management information. The Computer Journal, 30(6):485–486, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/485.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/485.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/486.tif.

Rushby:1986:KEA

[Rus86] N. Rushby. A knowledge engineering approach to instructional design. The Computer Journal, 29(5):385–389, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/5/385.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/385.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/386.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/387.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_05/
tiff/388.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_05/tiff/389.tif.

Ruskey:1989:C

[RvB89] F. Ruskey and D. R. van Baronaigien. Correspondence. The Computer Journal, 32(3):282–283, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/3/282.2.full.pdf+html.

Rennhackkamp:1988:SNU

[RvS88] M. H. Rennhackkamp and S. H. von Solms. Short notes: on using an abstract model of distributed database concurrency control methods. The Computer Journal, 31(6):570–572, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/570.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/570.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/571.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/572.tif.

Rzevski:1985:CAI

[Rze85] George Rzevski. On criteria for assessing an information theory. The Computer Journal, 28(3):200–202, July 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/200.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/201.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/202.tif.

Stecher:1981:ACC

[SA81] P. Stecher and Volker Allenstein. The application controller concept: a first experience. The Computer Journal, 24(2):97–106, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/2/97.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/100.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/102.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/105.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/106.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/97.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/98.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/99.tif.

Suraweera:1988:AMA

[SAA88] F. Suraweera and J. M. Al-Anzy. Analysis of a modified address calculation sorting algorithm. The Computer Journal, 31(6):561– 563, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/6/561.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/561.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/562.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_06/tiff/563.tif.

Salomon:1987:SND

[Sal87] D. Salomon. Short note: a design for an efficient NOR-gate only, binary-ripple adder with carry-completion-detection logic. The Computer Journal, 30(3):283–285, June 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/3/283.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/284.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/285.tif.

Samson:1981:HTC

[Sam81] W. B. Samson. Hash table collision handling on storage devices with latency. The Computer Journal, 24(2):130–131, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/130.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/130.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/131.tif.

Subieta:1988:SNO

[SAM88] K. Subieta, K. Anacki, and M. Missala. Short note: Overlay clusters for C programs. The Computer Journal, 31(3):285– 286, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/285.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/285.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/286.tif.

Stephenson:1981:CIB

[SB81] J. Stephenson and G. Brown. Correspondence: Implementation of BEDSOCS: An interactive simulation language. The Computer Journal, 24(2):191–192, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/24/2/191.2.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
02/tiff/191.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/192.tif. See [ER80].

Samson:1985:ROD

[SB85] W. B. Samson and A. Bendell. Rank order distributions and secondary key indexing. The Computer Journal, 28(3):309– 312, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/309.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/309.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/310.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/311.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/312.tif.

Sentance:1981:DES

[SC81] W. A. Sentance and I. P. Cliff. The determination of eigenvalues of symmetric quindiagonal matrices. The Computer Journal, 24(2):

177–179, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/177.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/177.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/178.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/179.tif.

Sernadas:1987:RCM

[SC87] Cristina Sernadas and R. Carapuça. The role of conceptual modelling abstractions in compiler development. The Computer Journal, 30(4):328–336, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/4/328.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/329.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/330.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/331.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/332.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/333.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/334.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/335.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/336.tif.

Scarrott:1985:ILB

[Sca85] G. G. Scarrott. Information, the life blood of organisation. The Computer Journal, 28(3):203–205, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/203.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/203.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/204.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/205.tif.

Scarrott:1989:NI

[Sca89] G. G. Scarrott. The nature of information. The Computer Journal, 32(3):262–266, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/3/262.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/262.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/263.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/264.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/265.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/266.
tif.

Schagen:1982:ACS

[Sch82a] I. P. Schagen. Automatic contouring from scattered data points. The Computer Journal, 25(1):7–11, February 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/1/7.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/10.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/11.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/7.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
9.tif.

Schmidt:1982:SDS

[Sch82b] B. Schmidt. Simulation of discrete systems using GPSS-Fortran. The Computer Journal, 25(1):84–86, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/84.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/84.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/85.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/86.tif.

Schlorer:1983:ILP

[Sch83] Jan Schlörer. Information loss in partitioned statistical databases. The Computer Journal, 26(3):218–223, August 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/26/3/218.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/218.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/219.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

26/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/221.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/222.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/223.tif.

Schneider:1985:WDT

[Sch85] M. L. Schneider. Weighted decision tables — an alternative solution for ambiguity. The Computer Journal, 28(4):366–371, August 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/4/366.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/367.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/368.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/369.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/370.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/371.
tif.

Scoltock:1982:SLC

[Sco82] Julie Scoltock. A survey of the literature of cluster analysis. The Computer Journal, 25(1):130–134, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/1/130.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/131.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/132.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/133.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/134.
tif.

Sacks-Davis:1982:ARN

[SD82] R. Sacks-Davis. Applications of redundant number representations to decimal arithmetic. The Computer Journal, 25(4):471– 477, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/471.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/471.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

25/Issue_04/tiff/472.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/473.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/474.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/475.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/476.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/477.tif.

Smith:1986:GCC

[SD86] G. W. Smith and J. B. H. Du Boulay. The generation of cryptic crossword clues. The Computer Journal, 29(3):282– 284, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/282.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/282.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/284.tif.

Shanehchi:1982:MBA

[SE82] J. Shanehchi and D. J. Evans. A modified bisection algorithm for the determination of the eigenvalues of a symmetric quindiagonal matrix. The Computer Journal, 25(4):493–494, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/493.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/493.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/494.tif.

Segev:1987:OTW

[Seg87] A. Segev. Optimising two-way joins in horizontally partitioned database systems. The Computer Journal, 30(5):458– 468, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/5/458.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/458.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/459.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/460.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/461.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_30/Issue_05/tiff/462.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/463.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/464.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/465.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_05/
tiff/466.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_05/tiff/467.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_05/tiff/468.
tif.

Selim:1988:SVV

[Sel88] S. M. Selim. Split vertices in vertex colouring and their application in developing a solution to the faculty timetable problem. The Computer Journal, 31(1):76–82, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/31/1/76.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_31/Issue_01/tiff/76.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_31/Issue_01/tiff/77.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_
01/tiff/78.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_01/tiff/79.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/80.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/81.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/82.tif.

Sideri:1989:SNS

[SEPK89] M. Sideri, S. Efremidis, and G. Papa-Konstantinou. Short notes: Semantically driven parsing of context-free languages. The Computer Journal, 32(1):91–93, February 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/32/1/91.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/92.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/93.tif.

Sernadas:1981:SSSa

[Ser81a] Amilcar Sernadas. SYSTEMATICS: its syntax and semantics as a query language. The Computer Journal, 24(1):56–61, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/1/56.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/56.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
57.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/59.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
60.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/61.tif.

Sernadas:1981:SSSb

[Ser81b] Amilcar Sernadas. SYSTEMATICS: Its syntax and semantics as a query language (2). The Computer Journal, 24(2):125– 129, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/125.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/125.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/126.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/127.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/128.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/129.tif.

Shafibegly-Gray:1982:DFS

[SGW82] A. Shafibegly-Gray and R. W. Whitty. Decomposition of flowchart schemata. The Computer Journal, 25(4):495, November 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/25/4/495.
full.pdf+html.

Shand:1981:BRa

[Sha81a] I. M. C. Shand. Book review. The Computer Journal, 24(1):47, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/47.1.full.pdf+html.

Shand:1981:BRb

[Sha81b] I. M. C. Shand. Book review. The Computer Journal, 24(3): 199, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/199.1.full.pdf+html.

Shave:1981:EFB

[Sha81c] M. J. R. Shave. Entities, functions, and binary relations: Steps to a conceptual schema. The Computer Journal, 24(1):42–47, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
24/1/42.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/42.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
43.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
46.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/47.tif.

Shackel:1982:BSP

[Sha82a] B. Shackel. The BLEND system: programme for the study of some ‘electronic journals’. The Computer Journal, 25(2):161– 168, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/2/161.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/161.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/165.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/166.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/167.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/168.tif.

Shand:1982:BRF

[Sha82b] I. M. C. Shand. Book review: F. F. Kuo (editor), Protocols and Techniques for Data Communication Networks. The Computer Journal, 25(4):494, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/4/494.full.pdf+html.

Shackelton:1985:C

[Sha85] P. Shackelton. Correspondence. The Computer Journal, 28(1):95, ???? 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/1/95.1.full.pdf+html.

Shackel:1986:IMU

[Sha86] B. Shackel. IBM makes usability as important as functionality. The Computer Journal, 29(5):475–476, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/29/5/475.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_05/tiff/475.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/476.tif.

Shebini:1985:FFE

[She85] Nader Shebini. FEDOS: a Finite Element Database of Offshore Structures. The Computer Journal, 28(5):498–500, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/5/498.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/498.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/499.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/500.tif.

Shields:1985:CM

[Shi85] M. W. Shields. Concurrent machines. The Computer Journal, 28(5):449–465, November 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/28/5/449.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
05/tiff/449.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/450.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/451.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/452.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/453.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/454.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/455.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/456.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/457.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/458.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/459.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/460.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/461.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_05/tiff/462.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/463.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/
tiff/464.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/465.tif.

Shields:1989:ISS

[Shi89] M. W. Shields. Implicit system specification and the interface equation. The Computer Journal, 32(5):399–412, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/5/399.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/399.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/400.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/401.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/402.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/403.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/404.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/405.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/406.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/407.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/408.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/409.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/410.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/411.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/412.tif.

Short:1981:PDT

[Sho81] K. W. Short. Protection in data type abstractions using constraints on data values. The Computer Journal, 24(2):118– 124, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/2/118.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/118.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/119.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/120.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/121.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/122.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/123.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/124.tif.

Sarda:1980:DCA

[SI80] N. L. Sarda and J. R. Isaac. Discussion and correspondence: Automatic analysis and design of information processing systems. The Computer Journal, 23(1):90–94, February 1980. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/23/1/90.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/90.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/91.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/92.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/93.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/94.
tif.

Silberschatz:1981:PDC

[Sil81] A. Silberschatz. Port directed communication. The Computer Journal, 24(1):78–82, February 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/24/1/78.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_24/Issue_01/tiff/78.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/79.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/80.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/82.
tif.

Silberschatz:1982:PQS

[Sil82] Abraham Silberschatz. Priority and queuing specification in ‘distributed processes’. The Computer Journal, 25(1):34–36, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/1/34.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/34.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
35.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/36.tif.

Simpson:1981:BR

[Sim81] D. Simpson. Book review. The Computer Journal, 24(2):172, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/172.2.full.pdf+html.

Sipala:1988:EIB

[Sip88] P. Sipala. Efficient implementation of binary trees in LISP systems. The Computer Journal, 31(4):330–335, August 1988. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/31/4/330.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/330.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/331.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/332.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/333.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/334.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/335.tif.

Santoro:1985:LIR

[SK85] Nicola Santoro and Ramez Khatib. Labelling and implicit routing in networks. The Computer Journal, 28(1):5–8, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/1/5.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_01/tiff/5.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/6.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/7.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_28/Issue_01/tiff/8.tif.

Sklenar:1981:AS

[Skl81] I. Sklenář. Affine slices. The Computer Journal, 24(4):374–377, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/374.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/374.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/375.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/376.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/377.tif.

Styne:1985:PSR

[SKW85] B. A. Styne, T. R. King, and N. E. Wiseman. Pad structures for the Rainbow Workstation. The Computer Journal, 28(1):68–72, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/28/1/68.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_01/tiff/68.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/69.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/71.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/72.tif.

Smith:1983:IUF

[SLH83] A. M. R. Smith, L. S. Lee, and D. J. Hand. Interactive user-friendly interfaces to statistical packages. The Computer Journal, 26(3): 199–204, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/199.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/199.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/201.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/203.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/204.
tif.

Stojmenovic:1988:ASG

[SM88] I. Stojmenović and M. Miyakawa. Applications of a subsetgenerating algorithm to base enumeration, knapsack and minimal covering problems. The Computer Journal, 31(1):65–70, February 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
31/1/65.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/65.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
66.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/67.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_01/tiff/68.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_01/tiff/
69.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_01/tiff/70.tif.

Small:1981:BR

[Sma81] R. E. Small. Book review. The Computer Journal, 24(2):106, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/106.2.full.pdf+html.

Small:1988:IES

[Sma88] Carol Small. The implementation of the Exegesis system. The Computer Journal, 31(2):125–132, April 1988. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/2/125.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/125.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/126.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/127.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/128.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/129.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/131.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/132.tif.

Smith:1983:XCA

[Smi83] P. D. Smith. XENO: Computer-assisted compilation of crossword puzzles. The Computer Journal, 26(4):296–302, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/296.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/296.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/297.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/298.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/299.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/300.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/301.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/302.tif.

Smith:1986:ISP

[Smi86] Joan M. Smith. The implications of SGML for the preparation of scientific publications. The Computer Journal, 29(3):193– 200, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/193.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/193.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/195.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/197.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/198.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/199.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/200.tif.

Snell:1980:CEP

[Sne80] J. Snell. Correspondence: On “An efficient predictor-corrector algorithm”. The Computer Journal, 23(4):382, November 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/382.tif. See [Wes80a, Ano80c].

Sowa:1987:MSS

[Sow87] M. Sowa. A method for speeding up serial processing in dataflow computers by means of a program counter. The Computer Journal, 30(4):289–294, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/4/289.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/289.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_30/Issue_04/tiff/290.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/291.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/292.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/293.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/294.tif.

Skordalakis:1982:CSV

[SP82] E. Skordalakis and G. Papakonstantinou. A control structure for a variable number of nested loops. The Computer Journal, 25(1):48–51, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/48.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/48.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/50.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/51.tif.

Shepherd:1989:FSB

[SPC89] M. A. Shepherd, W. J. Phillips, and C.-K. Chu. A fixedsize Bloom filter for searching textual documents. The Computer Journal, 32(3):212–219, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/3/212.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/212.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/213.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/214.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/215.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/216.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/217.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/218.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/219.tif.

Spivey:1983:LP

[Spi83] J. M. Spivey. Logic programming. The Computer Journal, 26(4):383a–383, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/4/383-a.full.pdf+html.

Shackel:1983:BLP

[SPMD83] B. Shackel, D. J. Pullinger, T. I. Maude, and W. P. Dodd. The BLEND-LINC project on ‘electronic journals’ after two years. The Computer Journal, 26(3):247–254, August 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/3/247.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/247.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/248.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/249.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/250.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/251.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/252.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/253.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/254.tif.

Spoletini:1985:AGP

[Spo85] Enrico Spoletini. Algorithm 122: Generation of permutations in direct lexicographic ordering by arithmetic methods. The Computer Journal, 28(4):444–445, November 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/4/444.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/444.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/445.tif.

Spurr:1988:DDA

[Spu88] Kathleen Spurr. A data dictionary approach to teaching information systems analysis. The Computer Journal, 31(2):182– 189, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/2/182.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/182.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/183.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/184.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/

tiff/185.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/186.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/187.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/188.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/189.tif.

Schmidt:1980:TCA

[SS80] G. Schmidt and T. Ströhlein. Timetable construction — an annotated bibliography. The Computer Journal, 23(4):307–316, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/307.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/308.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/309.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/310.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/311.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/312.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/
tiff/313.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_04/tiff/314.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_04/tiff/315.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_04/tiff/316.tif.

Smith:1981:PCC

[SS81] P. D. Smith and S. Y. Steen. A prototype crossword compiler. The Computer Journal, 24(2):107–111, May 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/107.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/107.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/108.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/109.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/110.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/111.
tif.

Schiper:1989:EIR

[SSDS89] A. Schiper, R. Simon, Ph. Desarzens, and J.-A. Sengstag. Efficient implementation of rendezvous. The Computer Journal, 32(3):

267–272, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/3/267.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/267.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/268.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/269.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/270.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/271.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/272.
tif.

Sibson:1981:SQE

[ST81] Robin Sibson and Graeme D. Thomson. A seamed quadratic element for contouring. The Computer Journal, 24(4):378–382, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/4/378.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_04/tiff/378.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_04/tiff/379.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_04/tiff/380.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_04/
tiff/381.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_04/tiff/382.tif.

Symons:1982:SPA

[ST82] C. R. Symons and P. Tijsma. A systematic and practical approach to the definition of data. The Computer Journal, 25(4):410–422, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/410.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/410.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/411.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/412.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/413.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/414.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/415.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/416.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/417.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/418.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/419.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/420.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/421.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/422.tif.

Sommerville:1989:ASS

[ST89] I. Sommerville and R. Thomson. An approach to the support of software evolution. The Computer Journal, 32(5):386–398, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/5/386.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/386.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/387.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/388.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/389.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/390.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/391.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/392.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/393.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/394.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/395.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/396.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/397.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/398.tif.

Stamper:1985:TTI

[Sta85] R. K. Stamper. Towards a theory of information. information: Mystical fluid or a subject for scientific enquiry? The Computer Journal, 28(3):195–199, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/195.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/195.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/196.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/

tiff/197.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/198.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/199.
tif.

Stewart:1986:QSS

[Ste86] I. P. Stewart. Quadtrees: Storage and scan conversion. The Computer Journal, 29(1):60–75, February 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/29/1/60.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/60.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/61.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/62.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/63.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/64.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/65.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/66.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
67.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/68.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/69.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
70.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/71.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/72.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
73.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/74.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/75.tif.

Stone:1986:CGP

[Sto86] R. G. Stone. On the choice of grammar and parser for the compact analytical encoding of programs. The Computer Journal, 29(4): 307–314, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/307.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/307.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/308.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/309.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/

tiff/310.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/311.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/312.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/314.tif.

Stubbs:1984:ERS

[Stu84] Michael Stubbs. An examination of the resolution of structure clashes by structure inversion. The Computer Journal, 27(4):354– 361, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/354.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/354.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/355.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/356.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/357.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/358.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/359.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/360.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/361.tif.

Subieta:1985:SMD

[Sub85] Kazimierz Subieta. A simple method of data correction. The Computer Journal, 28(4):372–374, August 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/4/372.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/372.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/373.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/374.tif.

Suraweera:1986:UDC

[Sur86] F. Suraweera. Use of doubly chained tree structures in file organisation for optimal searching. The Computer Journal, 29(1):52–59, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/1/52.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/52.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/55.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/56.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/57.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/
58.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_01/tiff/59.tif.

Smeaton:1983:REQ

[SvR83] A. F. Smeaton and C. J. van Rijsbergen. The retrieval effects of query expansion on a feedback document retrieval system. The Computer Journal, 26(3):239–246, August 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/3/239.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/239.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/240.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/241.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/242.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/243.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/244.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/245.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/246.tif.

Slape:1983:CFA

[SW83] John K. Slape and Peter J. L. Wallis. Conversion of Fortran to Ada using an intermediate tree representation. The Computer Journal, 26(4):344–353, November 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/26/4/344.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_04/tiff/344.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/345.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/346.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/347.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/348.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

26/Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/350.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/
tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/352.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/353.
tif.

Six:1984:SRA

[SW84] H. W. Six and L. Wegner. Sorting a random access file in situ. The Computer Journal, 27(3):270–275, August 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/3/270.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/270.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/271.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/272.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/273.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/274.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/275.tif.

Sommerville:1987:DSD

[SWB87] I. Sommerville, R. Welland, and S. Beer. Describing software design methodologies. The Computer Journal, 30(2):128– 133, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/2/128.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/128.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/129.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/130.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/131.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/132.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/133.
tif.

Takaoka:1986:SNW

[Tak86] Tadao Takaoka. The semantics of new while loop. The Computer Journal, 29(1):33–35, February 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/29/1/33.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/34.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/35.tif.

Talbot:1984:ILO

[Tal84] Steve Talbot. An investigation into logical optimization of relational query languages. The Computer Journal, 27(4):301–309, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/4/301.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/301.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/302.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/303.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/304.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/305.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/306.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/307.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/308.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/309.tif.

Tsanakas:1989:FBD

[TAP89] P. Tsanakas, N. Alexandridis, and G. Papakonstantinou. An FPbased design methodology for problem-oriented architectures. The Computer Journal, 32(5):453–460, October 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/5/453.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/453.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/454.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/455.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/456.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/457.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/458.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/459.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/
tiff/460.tif.

Taylor:1980:CJS

[Tay80] N. B. Taylor. Correspondence: Jumping to some purpose. The Computer Journal, 23(2):189–190, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_02/tiff/189.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_02/tiff/190.tif.

See [ASG79, Hil80, Rob80a, MR82].

Taylor:1981:CGP

[Tay81] D. J. Taylor. Correspondence: GOTOs in practice. The Computer Journal, 24(1):95, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_24/Issue_01/tiff/95.tif.

Thomas:1982:CDD

[TB82] R. C. Thomas and A. Burns. The case for distributed decision making systems. The Computer Journal, 25(1):148–152, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/148.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/150.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/151.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/152.tif.

Taylor:1986:LCB

[TB86] D. J. Taylor and J. P. Black. A locally correctable Btree implementation. The Computer Journal, 29(3):269–276, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/269.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/270.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/271.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/

tiff/272.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/273.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/274.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/275.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/276.tif.

Tan:1989:ADI

[TCL89] K. P. Tan, T. S. Chua, and P. T. Lee. AUTO-DFD: an intelligent data flow processor. The Computer Journal, 32(3):194– 201, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/3/194.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/194.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/195.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/196.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/197.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/198.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/199.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/201.tif.

Teare:1982:E

[Tea82] K. P. Teare. Eurographics 80. The Computer Journal, 25(1):129a– 129, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/129-a.full.pdf+html.

Teufel:1987:SDR

[Teu87] B. Teufel. System design for the remote execution of library routines. The Computer Journal, 30(3):254–257, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/3/254.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/255.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/256.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/257.tif.

Tsutsui:1987:DPP

[TF87] S. Tsutsui and Y. Fujimoto. Deadlock prevention in process control computer systems. The Computer Journal, 30(1):20–26, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
30/1/20.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
21.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/23.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
24.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_01/tiff/25.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_01/tiff/26.tif.

Tan:1982:BSL

[TH82] K. P. Tan and L. S. Hsu. Block sorting of a large file in external storage by a 2-component key. The Computer Journal, 25(3):327– 330, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/327.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/327.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/329.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/330.tif.

Thimbleby:1981:WBA

[Thi81] H. W. Thimbleby. A word boundary algorithm for text processing. The Computer Journal, 24(3):249–255, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/24/3/249.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/249.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/251.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/253.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/254.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/255.tif.

Thimbleby:1986:ELP

[Thi86] H. Thimbleby. Experiences of ‘literate programming’ using Cweb (a variant of Knuth’s Web). The Computer Journal, 29(3):201– 211, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/201.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/201.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/202.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/203.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/205.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/206.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/207.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/208.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/209.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/210.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/211.
tif.

Thimbleby:1987:OSR

[Thi87] H. Thimbleby. Optimising self-replicating programs. The Computer Journal, 30(5):475–476, October 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/5/475.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_05/tiff/475.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_05/tiff/476.tif.

Thoma:1980:CUB

[Tho80a] Martyn Thoma. Correspondence: The unconditional branch — a modest proposal. The Computer Journal, 23(4):380–381, November 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_04/tiff/380.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_04/

tiff/381.tif. See correspondence [Wig81b].

Thomas:1980:CS

[Tho80b] A. J. Thomas. Computer security. The Computer Journal, 23 (3):242a–242, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/3/242-a.full.pdf+html.

Thomas:1980:BRa

[Tho80c] J. O. Thomas. Book review. The Computer Journal, 23(1):84, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/84.2.full.pdf+html.

Thomas:1980:BRb

[Tho80d] M. Thomas. Book review. The Computer Journal, 23(1):84, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/84.3.full.pdf+html.

Thomas:1981:BR

[Tho81] R. E. Thomas. Book review. The Computer Journal, 24(4):342, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/342.4.full.pdf+html.

Treleaven:1982:CDF

[THR82] Philip C. Treleaven, Richard P. Hopkins, and Paul W. Rautenbach. Combining data flow and control flow computing. The Computer Journal, 25(2):207–217, May 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/207.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/207.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/208.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/209.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/210.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/211.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/212.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/213.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/214.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_25/Issue_02/tiff/215.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/216.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/217.tif.

Triance:1985:LEF

[TL85a] J. M. Triance and P. J. Layzell. Language enhancement facility for COBOL — its design and implementation. The Computer Journal, 28(2):128–133, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/2/128.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/128.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/129.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/130.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/131.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/132.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/133.tif.

Triance:1985:MPE

[TL85b] J. M. Triance and P. J. Layzell. Macro processors for enhancing high-level languages — some design principles. The Computer Journal, 28(1):34–43, February 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/28/1/34.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_28/Issue_01/tiff/34.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
01/tiff/36.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_01/tiff/37.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/38.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/40.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
41.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/43.tif.

Turner:1985:IFS

[TL85c] R. Turner and B. G. T. Lowden. An introduction to the formal specification of relational query languages. The Computer Journal, 28(2):162–169, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/2/162.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/165.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/166.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/167.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/168.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/169.tif.

Tomlinson:1982:AHS

[Tom82] D. J. Tomlinson. An aid to hidden surface removal in real time CGI systems. The Computer Journal, 25(4):429–441, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/4/429.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/429.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/430.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/431.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/432.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/433.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/434.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/435.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/436.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/437.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/438.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/439.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/440.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/441.tif.

Topor:1982:FPG

[Top82] R. W. Topor. Functional programs for generating permutations.

The Computer Journal, 25(2):257–263, May 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/257.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/257.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/258.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/259.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/260.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/261.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/262.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/263.tif.

Townsend:1987:C

[Tow87] H. R. A. Townsend. Correspondence. The Computer Journal, 30 (2):191, ???? 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/2/191.2.full.pdf+html.

Tse:1989:TFF

[TP89] T. H. Tse and L. Pong. Towards a formal foundation for De-Marco data flow diagrams. The Computer Journal, 32(1):1–12, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/1.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_32/Issue_01/tiff/1.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/10.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/11.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/12.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/2.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
4.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/tiff/
7.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/8.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/9.tif.

Travis:1982:APR

[Tra82] A. J. B. Travis. An aid to pattern recognition. The Computer Journal, 25(1):37–44, February 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/25/1/37.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_01/tiff/37.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/39.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/40.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/41.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/43.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/
44.tif.

Travis:1984:APR

[Tra84] A. J. B. Travis. An aid to pattern recognition. Part 2. The Computer Journal, 27(4):362–367, November 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/4/362.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/362.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/363.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/364.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_04/tiff/365.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_04/tiff/367.tif.

Treleaven:1982:WRV

[Tre82] Philip C. Treleaven. Workshop report: VLSI Machine Architectures and Very High Level Languages. The Computer Journal, 25(1):153–157, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/1/153.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/153.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/154.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/

tiff/155.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/156.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/157.
tif.

Triance:1980:SPC

[Tri80] J. M. Triance. Structured programming in COBOL — the current options. The Computer Journal, 23(3):194–200, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/194.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/194.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/195.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/196.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/
tiff/197.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/198.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/199.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/200.tif.

Triance:1983:PCS

[Tri83] J. M. Triance. The proposed COBOL Standard — its significance. The Computer Journal, 26(1):60–67, February 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/26/1/60.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/60.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/61.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/62.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/63.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/64.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/65.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/66.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
67.tif.

Tse:1987:IPU

[Tse87a] T. H. Tse. The identification of program unstructuredness: a formal approach. The Computer Journal, 30(6):507–511, December 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/6/507.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_06/tiff/507.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_06/tiff/508.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_06/tiff/509.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/510.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/511.tif.

Tse:1987:SNT

[Tse87b] T. H. Tse. Short notes: Towards a single criterion for identifying program unstructuredness*. The Computer Journal, 30(4):378– 380, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/378.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/378.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/379.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/380.tif.

Tsin:1986:C

[Tsi86] Y. H. Tsin. Correspondence. The Computer Journal, 29(4):382, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/4/382.full.pdf+html.

Tully:1985:IHA

[Tul85] C. J. Tully. Information, human activity and the nature of relevant theories. The Computer Journal, 28(3):206–210, July 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/3/206.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/206.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/207.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/208.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/209.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/210.tif.

Tyldesley:1988:EUE

[Tyl88] D. A. Tyldesley. Employing usability engineering in the development of office products. The Computer Journal, 31(5):431– 436, October 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/5/431.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/431.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_05/tiff/432.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_05/tiff/433.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_05/
tiff/434.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_05/tiff/435.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_05/tiff/436.
tif.

Uckan:1985:RSD

[Uck85] Yuksel Uckan. A relational schema description and manipulation facility. The Computer Journal, 28(2):170–178, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/2/170.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/170.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/171.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/172.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/173.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/174.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/175.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/176.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/177.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/178.tif.

Ullmann:1985:CAS

[UHS85] Julian R. Ullmann, Robert M. Haralick, and Linda G. Shapiro. Computer architecture for solving consistent labelling problems. The Computer Journal, 28(2):105–110, May 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/2/105.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/105.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/106.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/107.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_28/Issue_02/tiff/108.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/109.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/110.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/111.tif.

Ullmann:1988:FIR

[Ull88] J. R. Ullmann. Fast implementation of relational operations via inverse projections. The Computer Journal, 31(2):147– 154, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/2/147.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/147.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/148.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/149.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/150.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/151.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/152.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/153.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/154.tif.

Umphress:1989:SFP

[UPT89] D. A. Umphress, U. W. Pooch, and M. Tanik. Simulation: Fast prototyping of a goal-oriented simulation environment system. The Computer Journal, 32(6):541–548, December 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/6/541.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/541.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/542.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/543.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_06/
tiff/544.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_06/tiff/545.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_06/tiff/546.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_06/tiff/547.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_06/tiff/548.tif.

Unnikrishnan:1987:CCL

[UVS87] A. Unnikrishnan, Y. V. Venkatesh, and Priti Shankar. Connected component labelling using quadtrees — a bottomup approach. The Computer Journal, 30(2):176–182, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/30/2/176.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/176.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/177.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/180.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/181.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/182.tif.

Vajda:1983:NO

[Vaj83] S. Vajda. Nonlinear optimization 1981. The Computer Journal, 26(4):382a–382, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/26/4/382-a.full.pdf+html.

Vallance:1980:BR

[Val80] D. M. Vallance. Book review. The Computer Journal, 23(2): 152, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/152.1.full.pdf+html.

vanRijsbergen:1986:NCL

[Van86] C. J. Van Rijsbergen. A non-classical logic for information retrieval. The Computer Journal, 29(6):481–485, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/6/481.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_06/tiff/481.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_06/tiff/482.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_06/tiff/483.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_06/
tiff/484.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/485.tif.

Varol:1980:SAP

[Var80] Y. L. Varol. On a subset of all the permutations of n marks. The Computer Journal, 23(4):344–346, November 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_04/tiff/344.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_23/Issue_04/tiff/345.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_04/tiff/346.tif.

vanderWeide:1988:ITT

[vdW88] T. P. van der Weide. Iterative tree traversal using refinements and pointer rotation. The Computer Journal, 31(3):269– 275, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/3/269.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/270.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_03/tiff/271.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_03/
tiff/272.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_03/tiff/273.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_03/tiff/274.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_03/tiff/275.tif.

Velasevic:1982:RLC

[Vel82] Dusan M. Velasević. Right-to-left code generation for arithmetic expressions. The Computer Journal, 25(3):316–326, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/316.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/316.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/317.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/318.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/319.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/320.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/321.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/322.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_03/tiff/323.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/324.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/325.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/326.
tif.

Verkamo:1987:PQA

[Ver87] A. I. Verkamo. Performance of Quicksort adapted for virtual memory use. The Computer Journal, 30(4):362–371, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/4/362.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/362.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/363.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/364.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/365.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/366.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/367.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/368.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/369.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/370.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/371.tif.

Vallance:1980:BRB

[VP80] D. M. Vallance and Peter H. Prowse. Book reviews: Programming in Standard Fortran 77, by A. Balfour and D. Marwick, 1979; Issues in Data Base Management, by Herbert Weber and Anthony I. Wasserman, 1979. The Computer Journal, 23(2):152, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/152.tif.

Varol:1981:AGA

[VR81a] Y. L. Varol and D. Rotem. An algorithm to generate all topological sorting arrangements. The Computer Journal, 24(1):83–84, February 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/83.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/83.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/84.tif.

Varol:1981:SEJ

[VR81b] Yaakov L. Varol and Doron Rotem. Some experiences with the Johnson-Trotter permutation generation algorithm. The Computer Journal, 24(2):173–176, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/173.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/173.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/174.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/175.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/176.tif.

Verrijn-Stuart:1987:TTI

[VS87] A. Verrijn-Stuart. Themes and trends in information systems: TC8, 1975–1985. The Computer Journal, 30(2):97–109, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
30/2/97.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/100.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/102.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/105.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/106.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/107.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/108.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/109.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_02/tiff/97.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/
98.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_02/tiff/99.tif.

VanLeeuwen:1987:IR

[VT87] J. Van Leeuwen and R. B. Tan. Interval routing. The Computer Journal, 30(4):298–307, August 1987. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/4/298.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/298.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/299.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/301.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/302.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_04/tiff/304.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/305.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/306.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_04/tiff/307.
tif.

Wadler:1989:ELF

[Wad89] P. Wadler. Editorial — lazy functional programming. The Computer Journal, 32(2):97, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/2/97.full.pdf+html.

Wakeman:1981:C

[Wak81] M. Wakeman. Correspondence. The Computer Journal, 24(3): 270, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/270.3.full.pdf+html.

Walker:1980:BR

[Wal80] B. S. Walker. Book review. The Computer Journal, 23(1):60, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/1/60.2.full.pdf+html.

Walker:1981:BR

[Wal81] B. S. Walker. Book review. The Computer Journal, 24(4):342, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/342.2.full.pdf+html.

Wallis:1982:PGP

[Wal82] Peter J. L. Wallis. The preparation of guidelines for portable programming in high-level languages. The Computer Journal, 25(3): 375–378, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/3/375.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/375.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/376.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/377.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/378.tif.

Wand:1980:DRA

[Wan80] I. C. Wand. Dynamic resource allocation and supervision with the programming language MODULA. The Computer Journal, 23(2): 147–152, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/2/147.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/147.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/148.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_02/tiff/149.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/150.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/151.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_02/tiff/152.
tif.

Ward:1986:PCC

[War86] R. K. Ward. Parity check codes for logic processors. The Computer Journal, 29(1):12–16, February 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/29/1/12.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/12.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/14.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/16.
tif.

Watts:1980:DNA

[Wat80] K. Watts. Discussion on ‘New approaches to systems analysis and design’. The Computer Journal, 23(1):28–33, February 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/1/28.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/29.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/30.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/31.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
32.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/33.tif.

Watson:1981:CDD

[Wat81] D. F. Watson. Computing the n-dimensional Delaunay tessellation with application to Voronoĭ polytopes. The Computer Journal, 24(2):167–172, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/2/167.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/167.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_02/tiff/168.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_02/
tiff/169.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_02/tiff/170.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_02/tiff/171.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_02/tiff/172.tif.

Watt:1982:CPI

[Wat82] J. Watt. Correspondence: The possible impact of minicomputers and microprocessors [on] mainframe computer manufacturers. The Computer Journal, 25(2):286, May 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/286.tif. See [Hea81].

Williams:1985:RPP

[WC85] M. H. Williams and G. Chen. Restructuring Pascal programs containing GOTO statements. The Computer Journal, 28(2):134– 137, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/28/2/134.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/134.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/135.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/136.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/
tiff/137.tif.

Williams:1986:TPE

[WC86] M. H. Williams and G. Chen. Translating Pascal for execution on a Prolog-based system. The Computer Journal, 29(3):246– 252, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/246.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/246.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/247.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/248.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/249.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/250.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/251.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/252.tif.

Wise:1980:BRB

[WD80] A. H. Wise and A. Colin Day. Book reviews: Machine and Assembly Language Programming of the PDP-11, by Arthur Gill, 1979; Assembly Language Fundamentals, by Rina Yarmish and J. Yarmish, 1978. The Computer Journal, 23(2):164, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_02/tiff/164.tif.

Weldon:1980:UDB

[Wel80a] Jay Louise Weldon. Using data base abstractions for logical design. The Computer Journal, 23(1):41–45, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/23/1/41.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/41.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/42.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/43.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/44.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/45.

tif. See correspondence [JL80, Ree80c].

Wells:1980:BR

[Wel80b] M. Wells. Book review. The Computer Journal, 23(1):45, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
23/1/45.2.full.pdf+html.

Wells:1982:CN

[Wel82] M. Wells. Computer networks. The Computer Journal, 25 (3):337a–337, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/25/3/337-a.full.pdf+html.

Wells:1984:BRa

[Wel84a] M. Wells. Book review. The Computer Journal, 27(3):286, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
27/3/286.2.full.pdf+html.

Wells:1984:BRb

[Wel84b] M. Wells. Book review. The Computer Journal, 27(3):286, ???? 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
27/3/286.3.full.pdf+html.

Westreich:1980:EPC

[Wes80a] D. Westreich. An efficient predictor-corrector algorithm. The Computer Journal, 23(2):186, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/23/2/186.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

23/Issue_02/tiff/186.tif. See erratum [Ano80c] and correspondence [Sne80].

Westreich:1980:IPE

[Wes80b] D. Westreich. Improving polynomial evaluation at an approximate root. The Computer Journal, 23(2):187, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/2/187.
1.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_02/tiff/187.tif. See correspondence

[Bre81b, CF81].

Wettstein:1981:LOM

[Wet81] H. Wettstein. Locking operations for maximum concurrency. The Computer Journal, 24(3):243–248, August 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/24/3/243.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/243.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/244.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/246.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/247.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/248.tif.

Weyuker:1982:TNT

[Wey82] Elaine J. Weyuker. On testing non-testable programs. The Computer Journal, 25(4):465–470, November 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/4/465.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/465.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_04/tiff/466.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/467.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/468.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/469.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_04/tiff/470.tif.

Wray:1989:NSL

[WF89] S. C. Wray and J. Fairbairn. Non-strict languages — programming and implementation. The Computer Journal, 32(2):142– 151, April 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/2/142.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/142.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

32/Issue_02/tiff/143.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/144.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/145.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/146.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_02/tiff/147.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_02/tiff/148.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_02/tiff/149.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_02/
tiff/150.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_02/tiff/151.tif.

Wampler:1983:CEI

[WG83] S. B. Wampler and R. E. Griswold. Co-expressions in Icon. The Computer Journal, 26(1):72–78, February 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/26/1/72.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/76.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/78.tif.

Wade:1987:NAB

[WH87] S. Wade and J. Henstridge. Note on Algorithm 117: Buffered output in FORTRAN. The Computer Journal, 30(3):276, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_03/tiff/276.tif. See [HPB84].

Wheeler:1989:C

[Whe89] D. J. Wheeler. Correspondence. The Computer Journal, 32(1):95, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/32/1/95.1.full.pdf+html.

Wood-Harper:1982:TCA

[WHF82] A. T. Wood-Harper and G. Fitzgerald. A taxonomy of current approaches to systems analysis. The Computer Journal, 25(1):12–16, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/25/1/12.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_25/Issue_01/tiff/12.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
01/tiff/13.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/14.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/15.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/16.tif.

Wood-Harper:1983:ALT

[WHF83] A. T. Wood-Harper and D. J. Flynn. Action learning for teaching information systems. The Computer Journal, 26(1):79–82, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/1/79.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/79.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/80.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/82.
tif.

Wong:1984:PSR

[WI84] F. S. Wong and M. R. Ito. Parallel sorting on a re-circulating systolic sorter. The Computer Journal, 27(3):260–269, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/260.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/260.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/261.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/262.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/263.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/264.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/265.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/266.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_27/Issue_03/tiff/267.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/268.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/269.tif.

Wichmann:1982:CPA

[Wic82] B. A. Wichmann. A comparison of PASCAL and Ada. The Computer Journal, 25(2):248–252, May 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/2/248.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/248.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/249.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/250.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/251.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/252.
tif.

Wichmann:1987:NAR

[Wic87] B. A. Wichmann. Note on Algorithm 121: RSA key calculation in Ada. The Computer Journal, 30(3):276, June 1987. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/276.tif. See [Hun85].

Wichmann:1989:EFM

[Wic89a] B. Wichmann. Editorial — formal methods in computing. The Computer Journal, 32(5):385, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/5/385.full.pdf+html.

Wichmann:1989:TFS

[Wic89b] B. A. Wichmann. Towards a formal specification of floating point. The Computer Journal, 32(5):432–436, October 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/32/5/432.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_05/tiff/432.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_05/tiff/433.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_05/tiff/434.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_05/

tiff/435.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_05/tiff/436.tif.

Wigg:1981:C

[Wig81a] J. D. Wigg. Correspondance. The Computer Journal, 24(2):192, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/2/192.1.full.pdf+html.

Wigg:1981:CUB

[Wig81b] J. D. Wigg. Correspondence: Unconditional branches. The Computer Journal, 24(2):192, May 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
02/tiff/192.tif. See [Tho80a].

Williams:1980:IMS

[Wil80] I. R. Williams. Information Management Systems/Virtual Storage. The Computer Journal, 23(3):285, August 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/3/285.
full.pdf+html.

Williams:1982:CDF

[Wil82a] M. H. Williams. A comment on the decomposition of flowchart schemata. The Computer Journal, 25(3):393–396, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/25/3/393.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/393.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/394.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/395.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/396.tif. See [PG81].

Williams:1982:PLB

[Wil82b] M. H. Williams. The programming language BPL. The Computer Journal, 25(3):289–306, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/25/3/289.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/289.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_25/Issue_03/tiff/290.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/291.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/292.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/293.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/294.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/295.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/296.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/297.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/298.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/299.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/300.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/301.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/302.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_03/tiff/303.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_03/tiff/304.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/305.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_03/
tiff/306.tif.

Williams:1983:BRa

[Wil83a] I. Williams. Book review. The Computer Journal, 26(2):191, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/2/191.1.full.pdf+html.

Williams:1983:BRb

[Wil83b] I. Williams. Book review. The Computer Journal, 26(2):191, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/2/191.4.full.pdf+html.

Williams:1983:FSP

[Wil83c] M. H. Williams. Flowchart schemata and the problem of nomenclature. The Computer Journal, 26(3):270–276, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/270.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/270.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/271.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/272.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/273.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/274.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/275.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/276.tif.

Wilson:1983:AIP

[Wil83d] J. M. Wilson. Algorithm 114: Interrupted permutations in lexicographic order. The Computer Journal, 26(1):92, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilson:1983:AS

[Wil83e] J. M. Wilson. Algorithms supplement. The Computer Journal, 26 (1):92, ???? 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/1/92.full.pdf+html.

Wilkes:1984:SMP

[Wil84] Maurice V. Wilkes. Security management and protection — a personal approach. The Computer Journal, 27(1):3–7, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/27/1/3.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_27/Issue_01/tiff/3.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_
01/tiff/4.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_01/tiff/5.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_27/Issue_01/tiff/6.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_27/
Issue_01/tiff/7.tif.

Williams:1985:LAC

[Wil85a] M. H. Williams. A linear algorithm for colouring planar graphs with five colours. The Computer Journal, 28(1):78–81, February 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
28/1/78.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_28/Issue_01/tiff/78.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_01/tiff/
79.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_01/tiff/81.tif.

Williams:1985:STL

[Wil85b] M. H. Williams. Structuring two-level grammar specifications. The Computer Journal, 28(3):250–256, July 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/3/250.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/251.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_03/
tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_03/tiff/253.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_03/tiff/255.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_03/tiff/256.tif.

Wilson:1986:SNW

[Wil86a] J. M. Wilson. Short note: a weighted mark-scaling algorithm. The Computer Journal, 29(5):477, ???? 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/29/5/477.full.pdf+html.

Wilson:1986:WMS

[Wil86b] J. M. Wilson. A weighted mark-scaling algorithm. The Computer Journal, 29(5):477, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
05/tiff/477.tif.

Williams:1988:GQ

[Wil88] R. Williams. The Goblin quadtree. The Computer Journal, 31(4): 358–363, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/4/358.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/358.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/359.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_31/Issue_04/tiff/360.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/361.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/362.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/363.
tif.

Wilson:1989:CCU

[Wil89] J. M. Wilson. Crossword compilation using integer programming. The Computer Journal, 32(3):273–275, June 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/32/3/273.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_03/tiff/273.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/274.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/275.tif.

Wise:1980:BR

[Wis80] A. H. Wise. Book review. The Computer Journal, 23(2):164, ???? 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/2/164.1.full.pdf+html.

Wise:1981:BR

[Wis81] A. H. Wise. Book review. The Computer Journal, 24(1):77, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/1/77.2.full.pdf+html.

Wise:1982:BRB

[Wis82] A. H. Wise. Book review: Advances in Data Processing Management — Vol.1. The Computer Journal, 25(3):400, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
25/3/400.full.pdf+html.

Witten:1980:AAL

[Wit80] I. H. Witten. Algorithms for adaptive linear prediction. The Computer Journal, 23(1):78–84, February 1980. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/23/1/78.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_23/Issue_01/tiff/78.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/79.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/80.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/82.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/84.tif.

Wey:1987:NST

[WL87] C.-L. Wey and F. Lombardi. On a novel self-test approach to digital testing. The Computer Journal, 30(3):258–267, June 1987. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/30/3/258.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/258.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/259.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/260.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/261.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/262.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_03/tiff/263.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
30/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_30/Issue_03/tiff/265.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_30/Issue_03/
tiff/266.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_03/tiff/267.tif.

Wlodarczyk:1989:C

[Wlo89] A. J. Wlodarczyk. Correspondence. The Computer Journal, 32(4): 380–381, ???? 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/380.2.full.pdf+html.

Watt:1983:EAG

[WM83] David A. Watt and Ole Lehrmann Madsen. Extended attribute grammars. The Computer Journal, 26(2):142–153, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/2/142.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/142.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/143.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/144.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/145.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/146.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/147.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_02/tiff/148.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/149.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_02/
tiff/150.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_02/tiff/151.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_02/tiff/152.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

26/Issue_02/tiff/153.tif. See also: report 10, Computer Science Department, University of Glasgow (July 1977) and report DAIMI PB-105, Computer Science Department, Aarhus University (November 1979).

Williams:1984:PAC

[WM84] M. H. Williams and K. T. Milne. The performance of algorithms for colouring planar graphs. The Computer Journal, 27(2):165– 170, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/2/165.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/165.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/166.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/167.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/168.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/169.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/170.
tif.

Witten:1981:ILF

[WN81] I. H. Witten and Y. H. Ng. An ideographic language front end processor for accessing English language computer systems. The Computer Journal, 24(1):62–70, February 1981. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/24/1/62.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_24/Issue_01/tiff/62.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_24/Issue_01/tiff/63.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_
01/tiff/64.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_01/tiff/65.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/66.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/67.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_01/tiff/68.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_24/Issue_01/tiff/
69.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_01/tiff/70.tif.

Wiseman:1986:CRF

[WN86] N. E. Wiseman and S. Nedunuri. Computing random fields. The Computer Journal, 29(4):373–377, August 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/4/373.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/373.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/374.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/375.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/376.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/377.
tif.

Williams:1988:AHI

[WN88] M. H. Williams and K. A. Nicholson. An approach to handling incomplete information in databases. The Computer Journal, 31(2): 133–140, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/2/133.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/133.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/134.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/135.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/136.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/137.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/138.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/139.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/140.tif.

Wootton:1981:BR

[Woo81] N. P. Wootton. Book review. The Computer Journal, 24(4):342, ???? 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/24/4/342.3.full.pdf+html.

Woodwark:1982:EQT

[Woo82] J. R. Woodwark. The explicit quad tree as a structure for computer graphics. The Computer Journal, 25(2):235– 238, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/2/235.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/235.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/236.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/237.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/

tiff/238.tif. See comment [BK83a].

Woodwark:1984:CQT

[Woo84] J. R. Woodwark. Compressed quad trees. The Computer Journal, 27(3):225–229, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/3/225.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/225.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/226.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/227.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/228.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/229.
tif.

Woodall:1986:TPD

[Woo86] A. D. Woodall. Techniques of program design. The Computer Journal, 29(5):434–435, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/5/434.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_05/tiff/434.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_05/tiff/435.tif.

Williams:1986:SID

[WR86] M. H. Williams and J. Riddall. The segregation of incompatible duties. The Computer Journal, 29(1):9–11, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/29/1/9.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_01/tiff/9.tif.

Wright:1982:SVD

[Wri82] William E. Wright. Single versus double buffering in constrained merging. The Computer Journal, 25(2):227–230, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/2/227.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/227.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/228.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/229.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/230.tif.

Wright:1983:SFS

[Wri83] William E. Wright. Some file structure considerations pertaining to magnetic bubble memory. The Computer Journal, 26(1):43–51, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/26/1/43.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_26/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_01/tiff/45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/46.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_01/tiff/
49.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_01/tiff/51.tif.

Wong:1986:SRD

[WS86] E. Wong and W. B. Samson. The specification of a relational database (PRECI) as an abstract data type and its realisation in HOPE. The Computer Journal, 29(3):261–268, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/3/261.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/261.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/262.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/263.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/
tiff/264.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/265.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/266.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_03/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/268.tif.

Wilkes:1984:RW

[WSG^{+}84] A. J. Wilkes, D. W. Singer, J. J. Gibbons, T. R. King, P. Robinson, and N. E. Wiseman. The Rainbow Workstation. The Computer Journal, 27(2):112–120, May 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/27/2/112.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/112.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/113.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/114.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/115.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_02/tiff/116.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_02/tiff/117.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_02/tiff/118.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_02/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_02/tiff/120.tif.

Ward:1984:ECD

[WT84] R. K. Ward and M. Tabandeh. Error correction and detection, a geometric approach. The Computer Journal, 27(3):246– 253, August 1984. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/3/246.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/246.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/247.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/248.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_03/
tiff/249.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_27/Issue_03/tiff/250.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_27/Issue_03/tiff/251.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
27/Issue_03/tiff/252.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_03/tiff/253.tif.

Wilkes:1982:SEC

[WW82] A. J. Wilkes and N. E. Wiseman. A soft-edged character set and its derivation. The Computer Journal, 25(1):140–147, February 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/1/140.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/140.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/141.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/142.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_01/
tiff/143.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_01/tiff/144.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_01/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_01/tiff/146.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_01/tiff/147.tif.

Woods:1983:MAM

[WW83] J. V. Woods and A. J. T. Wheen. Mu6P: an advanced microprocessor architecture. The Computer Journal, 26(3):208– 217, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/3/208.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/208.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/209.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/210.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/

tiff/211.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/212.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/213.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
26/Issue_03/tiff/214.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/215.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_03/
tiff/216.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/217.tif.

Yannakoudakis:1980:CCB

[YAH80] E. J. Yannakoudakis, F. H. Ayres, and J. A. W. Huggill. Character coding for bibliographical record control. The Computer Journal, 23(1):53–60, February 1980. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/23/1/53.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/55.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/57.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/59.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
60.tif.

Yelland:1989:FST

[Yel89] P. M. Yelland. First steps towards fully abstract semantics for object-oriented languages. The Computer Journal, 32(4):290– 296, August 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/4/290.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/290.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/291.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_04/tiff/292.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_04/
tiff/293.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_04/tiff/294.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_04/tiff/295.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_04/tiff/296.tif.

Young:1986:UEF

[YP86] S. J. Young and C. Proctor. UFL: an experimental frame language based on abstract data types. The Computer Journal, 29(4):340– 347, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/29/4/340.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/343.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/344.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_04/tiff/345.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
29/Issue_04/tiff/346.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_04/tiff/347.tif.

Yannakoudakis:1982:QEG

[YW82] E. J. Yannakoudakis and A. K. P. Wu. Quasi-equifrequent group generation and evaluation. The Computer Journal, 25(2):183– 187, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/25/2/183.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_02/tiff/183.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
25/Issue_02/tiff/184.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_02/tiff/185.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_02/
tiff/186.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_02/tiff/187.tif.

Yu:1989:DEB

[YY89] K. W. Yu and T. L. Yu. Data encryption based upon time reversal transformations. The Computer Journal, 32(3):241– 245, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/3/241.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/241.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

32/Issue_03/tiff/242.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_03/tiff/243.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/244.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/245.tif.

Zahran:1981:UDM

[Zah81] F. S. Zahran. Universal data model holder for data dictionary systems. The Computer Journal, 24(3):193–199, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/193.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/193.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/195.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/197.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/198.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/199.tif.

Zaki:1984:ASS

[Zak84] Ahmed S. Zaki. Algorithm 119: a space saving insertion algorithm for 2-3 trees. The Computer Journal, 27(4):368–372, November 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_27/Issue_04/tiff/368.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_27/Issue_04/tiff/369.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_27/
Issue_04/tiff/370.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_27/Issue_04/tiff/371.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_27/Issue_04/
tiff/372.tif.

Zaremba:1980:DCA

[Zar80] W. A. Zaremba. Discussion and correspondence: An alternative for the ‘completer’ function. The Computer Journal, 23(3):286– 287, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/23/3/286.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/286.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

23/Issue_03/tiff/287.tif. See correspondence [Hun81b].

Zdrahal:1981:RCP

[ZBS81] Z. Zdrahal, I. Bratko, and A. Shapiro. Recognition of complex patterns using cellular arrays. The Computer Journal, 24(3):263– 270, August 1981. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/24/3/263.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/263.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/265.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_24/Issue_03/
tiff/266.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_24/Issue_03/tiff/267.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_24/Issue_03/tiff/268.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
24/Issue_03/tiff/269.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_03/tiff/270.tif.

Zaki:1988:ARM

[ZE88] M. Zaki and M. M. Elboraey. Analysis of reliability models for interconnecting MIMD systems. The Computer Journal, 31(4):304– 312, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/31/4/304.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/304.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/305.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/306.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/307.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/308.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/309.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/310.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/311.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/312.tif.

Zivani:1985:AIS

[ZOG85] Nivio Zivani, Henk J. Olivié, and Gaston H. Gonnet. The analysis of an improved symmetric binary B-tree algorithm. The Computer Journal, 28(4):417–425, August 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/28/4/417.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/417.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/418.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/419.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/420.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_28/Issue_04/tiff/421.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_04/tiff/422.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_04/tiff/423.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_04/
tiff/424.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_04/tiff/425.tif.

Zhao:1988:OOD

[ZR88] Liping Zhao and S. A. Roberts. An object-oriented data model for database modelling, implementation and access. The Computer Journal, 31(2):116–124, April 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/31/2/116.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/117.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/119.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_31/Issue_02/tiff/120.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_02/tiff/121.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_02/tiff/122.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_02/
tiff/123.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_02/tiff/124.tif.

Zvegintzov:1980:PMR

[Zve80] N. Zvegintzov. Partial-match retrieval in an index sequential directory. The Computer Journal, 23(1):37–40, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/23/1/37.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_23/Issue_01/tiff/37.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/39.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/40.tif.